GNU Automake

For version 1.8.5, 13 May 2004

David MacKenzie Tom Tromey Alexandre Duret-Lutz

This manual is for GNU Automake (version 1.8.5, 13 May 2004), a program which creates GNU standards-compliant Makefiles from template files. Copyright \begin{matrix}c \\ ◯ \end{matrix} 1995, 1996, 1997, 1998, 1999, 2000, 2001, 2002, 2003, 2004 Free Software Foundation, Inc.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.1 or any later version published by the Free Software Foundation; with no Invariant Sections, with the Front-Cover texts being “A GNU Manual,” and with the Back-Cover Texts as in (a) below. A copy of the license is included in the section entitled “GNU Free Documentation License.”

(a) The FSF’s Back-Cover Text is: “You have freedom to copy and modify this GNU Manual, like GNU software. Copies published by the Free Software Foundation raise funds for GNU development.”

1 Introduction

Automake is a tool for automatically generating Makefile.ins from files called Makefile.am. Each Makefile.am is basically a series of make variable definitions^{1}, with rules being thrown in occasionally. The generated Makefile.ins are compliant with the GNU Makefile standards.

The GNU Makefile Standards Document (see Section “Makefile Conventions” in The GNU Coding Standards) is long, complicated, and subject to change. The goal of Automake is to remove the burden of Makefile maintenance from the back of the individual GNU maintainer (and put it on the back of the Automake maintainer).

The typical Automake input file is simply a series of variable definitions. Each such file is processed to create a Makefile.in. There should generally be one Makefile.am per directory of a project.

Automake does constrain a project in certain ways; for instance it assumes that the project uses Autoconf (see Section “Introduction” in The Autoconf Manual), and enforces certain restrictions on the configure.ac contents^{2}.

Automake requires perl in order to generate the Makefile.ins. However, the distributions created by Automake are fully GNU standards-compliant, and do not require perl in order to be built.

Mail suggestions and bug reports for Automake to bug-automake@gnu.org.

2 General ideas

The following sections cover a few basic ideas that will help you understand how Automake works.

2.1 General Operation

Automake works by reading a Makefile.am and generating a Makefile.in. Certain variables and rules defined in the Makefile.am instruct Automake to generate more specialized code; for instance, a ‘bin_PROGRAMS’ variable definition will cause rules for compiling and linking programs to be generated.

The variable definitions and rules in the Makefile.am are copied verbatim into the generated file. This allows you to add arbitrary code into the generated Makefile.in. For instance the Automake distribution includes a non-standard rule for the cvs-dist target, which the Automake maintainer uses to make distributions from his source control system.

Note that most GNU make extensions are not recognized by Automake. Using such extensions in a Makefile.am will lead to errors or confusing behavior.

A special exception is that the GNU make append operator, ‘+=’, is supported. This operator appends its right hand argument to the variable specified on the left. Automake will translate the operator into an ordinary ‘=’ operator; ‘+=’ will thus work with any make program.

1 These variables are also called make macros in Make terminology, however in this manual we reserve the

term macro for Autoconf’s macros.

2 Older Autoconf versions used configure.in. Autoconf 2.50 and greater promotes configure.ac over

configure.in. The rest of this documentation will refer to configure.ac, but Automake also supports

configure.in for backward compatibility.

Automake tries to keep comments grouped with any adjoining rules or variable definitions.

A rule defined in Makefile.am generally overrides any such rule of a similar name that would be automatically generated by automake. Although this is a supported feature, it is generally best to avoid making use of it, as sometimes the generated rules are very particular.

Similarly, a variable defined in Makefile.am or AC_SUBST’ed from configure.ac will override any definition of the variable that automake would ordinarily create. This feature is more often useful than the ability to override a rule. Be warned that many of the variables generated by automake are considered to be for internal use only, and their names might change in future releases.

When examining a variable definition, Automake will recursively examine variables referenced in the definition. For example, if Automake is looking at the content of foo_SOURCES in this snippet

     xs = a.c b.c
     foo_SOURCES = c.c $(xs)

it would use the files a.c, b.c, and c.c as the contents of foo_SOURCES.

Automake also allows a form of comment which is not copied into the output; all lines beginning with ‘##’ (leading spaces allowed) are completely ignored by Automake.

It is customary to make the first line of Makefile.am read:

     ## Process this file with automake to produce Makefile.in

2.2 Strictness

While Automake is intended to be used by maintainers of GNU packages, it does make some effort to accommodate those who wish to use it, but do not want to use all the GNU conventions.

To this end, Automake supports three levels of strictness—the strictness indicating how stringently Automake should check standards conformance.

The valid strictness levels are:

‘foreign’ Automake will check for only those things which are absolutely required for proper operations. For instance, whereas GNU standards dictate the existence of a NEWS file, it will not be required in this mode. The name comes from the fact that Automake is intended to be used for GNU programs; these relaxed rules are not the standard mode of operation.

‘gnu’ Automake will check—as much as possible—for compliance to the GNU standards for packages. This is the default.

‘gnits’ Automake will check for compliance to the as-yet-unwritten Gnits standards. These are based on the GNU standards, but are even more detailed. Unless you are a Gnits standards contributor, it is recommended that you avoid this option until such time as the Gnits standard is actually published (which may never happen).

For more information on the precise implications of the strictness level, see Chapter 21 [Gnits], page 73.

Automake also has a special “cygnus” mode which is similar to strictness but handled differently. This mode is useful for packages which are put into a “Cygnus” style tree (e.g., the GCC tree). For more information on this mode, see Chapter 22 [Cygnus], page 74.

2.3 The Uniform Naming Scheme

Automake variables generally follow a uniform naming scheme that makes it easy to decide how programs (and other derived objects) are built, and how they are installed. This scheme also supports configure time determination of what should be built.

At make time, certain variables are used to determine which objects are to be built. The variable names are made of several pieces which are concatenated together.

The piece which tells automake what is being built is commonly called the primary. For instance, the primary PROGRAMS holds a list of programs which are to be compiled and linked.

A different set of names is used to decide where the built objects should be installed. These names are prefixes to the primary which indicate which standard directory should be used as the installation directory. The standard directory names are given in the GNU standards (see Section “Directory Variables” in The GNU Coding Standards). Automake extends this list with pkglibdir, pkgincludedir, and pkgdatadir; these are the same as the non-‘pkg’ versions, but with ‘$(PACKAGE)’ appended. For instance, pkglibdir is

defined as $(libdir)/$(PACKAGE).

For each primary, there is one additional variable named by prepending ‘EXTRA_’ to the primary name. This variable is used to list objects which may or may not be built, depending on what configure decides. This variable is required because Automake must statically know the entire list of objects that may be built in order to generate a Makefile.in that will work in all cases.

For instance, cpio decides at configure time which programs are built. Some of the programs are installed in bindir, and some are installed in sbindir:

     EXTRA_PROGRAMS = mt rmt
     bin_PROGRAMS = cpio pax
     sbin_PROGRAMS = $(MORE_PROGRAMS)

Defining a primary without a prefix as a variable, e.g., PROGRAMS, is an error.

Note that the common ‘dir’ suffix is left off when constructing the variable names; thus

one writes ‘bin_PROGRAMS’ and not ‘bindir_PROGRAMS’.

Not every sort of object can be installed in every directory. Automake will flag those attempts it finds in error. Automake will also diagnose obvious misspellings in directory names.

Sometimes the standard directories—even as augmented by Automake— are not enough. In particular it is sometimes useful, for clarity, to install objects in a subdirectory of some predefined directory. To this end, Automake allows you to extend the list of possible installation directories. A given prefix (e.g. ‘zar’) is valid if a variable of the same name with ‘dir’ appended is defined (e.g. zardir).

For instance, installation of HTML files is part of Automake, you could use this to install raw HTML documentation:

     htmldir = $(prefix)/html
     html_DATA = automake.html

The special prefix ‘noinst’ indicates that the objects in question should be built but not installed at all. This is usually used for objects required to build the rest of your package, for instance static libraries (see Section 8.2 [A Library], page 30), or helper scripts.

The special prefix ‘check’ indicates that the objects in question should not be built until the make check command is run. Those objects are not installed either.

The current primary names are ‘PROGRAMS’, ‘LIBRARIES’, ‘LISP’, ‘PYTHON’, ‘JAVA’,

‘SCRIPTS’, ‘DATA’, ‘HEADERS’, ‘MANS’, and ‘TEXINFOS’.

Some primaries also allow additional prefixes which control other aspects of automake’s behavior. The currently defined prefixes are ‘dist_’, ‘nodist_’, and ‘nobase_’. These prefixes are explained later (see Section 8.4 [Program and Library Variables], page 35).

2.4 How derived variables are named

Sometimes a Makefile variable name is derived from some text the maintainer supplies. For instance, a program name listed in ‘_PROGRAMS’ is rewritten into the name of a ‘_SOURCES’ variable. In cases like this, Automake canonicalizes the text, so that program names and the like do not have to follow Makefile variable naming rules. All characters in the name except for letters, numbers, the strudel (@), and the underscore are turned into underscores when making variable references.

For example, if your program is named sniff-glue, the derived variable name would be sniff_glue_SOURCES, not sniff-glue_SOURCES. Similarly the sources for a library named

libmumble++.a should be listed in the libmumble___a_SOURCES variable.

The strudel is an addition, to make the use of Autoconf substitutions in variable names less obfuscating.

2.5 Variables reserved for the user

Some Makefile variables are reserved by the GNU Coding Standards for the use of the “user” – the person building the package. For instance, CFLAGS is one such variable.

Sometimes package developers are tempted to set user variables such as CFLAGS because it appears to make their job easier. However, the package itself should never set a user variable, particularly not to include switches which are required for proper compilation of the package. Since these variables are documented as being for the package builder, that person rightfully expects to be able to override any of these variables at build time.

To get around this problem, automake introduces an automake-specific shadow variable for each user flag variable. (Shadow variables are not introduced for variables like CC, where they would make no sense.) The shadow variable is named by prepending ‘AM_’ to the user variable’s name. For instance, the shadow variable for YFLAGS is AM_YFLAGS.

2.6 Programs automake might require

Automake sometimes requires helper programs so that the generated Makefile can do its work properly. There are a fairly large number of them, and we list them here.

ansi2knr.c
ansi2knr.1

These two files are used by the automatic de-ANSI-fication support (see Section 8.14 [ANSI], page 47).

compile This is a wrapper for compilers which don’t accept both ‘-c’ and ‘-o’ at the same time. It is only used when absolutely required. Such compilers are rare.

config.guess
config.sub

These programs compute the canonical triplets for the given build, host, or target architecture. These programs are updated regularly to support new architectures and fix probes broken by changes in new kernel versions. You are encouraged to fetch the latest versions of these files from ftp://ftp.gnu.org/ gnu/config/ before making a release.

depcomp This program understands how to run a compiler so that it will generate not only the desired output but also dependency information which is then used by the automatic dependency tracking feature.

elisp-comp

This program is used to byte-compile Emacs Lisp code.

install-sh

This is a replacement for the install program which works on platforms where install is unavailable or unusable.

mdate-sh This script is used to generate a version.texi file. It examines a file and prints some date information about it.

missing This wraps a number of programs which are typically only required by maintainers. If the program in question doesn’t exist, missing prints an informative warning and attempts to fix things so that the build can continue.

mkinstalldirs

This script used to be a wrapper around mkdir -p, which is not portable. Now we use prefer to use install-sh -d when configure finds that mkdir -p does not work, this makes one less script to distribute. For backward compatibility mkinstalldirs is still used and distributed when automake finds it in a package. But it is no longer installed automatically, and it should be safe to remove it.

py-compile

This is used to byte-compile Python scripts.

texinfo.tex

Not a program, this file is required for make dvi, make ps and make pdf to work when Texinfo sources are in the package. ylwrap This program wraps lex and yacc and ensures that, for instance, multiple yacc

instances can be invoked in a single directory in parallel.

3 Some example packages

3.1 A simple example, start to finish

Let’s suppose you just finished writing zardoz, a program to make your head float from vortex to vortex. You’ve been using Autoconf to provide a portability framework, but your Makefile.ins have been ad-hoc. You want to make them bulletproof, so you turn to Automake.

The first step is to update your configure.ac to include the commands that automake needs. The way to do this is to add an AM_INIT_AUTOMAKE call just after AC_INIT:

     AC_INIT(zardoz, 1.0)
     AM_INIT_AUTOMAKE
     ...

Since your program doesn’t have any complicating factors (e.g., it doesn’t use gettext, it doesn’t want to build a shared library), you’re done with this part. That was easy!

Now you must regenerate configure. But to do that, you’ll need to tell autoconf how to find the new macro you’ve used. The easiest way to do this is to use the aclocal program to generate your aclocal.m4 for you. But wait… maybe you already have an aclocal.m4, because you had to write some hairy macros for your program. The aclocal program lets you put your own macros into acinclude.m4, so simply rename and then run:

     mv aclocal.m4 acinclude.m4
     aclocal
     autoconf

Now it is time to write your Makefile.am for zardoz. Since zardoz is a user program, you want to install it where the rest of the user programs go: bindir. Additionally, zardoz has some Texinfo documentation. Your configure.ac script uses AC_REPLACE_FUNCS, so you need to link against ‘$(LIBOBJS)’. So here’s what you’d write:

     bin_PROGRAMS = zardoz
     zardoz_SOURCES = main.c head.c float.c vortex9.c gun.c
     zardoz_LDADD = $(LIBOBJS)
     info_TEXINFOS = zardoz.texi

Now you can run automake --add-missing to generate your Makefile.in and grab any auxiliary files you might need, and you’re done!

3.2 A classic program

GNU hello (ftp://prep.ai.mit.edu/pub/gnu/hello-1.3.tar.gz) is renowned for its

classic simplicity and versatility. This section shows how Automake could be used with the GNU Hello package. The examples below are from the latest beta version of GNU Hello, but with all of the maintainer-only code stripped out, as well as all copyright comments.

Of course, GNU Hello is somewhat more featureful than your traditional two-liner. GNU Hello is internationalized, does option processing, and has a manual and a test suite.

Here is the configure.ac from GNU Hello. Please note: The calls to AC_INIT and AM_INIT_AUTOMAKE in this example use a deprecated syntax. For the current approach, see the description of AM_INIT_AUTOMAKE in Section 5.6.1 [Public macros], page 18.

dnl Process this file with autoconf to produce a configure script.
AC_INIT(src/hello.c)
AM_INIT_AUTOMAKE(hello, 1.3.11)
AM_CONFIG_HEADER(config.h)
dnl Set of available languages.
ALL_LINGUAS="de fr es ko nl no pl pt sl sv"
dnl Checks for programs.
AC_PROG_CC
AC_ISC_POSIX
dnl Checks for libraries.
dnl Checks for header files.
AC_STDC_HEADERS
AC_HAVE_HEADERS(string.h fcntl.h sys/file.h sys/param.h)
dnl Checks for library functions.
AC_FUNC_ALLOCA
dnl Check for st_blksize in struct stat
AC_ST_BLKSIZE
dnl internationalization macros
AM_GNU_GETTEXT
AC_OUTPUT([Makefile doc/Makefile intl/Makefile po/Makefile.in \
           src/Makefile tests/Makefile tests/hello],
   [chmod +x tests/hello])

The ‘AM_’ macros are provided by Automake (or the Gettext library); the rest are standard Autoconf macros.

The top-level Makefile.am:

EXTRA_DIST = BUGS ChangeLog.O
SUBDIRS = doc intl po src tests

As you can see, all the work here is really done in subdirectories. The po and intl directories are automatically generated using gettextize; they will

not be discussed here.

In doc/Makefile.am we see:
info_TEXINFOS = hello.texi
hello_TEXINFOS = gpl.texi

This is sufficient to build, install, and distribute the GNU Hello manual.

   Here is tests/Makefile.am:
     TESTS = hello
     EXTRA_DIST = hello.in testdata

The script hello is generated by configure, and is the only test case. make check will run this test.

Last we have src/Makefile.am, where all the real work is done:

     bin_PROGRAMS = hello
     hello_SOURCES = hello.c version.c getopt.c getopt1.c getopt.h system.h
     hello_LDADD = $(INTLLIBS) $(ALLOCA)
     localedir = $(datadir)/locale
     INCLUDES = -I../intl -DLOCALEDIR=\"$(localedir)\"

3.3 Building true and false

Here is another, trickier example. It shows how to generate two programs (true and false) from the same source file (true.c). The difficult part is that each compilation of true.c requires different cpp flags.

     bin_PROGRAMS = true false
     false_SOURCES =
     false_LDADD = false.o
     true.o: true.c
             $(COMPILE) -DEXIT_CODE=0 -c true.c
     false.o: true.c
             $(COMPILE) -DEXIT_CODE=1 -o false.o -c true.c

Note that there is no true_SOURCES definition. Automake will implicitly assume that there is a source file named true.c, and define rules to compile true.o and link true. The true.o: true.c rule supplied by the above Makefile.am, will override the Automake generated rule to build true.o.

false_SOURCES is defined to be empty—that way no implicit value is substituted. Because we have not listed the source of false, we have to tell Automake how to link the program. This is the purpose of the false_LDADD line. A false_DEPENDENCIES variable, holding the dependencies of the false target will be automatically generated by Automake from the content of false_LDADD.

The above rules won’t work if your compiler doesn’t accept both ‘-c’ and ‘-o’. The simplest fix for this is to introduce a bogus dependency (to avoid problems with a parallel

make):
     true.o: true.c false.o
             $(COMPILE) -DEXIT_CODE=0 -c true.c
     false.o: true.c
             $(COMPILE) -DEXIT_CODE=1 -c true.c && mv true.o false.o

Also, these explicit rules do not work if the de-ANSI-fication feature is used (see Section 8.14 [ANSI], page 47). Supporting de-ANSI-fication requires a little more work:

     true_.o: true_.c false_.o
             $(COMPILE) -DEXIT_CODE=0 -c true_.c
     false_.o: true_.c
             $(COMPILE) -DEXIT_CODE=1 -c true_.c && mv true_.o false_.o

As it turns out, there is also a much easier way to do this same task. Some of the above techniques are useful enough that we’ve kept the example in the manual. However if you were to build true and false in real life, you would probably use per-program compilation flags, like so:

     bin_PROGRAMS = false true
     false_SOURCES = true.c
     false_CPPFLAGS = -DEXIT_CODE=1
     true_SOURCES = true.c
     true_CPPFLAGS = -DEXIT_CODE=0

In this case Automake will cause true.c to be compiled twice, with different flags. De-ANSI-fication will work automatically. In this instance, the names of the object files would be chosen by automake; they would be false-true.o and true-true.o. (The name of the object files rarely matters.)

4 Creating a Makefile.in

To create all the Makefile.ins for a package, run the automake program in the top level directory, with no arguments. automake will automatically find each appropriate Makefile.am (by scanning configure.ac; see Chapter 5 [configure], page 12) and generate the corresponding Makefile.in. Note that automake has a rather simplistic view of what constitutes a package; it assumes that a package has only one configure.ac, at the top. If your package has multiple configure.acs, then you must run automake in each directory holding a configure.ac. (Alternatively, you may rely on Autoconf’s autoreconf, which is able to recurse your package tree and run automake where appropriate.)

You can optionally give automake an argument; .am is appended to the argument and the result is used as the name of the input file. This feature is generally only used to automatically rebuild an out-of-date Makefile.in. Note that automake must always be run from the topmost directory of a project, even if being used to regenerate the Makefile.in in some subdirectory. This is necessary because automake must scan configure.ac, and because automake uses the knowledge that a Makefile.in is in a subdirectory to change its behavior in some cases.

Automake will run autoconf to scan configure.ac and its dependencies (aclocal.m4), therefore autoconf must be in your PATH. If there is an AUTOCONF variable in your environment it will be used instead of autoconf, this allows you to select a particular version of Autoconf. By the way, don’t misunderstand this paragraph: Automake runs autoconf to scan your configure.ac, this won’t build configure and you still have to run autoconf yourself for this purpose.

automake accepts the following options:

‘-a’
‘--add-missing’

Automake requires certain common files to exist in certain situations; for in-

stance config.guess is required if configure.ac runs AC_CANONICAL_HOST.

Automake is distributed with several of these files (see Section 2.6 [Auxiliary Programs], page 5); this option will cause the missing ones to be automatically added to the package, whenever possible. In general if Automake tells you a file is missing, try using this option. By default Automake tries to make a symbolic link pointing to its own copy of the missing file; this can be changed

with --copy.

Many of the potentially-missing files are common scripts whose location may be

specified via the AC_CONFIG_AUX_DIR macro. Therefore, AC_CONFIG_AUX_DIR’s

setting affects whether a file is considered missing, and where the missing file is added (see Section 5.2 [Optional], page 13).

‘--libdir=dir’

Look for Automake data files in directory dir instead of in the installation directory. This is typically used for debugging.

‘-c’

‘--copy’ When used with --add-missing, causes installed files to be copied. The default is to make a symbolic link.

‘--cygnus’

Causes the generated Makefile.ins to follow Cygnus rules, instead of GNU or Gnits rules. For more information, see Chapter 22 [Cygnus], page 74.

‘-f’
‘--force-missing’

When used with --add-missing, causes standard files to be reinstalled even if they already exist in the source tree. This involves removing the file from the source tree before creating the new symlink (or, with --copy, copying the new file).

‘--foreign’

Set the global strictness to ‘foreign’. For more information, see Section 2.2 [Strictness], page 2.

‘--gnits’ Set the global strictness to ‘gnits’. For more information, see Chapter 21 [Gnits], page 73.

‘--gnu’ Set the global strictness to ‘gnu’. For more information, see Chapter 21 [Gnits], page 73. This is the default strictness.

‘--help’ Print a summary of the command line options and exit.

‘-i’
‘--ignore-deps’

This disables the dependency tracking feature in generated Makefiles; see Section 8.15 [Dependencies], page 48.

‘--include-deps’

This enables the dependency tracking feature. This feature is enabled by default. This option is provided for historical reasons only and probably should not be used.

‘--no-force’

Ordinarily automake creates all Makefile.ins mentioned in configure.ac. This option causes it to only update those Makefile.ins which are out of date with respect to one of their dependents.

‘-o dir’
‘--output-dir=dir’

Put the generated Makefile.in in the directory dir. Ordinarily each Makefile.in is created in the directory of the corresponding Makefile.am. This option is deprecated and will be removed in a future release.

‘-v’
‘--verbose’

Cause Automake to print information about which files are being read or created.

‘--version’

Print the version number of Automake and exit.

‘-W CATEGORY’
‘--warnings=category’

Output warnings falling in category. category can be one of:

‘gnu’ warnings related to the GNU Coding Standards (see The GNU Coding Standards).

‘obsolete’

obsolete features or constructions

‘override’

user redefinitions of Automake rules or variables

‘portability’

portability issues (e.g., use of Make features which are known not portable)

‘syntax’ weird syntax, unused variables, typos

‘unsupported’

unsupported or incomplete features

‘all’ all the warnings

‘none’ turn off all the warnings ‘error’ treat warnings as errors

A category can be turned off by prefixing its name with ‘no-’. For instance ‘-Wno-syntax’ will hide the warnings about unused variables.

The categories output by default are ‘syntax’ and ‘unsupported’. Additionally, ‘gnu’ is enabled in ‘--gnu’ and ‘--gnits’ strictness.

‘portability’ warnings are currently disabled by default, but they will be enabled in ‘--gnu’ and ‘--gnits’ strictness in a future release.

The environment variable ‘WARNINGS’ can contain a comma separated list of categories to enable. It will be taken into account before the command-line switches, this way ‘-Wnone’ will also ignore any warning category enabled by ‘WARNINGS’. This variable is also used by other tools like autoconf; unknown categories are ignored for this reason.

5 Scanning configure.ac

Automake scans the package’s configure.ac to determine certain information about the package. Some autoconf macros are required and some variables must be defined in configure.ac. Automake will also use information from configure.ac to further tailor its output.

Automake also supplies some Autoconf macros to make the maintenance easier. These macros can automatically be put into your aclocal.m4 using the aclocal program.

5.1 Configuration requirements

The one real requirement of Automake is that your configure.ac call AM_INIT_AUTOMAKE. This macro does several things which are required for proper Automake operation (see Section 5.6 [Macros], page 18).

Here are the other macros which Automake requires but which are not run by AM_INIT_

AUTOMAKE:
AC_CONFIG_FILES
AC_OUTPUT

Automake uses these to determine which files to create (see Section “Creating Output Files” in The Autoconf Manual). A listed file is considered to be an Automake generated Makefile if there exists a file with the same name and the

.am extension appended. Typically, AC_CONFIG_FILES([foo/Makefile]) will

cause Automake to generate foo/Makefile.in if foo/Makefile.am exists.

When using AC_CONFIG_FILES with multiple input files, as in AC_CONFIG_

FILES([Makefile:top.in:Makefile.in:bot.in]), Automake will generate

the first .in input file for which a .am file exists. If no such file exists the output file is not considered to be Automake generated.

Files created by AC_CONFIG_FILES are removed by make distclean.

5.2 Other things Automake recognizes

Every time Automake is run it calls Autoconf to trace configure.ac. This way it can recognize the use of certain macros and tailor the generated Makefile.in appropriately. Currently recognized macros and their effects are:

AC_CONFIG_HEADERS

Automake will generate rules to rebuild these headers. Older versions of Automake required the use of AM_CONFIG_HEADER (see Section 5.6 [Macros], page 18); this is no longer the case today.

AC_CONFIG_LINKS

Automake will generate rules to remove configure generated links on make distclean and to distribute named source files as part of make dist.

AC_CONFIG_AUX_DIR

Automake will look for various helper scripts, such as install-sh, in the directory named in this macro invocation. (The full list of scripts is:

config.guess, config.sub, depcomp, elisp-comp, compile, install-sh,
ltmain.sh, mdate-sh, missing, mkinstalldirs, py-compile, texinfo.tex,

and ylwrap.) Not all scripts are always searched for; some scripts will only be sought if the generated Makefile.in requires them. If AC_CONFIG_AUX_DIR is not given, the scripts are looked for in their ‘standard’ locations. For mdate-sh, texinfo.tex, and ylwrap, the standard location is the source directory corresponding to the current Makefile.am. For the rest, the standard location is the first one of ., .., or ../.. (relative to the top source directory) that provides any one of the helper scripts. See Section “Finding ‘configure’ Input” in The Autoconf Manual. Required files from AC_CONFIG_AUX_DIR are automatically distributed, even if there is no Makefile.am in this directory.

AC_CANONICAL_HOST

Automake will ensure that config.guess and config.sub exist. Also, the Makefile variables ‘host_alias’ and ‘host_triplet’ are introduced. See Section “Getting the Canonical System Type” in The Autoconf Manual.

AC_CANONICAL_SYSTEM

This is similar to AC_CANONICAL_HOST, but also defines the Makefile variables ‘build_alias’ and ‘target_alias’. See Section “Getting the Canonical System Type” in The Autoconf Manual.

AC_LIBSOURCE
AC_LIBSOURCES
AC_LIBOBJ

Automake will automatically distribute any file listed in AC_LIBSOURCE or AC_

LIBSOURCES.

Note that the AC_LIBOBJ macro calls AC_LIBSOURCE. So if an Autoconf macro is documented to call AC_LIBOBJ([file]), then file.c will be distributed automatically by Automake. This encompasses many macros like AC_FUNC_

ALLOCA, AC_FUNC_MEMCMP, AC_REPLACE_FUNCS, and others.

By the way, direct assignments to LIBOBJS are no longer supported. You should always use AC_LIBOBJ for this purpose. See Section “AC_LIBOBJ vs. LIBOBJS” in The Autoconf Manual.

AC_PROG_RANLIB

This is required if any libraries are built in the package. See Section “Particular Program Checks” in The Autoconf Manual.

AC_PROG_CXX

This is required if any C++ source is included. See Section “Particular Program Checks” in The Autoconf Manual.

AC_PROG_F77

This is required if any Fortran 77 source is included. This macro is distributed with Autoconf version 2.13 and later. See Section “Particular Program Checks” in The Autoconf Manual.

AC_F77_LIBRARY_LDFLAGS

This is required for programs and shared libraries that are a mixture of languages that include Fortran 77 (see Section 8.11.3 [Mixing Fortran 77 With C and C++], page 44). See Section 5.6 [Autoconf macros supplied with Automake], page 18.

AC_PROG_LIBTOOL

Automake will turn on processing for libtool (see Section “Introduction” in The Libtool Manual).

AC_PROG_YACC

If a Yacc source file is seen, then you must either use this macro or define the variable ‘YACC’ in configure.ac. The former is preferred (see Section “Particular Program Checks” in The Autoconf Manual).

AC_PROG_LEX

If a Lex source file is seen, then this macro must be used. See Section “Particular Program Checks” in The Autoconf Manual.

AC_SUBST The first argument is automatically defined as a variable in each generated Makefile.in. See Section “Setting Output Variables” in The Autoconf Manual. If the Autoconf manual says that a macro calls AC_SUBST for var, or defines the output variable var then var will be defined in each Makefile.in generated by Automake. E.g. AC_PATH_XTRA defines X_CFLAGS and X_LIBS, so you can use these variables in any Makefile.am if AC_PATH_XTRA is called.

AM_C_PROTOTYPES

This is required when using automatic de-ANSI-fication; see Section 8.14 [ANSI], page 47.

AM_GNU_GETTEXT

This macro is required for packages which use GNU gettext (see Section 10.2 [gettext], page 55). It is distributed with gettext. If Automake sees this macro it ensures that the package meets some of gettext’s requirements.

AM_MAINTAINER_MODE
          This macro adds a ‘--enable-maintainer-mode’ option to configure. If this is

used, automake will cause ‘maintainer-only’ rules to be turned off by default in the generated Makefile.ins. This macro defines the ‘MAINTAINER_MODE’ conditional, which you can use in your own Makefile.am.

m4_include

Files included by configure.ac using this macro will be detected by Automake and automatically distributed. They will also appear as dependencies

          in Makefile rules.

m4_include is seldom used by configure.ac authors, but can appear in aclocal.m4 when aclocal detects that some required macros come from files local to your package (as opposed to macros installed in a system-wide directory, see Section 5.3 [Invoking aclocal], page 15).

5.3 Auto-generating aclocal.m4

Automake includes a number of Autoconf macros which can be used in your package (see Section 5.6 [Macros], page 18); some of them are actually required by Automake in certain situations. These macros must be defined in your aclocal.m4; otherwise they will not be

seen by autoconf.

The aclocal program will automatically generate aclocal.m4 files based on the contents of configure.ac. This provides a convenient way to get Automake-provided macros, without having to search around. The aclocal mechanism allows other packages to supply their own macros (see Section 5.7 [Extending aclocal], page 21). You can also use it to maintain your own set of custom macros (see Section 5.8 [Local Macros], page 22).

At startup, aclocal scans all the .m4 files it can find, looking for macro definitions (see Section 5.5 [Macro search path], page 16). Then it scans configure.ac. Any mention of one of the macros found in the first step causes that macro, and any macros it in turn requires, to be put into aclocal.m4.

Putting the file that contains the macro definition into aclocal.m4 is usually done by copying the entire text of this file, including unused macro definitions as well as both ‘#’ and ‘dnl’ comments. If you want to make a comment which will be completely ignored by aclocal, use ‘##’ as the comment leader.

When aclocal detects that the file containing the macro definition is in a subdirectory of your package, it will use m4_include instead of copying it; this makes the package smaller and eases dependency tracking. This only works if the subdirectory containing the macro was specified as a relative search path with aclocal’s -I argument. (see Section 5.8 [Local Macros], page 22, for an example.) Any macro which is found in a system-wide directory, or via an absolute search path will be copied.

The contents of acinclude.m4, if it exists, are also automatically included in aclocal.m4. We recommend against using acinclude.m4 in new packages (see Section 5.8 [Local Macros], page 22).

While computing aclocal.m4, aclocal runs autom4te (see Section “Using Autom4te” in The Autoconf Manual) in order to trace the macros which are really used, and omit from aclocal.m4 all macros which are mentioned but otherwise unexpanded (this can happen when a macro is called conditionally). autom4te is expected to be in the PATH, just as autoconf. Its location can be overridden using the AUTOM4TE environment variable.

5.4 aclocal options

aclocal accepts the following options:

--acdir=dir

Look for the macro files in dir instead of the installation directory. This is typically used for debugging.

--help Print a summary of the command line options and exit.

-I dir Add the directory dir to the list of directories searched for .m4 files.

--force Always overwrite the output file. The default is to overwrite the output file only when really needed, i.e., when its contents changes or if one of its dependencies is younger.

--output=file

Cause the output to be put into file instead of aclocal.m4.

--print-ac-dir

Prints the name of the directory which aclocal will search to find third-party .m4 files. When this option is given, normal processing is suppressed. This option can be used by a package to determine where to install a macro file.

--verbose

Print the names of the files it examines.

--version

Print the version number of Automake and exit.

5.5 Macro search path

By default, aclocal searches for .m4 files in the following directories, in this order:

acdir-APIVERSION

This is where the .m4 macros distributed with automake itself are stored. APIVERSION depends on the automake release used; for automake 1.6.x, APIVERSION = 1.6.

acdir This directory is intended for third party .m4 files, and is configured when automake itself is built. This is @datadir@/aclocal/, which typically expands to ${prefix}/share/aclocal/. To find the compiled-in value of acdir, use the --print-ac-dir option (see Section 5.4 [aclocal options], page 16).

As an example, suppose that automake-1.6.2 was configured with --prefix=/usr/local. Then, the search path would be:

 1. /usr/local/share/aclocal-1.6/
 2. /usr/local/share/aclocal/

As explained in (see Section 5.4 [aclocal options], page 16), there are several options that can be used to change or extend this search path.

5.5.1 Modifying the macro search path: --acdir

The most obvious option to modify the search path is --acdir=dir, which changes default directory and drops the APIVERSION directory. For example, if one specifies --acdir=/opt/private/, then the search path becomes:

 1. /opt/private/

Note that this option, --acdir, is intended for use by the internal automake test suite only; it is not ordinarily needed by end-users.

5.5.2 Modifying the macro search path: -I dir

Any extra directories specified using -I options (see Section 5.4 [aclocal options], page 16) are prepended to this search list. Thus, aclocal -I /foo -I /bar results in the following search path:

 1. /foo
 2. /bar

3. acdir-APIVERSION

4. acdir

5.5.3 Modifying the macro search path: dirlist

There is a third mechanism for customizing the search path. If a dirlist file exists in acdir, then that file is assumed to contain a list of directories, one per line, to be added to the search list. These directories are searched after all other directories.

For example, suppose acdir/dirlist contains the following:

     /test1
     /test2

and that aclocal was called with the -I /foo -I /bar options. Then, the search path would be

 1. /foo
 2. /bar

3. acdir-APIVERSION

4. acdir

 5. /test1
 6. /test2

If the --acdir=dir option is used, then aclocal will search for the dirlist file in dir. In the --acdir=/opt/private/ example above, aclocal would look for /opt/private/dirlist. Again, however, the --acdir option is intended for use by the internal automake test suite only; --acdir is not ordinarily needed by end-users.

dirlist is useful in the following situation: suppose that automake version 1.6.2 is installed with $prefix=/usr by the system vendor. Thus, the default search directories are

 1. /usr/share/aclocal-1.6/
 2. /usr/share/aclocal/

However, suppose further that many packages have been manually installed on the system, with $prefix=/usr/local, as is typical. In that case, many of these “extra” .m4 files are

in /usr/local/share/aclocal. The only way to force /usr/bin/aclocal to find these

“extra” .m4 files is to always call aclocal -I /usr/local/share/aclocal. This is inconvenient. With dirlist, one may create the file

   /usr/share/aclocal/dirlist

which contains only the single line

   /usr/local/share/aclocal

Now, the “default” search path on the affected system is

 1. /usr/share/aclocal-1.6/
 2. /usr/share/aclocal/
 3. /usr/local/share/aclocal/

without the need for -I options; -I options can be reserved for project-specific needs (my-source-dir/m4/), rather than using it to work around local system-dependent tool installation directories.

Similarly, dirlist can be handy if you have installed a local copy Automake on your account and want aclocal to look for macros installed at other places on the system.

5.6 Autoconf macros supplied with Automake

Automake ships with several Autoconf macros that you can use from your configure.ac. When you use one of them it will be included by aclocal in aclocal.m4.

5.6.1 Public macros

AM_CONFIG_HEADER

Automake will generate rules to automatically regenerate the config header. This obsolete macro is a synonym of AC_CONFIG_HEADERS today (see Section 5.2 [Optional], page 13).

AM_ENABLE_MULTILIB

This is used when a “multilib” library is being built. The first optional argument is the name of the Makefile being generated; it defaults to ‘Makefile’. The second option argument is used to find the top source directory; it defaults to the empty string (generally this should not be used unless you are familiar with the internals). See Section 18.3 [Multilibs], page 71.

AM_C_PROTOTYPES

Check to see if function prototypes are understood by the compiler. If so, define ‘PROTOTYPES’ and set the output variables ‘U’ and ‘ANSI2KNR’ to the empty string. Otherwise, set ‘U’ to ‘_’ and ‘ANSI2KNR’ to ‘./ansi2knr’. Automake uses these values to implement automatic de-ANSI-fication.

AM_HEADER_TIOCGWINSZ_NEEDS_SYS_IOCTL
          If the use of TIOCGWINSZ requires <sys/ioctl.h>, then define GWINSZ_IN_SYS_
          IOCTL. Otherwise TIOCGWINSZ can be found in <termios.h>.

AM_INIT_AUTOMAKE([OPTIONS])
AM_INIT_AUTOMAKE(PACKAGE, VERSION, [NO-DEFINE])

Runs many macros required for proper operation of the generated Makefiles.

This macro has two forms, the first of which is preferred. In this form, AM_ INIT_AUTOMAKE is called with a single argument — a space-separated list of Automake options which should be applied to every Makefile.am in the tree. The effect is as if each option were listed in AUTOMAKE_OPTIONS.

The second, deprecated, form of AM_INIT_AUTOMAKE has two required arguments: the package and the version number. This form is obsolete because the package and version can be obtained from Autoconf’s AC_INIT macro (which itself has an old and a new form).

If your configure.ac has:
     AC_INIT(src/foo.c)
     AM_INIT_AUTOMAKE(mumble, 1.5)

you can modernize it as follows:

     AC_INIT(mumble, 1.5)
     AC_CONFIG_SRCDIR(src/foo.c)
     AM_INIT_AUTOMAKE

Note that if you’re upgrading your configure.ac from an earlier version of Automake, it is not always correct to simply move the package and version arguments from AM_INIT_AUTOMAKE directly to AC_INIT, as in the example above. The first argument to AC_INIT should be the name of your package (e.g. ‘GNU Automake’), not the tarball name (e.g. ‘automake’) that you used to pass to AM_INIT_AUTOMAKE. Autoconf tries to derive a tarball name from the package name, which should work for most but not all package names. (If it doesn’t work for yours, you can use the four-argument form of AC_INIT — supported in Autoconf versions greater than 2.52g — to provide the tarball name explicitly).

By default this macro AC_DEFINE’s ‘PACKAGE’ and ‘VERSION’. This can be avoided by passing the ‘no-define’ option, as in:

     AM_INIT_AUTOMAKE([gnits 1.5 no-define dist-bzip2])

or by passing a third non-empty argument to the obsolete form.

AM_PATH_LISPDIR

Searches for the program emacs, and, if found, sets the output variable lispdir to the full path to Emacs’ site-lisp directory.

Note that this test assumes the emacs found to be a version that supports Emacs Lisp (such as gnu Emacs or XEmacs). Other emacsen can cause this test to hang (some, like old versions of MicroEmacs, start up in interactive mode, requiring ‘C-x C-c’ to exit, which is hardly obvious for a non-emacs user). In most cases, however, you should be able to use ‘C-c’ to kill the test. In order to avoid problems, you can set EMACS to “no” in the environment, or use the ‘--with-lispdir’ option to configure to explicitly set the correct path (if you’re sure you have an emacs that supports Emacs Lisp.

AM_PROG_AS

Use this macro when you have assembly code in your project. This will choose the assembler for you (by default the C compiler) and set CCAS, and will also set CCASFLAGS if required.

AM_PROG_CC_C_O

This is like AC_PROG_CC_C_O, but it generates its results in the manner required by automake. You must use this instead of AC_PROG_CC_C_O when you need this functionality.

AM_PROG_LEX

Like AC_PROG_LEX (see Section “Particular Program Checks” in The Autoconf Manual), but uses the missing script on systems that do not have lex. ‘HP-UX 10’ is one such system.

AM_PROG_GCJ

This macro finds the gcj program or causes an error. It sets ‘GCJ’ and ‘GCJFLAGS’. gcj is the Java front-end to the GNU Compiler Collection.

AM_SYS_POSIX_TERMIOS

Check to see if POSIX termios headers and functions are available on the system. If so, set the shell variable am_cv_sys_posix_termios to ‘yes’. If not, set the variable to ‘no’.

AM_WITH_DMALLOC
Add support for the dmalloc (ftp://ftp.letters.com/src/dmalloc/

dmalloc.tar.gz) package. If the user configures with ‘--with-dmalloc’, then

define WITH_DMALLOC and add ‘-ldmalloc’ to LIBS.
AM_WITH_REGEX

Adds ‘--with-regex’ to the configure command line. If specified (the default), then the ‘regex’ regular expression library is used, regex.o is put into ‘LIBOBJS’, and ‘WITH_REGEX’ is defined. If ‘--without-regex’ is given, then the ‘rx’ regular expression library is used, and rx.o is put into ‘LIBOBJS’.

5.6.2 Private macros

The following macros are private macros you should not call directly. They are called by the other public macros when appropriate. Do not rely on them, as they might be changed in a future version. Consider them as implementation details; or better, do not consider them at all: skip this section!

_AM_DEPENDENCIES
AM_SET_DEPDIR
AM_DEP_TRACK
AM_OUTPUT_DEPENDENCY_COMMANDS

These macros are used to implement Automake’s automatic dependency tracking scheme. They are called automatically by automake when required, and there should be no need to invoke them manually.

AM_MAKE_INCLUDE

This macro is used to discover how the user’s make handles include statements. This macro is automatically invoked when needed; there should be no need to invoke it manually.

AM_PROG_INSTALL_STRIP

This is used to find a version of install which can be used to strip a program at installation time. This macro is automatically included when required.

AM_SANITY_CHECK

This checks to make sure that a file created in the build directory is newer than a file in the source directory. This can fail on systems where the clock is set incorrectly. This macro is automatically run from AM_INIT_AUTOMAKE.

5.7 Writing your own aclocal macros

The aclocal program doesn’t have any built-in knowledge of any macros, so it is easy to extend it with your own macros.

This can be used by libraries which want to supply their own Autoconf macros for use by other programs. For instance the gettext library supplies a macro AM_GNU_GETTEXT which should be used by any package using gettext. When the library is installed, it installs this macro so that aclocal will find it.

A macro file’s name should end in .m4. Such files should be installed in $(datadir)/aclocal. This is as simple as writing:

     aclocaldir = $(datadir)/aclocal
     aclocal_DATA = mymacro.m4 myothermacro.m4

A file of macros should be a series of properly quoted AC_DEFUN’s (see Section “Macro Definitions” in The Autoconf Manual). The aclocal programs also understands AC_REQUIRE (see Section “Prerequisite Macros” in The Autoconf Manual), so it is safe to put each macro in a separate file. Each file should have no side effects but macro definitions. Especially, any call to AC_PREREQ should be done inside the defined macro, not at the beginning of the file.

Starting with Automake 1.8, aclocal will warn about all underquoted calls to AC_DEFUN. We realize this will annoy a lot of people, because aclocal was not so strict in the past and many third party macros are underquoted; and we have to apologize for this temporary inconvenience. The reason we have to be stricter is that a future implementation of aclocal (see Section 5.9 [Future of aclocal], page 23) will have to temporary include all these third party .m4 files, maybe several times, even those which are not actually needed. Doing so should alleviate many problem of the current implementation, however it requires a stricter style from the macro authors. Hopefully it is easy to revise the existing macros. For instance

     # bad style
     AC_PREREQ(2.57)
     AC_DEFUN(AX_FOOBAR,
     [AC_REQUIRE([AX_SOMETHING])dnl
     AX_FOO
     AX_BAR
     ])

should be rewritten as

     AC_DEFUN([AX_FOOBAR],
     [AC_PREREQ(2.57)dnl
     AC_REQUIRE([AX_SOMETHING])dnl
     AX_FOO
     AX_BAR
     ])

Wrapping the AC_PREREQ call inside the macro ensures that Autoconf 2.57 will not be required if AX_FOOBAR is not actually used. Most importantly, quoting the first argument of AC_DEFUN allows the macro to be redefined or included twice (otherwise this first argument would be expansed during the second definition).

If you have been directed here by the aclocal diagnostic but are not the maintainer of the implicated macro, you will want to contact the maintainer of that macro. Please make sure you have the last version of the macro and that the problem already hasn’t been reported before doing so: people tend to work faster when they aren’t flooded by mails.

Another situation where aclocal is commonly used is to manage macros which are used locally by the package, Section 5.8 [Local Macros], page 22.

5.8 Handling Local Macros

Feature tests offered by Autoconf do not cover all needs. People often have to supplement existing tests with their own macros, or with third-party macros.

There are two ways to organize custom macros in a package.

The first possibility (the historical practice) is to list all your macros in acinclude.m4. This file will be included in aclocal.m4 when you run aclocal, and its macro(s) will henceforth be visible to autoconf. However if it contains numerous macros, it will rapidly become difficult to maintain, and it will be almost impossible to share macros between packages.

The second possibility, which we do recommend, is to write each macro in its own file and gather all these files in a directory. This directory is usually called m4/. To build aclocal.m4, one should therefore instruct aclocal to scan m4/. From the command line, this is done with aclocal -I m4. The top-level Makefile.am should also be updated to define

      ACLOCAL_AMFLAGS = -I m4

ACLOCAL_AMFLAGS contains options to pass to aclocal when aclocal.m4 is to be rebuilt by make. This line is also used by autoreconf (see Section “Using autoreconf to Update configure Scripts” in The Autoconf Manual) to run aclocal with suitable options, or by autopoint (see Section “Invoking the autopoint Program” in GNU gettext tools) and gettextize (see Section “Invoking the gettextize Program” in GNU gettext tools) to locate the place where Gettext’s macros should be installed. So even if you do not really care about the rebuild rules, you should define ACLOCAL_AMFLAGS.

When aclocal -I m4 is run, it will build a aclocal.m4 that m4_includes any file from m4/ that defines a required macro. Macros not found locally will still be searched in systemwide directories, as explained in Section 5.5 [Macro search path], page 16.

Custom macros should be distributed for the same reason that configure.ac is: so that other people have all the sources of your package if they want to work on it. Actually, this distribution happens automatically because all m4_included files are distributed.

However there is no consensus on the distribution of third-party macros that your package may use. Many libraries install their own macro in the system-wide aclocal directory (see Section 5.7 [Extending aclocal], page 21). For instance Guile ships with a file called guile.m4 that contains the macro GUILE_FLAGS which can be used to define setup compiler and linker flags appropriate for using Guile. Using GUILE_FLAGS in configure.ac will cause aclocal to copy guile.m4 into aclocal.m4, but as guile.m4 is not part of the project, it will not be distributed. Technically, that means a user which needs to rebuild aclocal.m4 will have to install Guile first. This is probably OK, if Guile already is a requirement to build the package. However, if Guile is only an optional feature, or if your package might run on architectures where Guile cannot be installed, this requirement will hinder development. An easy solution is to copy such third-party macros in your local m4/ directory so they get distributed.

5.9 The Future of aclocal

aclocal is expected to disappear. This feature really should not be offered by Automake. Automake should focus on generating Makefiles; dealing with M4 macros really is Autoconf’s job. That some people install Automake just to use aclocal, but do not use automake otherwise is an indication of how that feature is misplaced.

The new implementation will probably be done slightly differently. For instance it could enforce the m4/-style layout discussed in Section 5.8 [Local Macros], page 22, and take care of copying (and even updating) third-party macros from /usr/share/aclocal/ into the local m4/ directory.

We have no idea when and how this will happen. This has been discussed several times in the past, but someone still has to commit itself to that non-trivial task.

From the user point of view, aclocal’s removal might turn out to be painful. There is a simple precaution that you may take to make that switch more seamless: never call aclocal yourself. Keep this guy under the exclusive control of autoreconf and Automake’s rebuild rules. Hopefully you won’t need to worry about things breaking, when aclocal disappears, because everything will have been taken care of. If otherwise you used to call aclocal directly yourself or from some script, you will quickly notice the change.

Many packages come with a script called bootstrap.sh or autogen.sh, that will just call

aclocal, libtoolize, gettextize or autopoint, autoconf, autoheader, and automake

in the right order. Actually this is precisely what autoreconf can do for you. If your package has such a bootstrap.sh or autogen.sh script, consider using autoreconf. That should simplify its logic a lot (less things to maintain, yum!), it’s even likely you will not need the script anymore, and more to the point you will not call aclocal directly anymore.

For the time being, third-party packages should continue to install public macros into /usr/share/aclocal/. If aclocal is replaced by another tool it might make sense to rename the directory, but supporting /usr/share/aclocal/ for backward compatibility should be really easy provided all macros are properly written (see Section 5.7 [Extending aclocal], page 21).

6 The top-level Makefile.am

6.1 Recursing subdirectories

In packages with subdirectories, the top level Makefile.am must tell Automake which subdirectories are to be built. This is done via the SUBDIRS variable.

The SUBDIRS variable holds a list of subdirectories in which building of various sorts can occur. The rules for many targets (e.g. all) in the generated Makefile will run commands both locally and in all specified subdirectories. Note that the directories listed in SUBDIRS are not required to contain Makefile.ams; only Makefiles (after configuration). This allows inclusion of libraries from packages which do not use Automake (such as gettext).

In packages that use subdirectories, the top-level Makefile.am is often very short. For instance, here is the Makefile.am from the GNU Hello distribution:

     EXTRA_DIST = BUGS ChangeLog.O README-alpha
     SUBDIRS = doc intl po src tests

When Automake invokes make in a subdirectory, it uses the value of the MAKE variable. It passes the value of the variable AM_MAKEFLAGS to the make invocation; this can be set in Makefile.am if there are flags you must always pass to make.

The directories mentioned in SUBDIRS must be direct children of the current directory. For instance, you cannot put ‘src/subdir’ into SUBDIRS. Instead you should put SUBDIRS = subdir into src/Makefile.am. Automake can be used to construct packages of arbitrary depth this way.

By default, Automake generates Makefiles which work depth-first (‘postfix’). However, it is possible to change this ordering. You can do this by putting ‘.’ into SUBDIRS. For instance, putting ‘.’ first will cause a ‘prefix’ ordering of directories. All ‘clean’ rules are run in reverse order of build rules.

6.2 Conditional subdirectories

It is possible to define the SUBDIRS variable conditionally if, like in the case of GNU Inetutils, you want to only build a subset of the entire package.

To illustrate how this works, let’s assume we have two directories src/ and opt/. src/ should always be built, but we want to decide in ./configure whether opt/ will be built or not. (For this example we will assume that opt/ should be built when the variable

$want_opt was set to yes.)

Running make should thus recurse into src/ always, and then maybe in opt/.

However make dist should always recurse into both src/ and opt/. Because opt/ should be distributed even if it is not needed in the current configuration. This means opt/Makefile should be created unconditionally.^{3}

There are two ways to setup a project like this. You can use Automake conditionals (see Chapter 20 [Conditionals], page 72) or use Autoconf AC_SUBST variables (see Section “Setting Output Variables” in The Autoconf Manual). Using Automake conditionals is the preferred solution.

3 Don’t try seeking a solution where opt/Makefile is created conditionally, this is a lot trickier than the

solutions presented here.

6.2.1 Conditional subdirectories with AM_CONDITIONAL

configure should output the Makefile for each directory and define a condition into which opt/ should be built.

...
AM_CONDITIONAL([COND_OPT], [test "$want_opt" = yes])
AC_CONFIG_FILES([Makefile src/Makefile opt/Makefile])
...

Then SUBDIRS can be defined in the top-level Makefile.am as follows.

if COND_OPT
  MAYBE_OPT = opt
endif
SUBDIRS = src $(MAYBE_OPT)

As you can see, running make will rightly recurse into src/ and maybe opt/. As you can’t see, running make dist will recurse into both src/ and opt/ directories

because make dist, unlike make all, doesn’t use the SUBDIRS variable. It uses the DIST_ SUBDIRS variable.

In this case Automake will define DIST_SUBDIRS = src opt automatically because it knows that MAYBE_OPT can contain opt in some condition.

6.2.2 Conditional subdirectories with AC_SUBST

Another idea is to define MAYBE_OPT from ./configure using AC_SUBST:

...
if test "$want_opt" = yes; then
  MAYBE_OPT=opt
else
  MAYBE_OPT=
fi
AC_SUBST([MAYBE_OPT])
AC_CONFIG_FILES([Makefile src/Makefile opt/Makefile])
...

In this case the top-level Makefile.am should look as follows.

SUBDIRS = src $(MAYBE_OPT)
DIST_SUBDIRS = src opt

The drawback is that since Automake cannot guess what the possible values of MAYBE_ OPT are, it is necessary to define DIST_SUBDIRS.

6.2.3 How DIST_SUBDIRS is used

As shown in the above examples, DIST_SUBDIRS is used in rules that need to recurse in all directories, even those which have been conditionally left out of the build.

Precisely, DIST_SUBDIRS is used by make dist, make distclean, and make

maintainer-clean. All other recursive rules use SUBDIRS.

Automake will define DIST_SUBDIRS automatically from the possibles values of SUBDIRS in all conditions.

If SUBDIRS contains AC_SUBST variables, DIST_SUBDIRS will not be defined correctly because Automake doesn’t know the possible values of these variables. In this case DIST_ SUBDIRS needs to be defined manually.

7 An Alternative Approach to Subdirectories

If you’ve ever read Peter Miller’s excellent paper, Recursive Make Considered Harmful

(http://www.pcug.org.au/~millerp/rmch/recu-make-cons-harm.html), the preceding

section on the use of subdirectories will probably come as unwelcome advice. For those who haven’t read the paper, Miller’s main thesis is that recursive make invocations are both slow and error-prone.

Automake provides sufficient cross-directory support^{4} to enable you to write a single Makefile.am for a complex multi-directory package.

By default an installable file specified in a subdirectory will have its directory name stripped before installation. For instance, in this example, the header file will be installed

as $(includedir)/stdio.h:
     include_HEADERS = inc/stdio.h

However, the ‘nobase_’ prefix can be used to circumvent this path stripping. In this example, the header file will be installed as $(includedir)/sys/types.h:

     nobase_include_HEADERS = sys/types.h

‘nobase_’ should be specified first when used in conjunction with either ‘dist_’ or ‘nodist_’ (see Chapter 14 [Dist], page 62). For instance:

     nobase_dist_pkgdata_DATA = images/vortex.pgm

8 Building Programs and Libraries

A large part of Automake’s functionality is dedicated to making it easy to build programs and libraries.

8.1 Building a program

In order to build a program, you need to tell Automake which sources are part of it, and which libraries it should be linked with.

This section also covers conditional compilation of sources or programs. Most of the comments about these also apply to libraries (see Section 8.2 [A Library], page 30) and libtool libraries (see Section 8.3 [A Shared Library], page 30).

4 We believe. This work is new and there are probably warts. See Chapter 1 [Introduction], page 1, for

information on reporting bugs.

8.1.1 Defining program sources

In a directory containing source that gets built into a program (as opposed to a library or a script), the ‘PROGRAMS’ primary is used. Programs can be installed in bindir, sbindir, libexecdir, pkglibdir, or not at all (‘noinst’). They can also be built only for make check, in which case the prefix is ‘check’.

For instance:

     bin_PROGRAMS = hello

In this simple case, the resulting Makefile.in will contain code to generate a program named hello.

Associated with each program are several assisting variables which are named after the program. These variables are all optional, and have reasonable defaults. Each variable, its use, and default is spelled out below; we use the “hello” example throughout.

The variable hello_SOURCES is used to specify which source files get built into an executable:

     hello_SOURCES = hello.c version.c getopt.c getopt1.c getopt.h system.h

This causes each mentioned ‘.c’ file to be compiled into the corresponding ‘.o’. Then all are linked to produce hello.

If ‘hello_SOURCES’ is not specified, then it defaults to the single file hello.c (see Section 8.5 [Default SOURCES], page 38).

Multiple programs can be built in a single directory. Multiple programs can share a single source file, which must be listed in each ‘_SOURCES’ definition.

Header files listed in a ‘_SOURCES’ definition will be included in the distribution but otherwise ignored. In case it isn’t obvious, you should not include the header file generated by configure in a ‘_SOURCES’ variable; this file should not be distributed. Lex (‘.l’) and Yacc (‘.y’) files can also be listed; see Section 8.8 [Yacc and Lex], page 40.

8.1.2 Linking the program

If you need to link against libraries that are not found by configure, you can use LDADD to do so. This variable is used to specify additional objects or libraries to link with; it is inappropriate for specifying specific linker flags, you should use AM_LDFLAGS for this purpose.

Sometimes, multiple programs are built in one directory but do not share the same linktime requirements. In this case, you can use the ‘prog_LDADD’ variable (where prog is the name of the program as it appears in some ‘_PROGRAMS’ variable, and usually written in lowercase) to override the global LDADD. If this variable exists for a given program, then that program is not linked using LDADD.

For instance, in GNU cpio, pax, cpio and mt are linked against the library libcpio.a. However, rmt is built in the same directory, and has no such link requirement. Also, mt and rmt are only built on certain architectures. Here is what cpio’s src/Makefile.am looks like (abridged):

     bin_PROGRAMS = cpio pax $(MT)
     libexec_PROGRAMS = $(RMT)
     EXTRA_PROGRAMS = mt rmt

     LDADD = ../lib/libcpio.a $(INTLLIBS)
     rmt_LDADD =
     cpio_SOURCES = ...
     pax_SOURCES = ...
     mt_SOURCES = ...
     rmt_SOURCES = ...

‘prog_LDADD’ is inappropriate for passing program-specific linker flags (except for ‘-l’, ‘-L’, ‘-dlopen’ and ‘-dlpreopen’). So, use the ‘prog_LDFLAGS’ variable for this purpose.

It is also occasionally useful to have a program depend on some other target which is not actually part of that program. This can be done using the ‘prog_DEPENDENCIES’ variable. Each program depends on the contents of such a variable, but no further interpretation is done.

If ‘prog_DEPENDENCIES’ is not supplied, it is computed by Automake. The automatically-assigned value is the contents of ‘prog_LDADD’, with most configure substitutions, ‘-l’, ‘-L’, ‘-dlopen’ and ‘-dlpreopen’ options removed. The configure substitutions that are left in are only ‘$(LIBOBJS)’ and ‘$(ALLOCA)’; these are left because it is known that they will not cause an invalid value for ‘prog_DEPENDENCIES’ to be generated.

8.1.3 Conditional compilation of sources

You can’t put a configure substitution (e.g., ‘@FOO@’ or ‘$(FOO)’ where FOO is defined via AC_SUBST) into a ‘_SOURCES’ variable. The reason for this is a bit hard to explain, but suffice to say that it simply won’t work. Automake will give an error if you try to do this.

Fortunately there are two other ways to achieve the same result. One is to use configure substitutions in _LDADD variables, the other is to use an Automake conditional.

8.1.3.1 Conditional compilation using _LDADD substitutions

Automake must know all the source files that could possibly go into a program, even if not all the files are built in every circumstance. Any files which are only conditionally built should be listed in the appropriate ‘EXTRA_’ variable. For instance, if hello-linux.c or hello-generic.c were conditionally included in hello, the Makefile.am would contain:

     bin_PROGRAMS = hello
     hello_SOURCES = hello-common.c
     EXTRA_hello_SOURCES = hello-linux.c hello-generic.c
     hello_LDADD = $(HELLO_SYSTEM)
     hello_DEPENDENCIES = $(HELLO_SYSTEM)

You can then setup the $(HELLO_SYSTEM) substitution from configure.ac:

     ...
     case $host in
       *linux*) HELLO_SYSTEM='hello-linux.$(OBJEXT)' ;;
       *)       HELLO_SYSTEM='hello-generic.$(OBJEXT)' ;;
     esac
     AC_SUBST([HELLO_SYSTEM])

     ...
   In this case, HELLO_SYSTEM should be replaced by hello-linux.o or hello-generic.o,

and added to hello_DEPENDENCIES and hello_LDADD in order to be built and linked in.

8.1.3.2 Conditional compilation using Automake conditionals

An often simpler way to compile source files conditionally is to use Automake conditionals. For instance, you could use this Makefile.am construct to build the same hello example:

     bin_PROGRAMS = hello
     if LINUX
     hello_SOURCES = hello-linux.c hello-common.c
     else
     hello_SOURCES = hello-generic.c hello-common.c
     endif

In this case, your configure.ac should setup the LINUX conditional using AM_CONDITIONAL (see Chapter 20 [Conditionals], page 72).

When using conditionals like this you don’t need to use the ‘EXTRA_’ variable, because Automake will examine the contents of each variable to construct the complete list of source files.

If your program uses a lot of files, you will probably prefer a conditional +=.

     bin_PROGRAMS = hello
     hello_SOURCES = hello-common.c
     if LINUX
     hello_SOURCES += hello-linux.c
     else
     hello_SOURCES += hello-generic.c
     endif

8.1.4 Conditional compilation of programs

Sometimes it is useful to determine the programs that are to be built at configure time. For instance, GNU cpio only builds mt and rmt under special circumstances. The means to achieve conditional compilation of programs are the same you can use to compile source files conditionally: substitutions or conditionals.

8.1.4.1 Conditional programs using configure substitutions

In this case, you must notify Automake of all the programs that can possibly be built, but at the same time cause the generated Makefile.in to use the programs specified by configure. This is done by having configure substitute values into each ‘_PROGRAMS’ definition, while listing all optionally built programs in EXTRA_PROGRAMS.

     bin_PROGRAMS = cpio pax $(MT)
     libexec_PROGRAMS = $(RMT)
     EXTRA_PROGRAMS = mt rmt

As explained in Section 8.16 [EXEEXT], page 48, Automake will rewrite bin_PROGRAMS,

libexec_PROGRAMS, and EXTRA_PROGRAMS, appending $(EXEEXT) to each binary. Ob-

viously it cannot rewrite values obtained at run-time through configure substitutions, therefore you should take care of appending $(EXEEXT) yourself, as in AC_SUBST([MT],

['mt${EXEEXT}']).

8.1.4.2 Conditional programs using Automake conditionals

You can also use Automake conditionals (see Chapter 20 [Conditionals], page 72) to select programs to be built. In this case you don’t have to worry about $(EXEEXT) or EXTRA_

PROGRAMS.
     bin_PROGRAMS = cpio pax
     if WANT_MT
       bin_PROGRAMS += mt
     endif
     if WANT_RMT
       libexec_PROGRAMS = rmt
     endif

8.2 Building a library

Building a library is much like building a program. In this case, the name of the primary is ‘LIBRARIES’. Libraries can be installed in libdir or pkglibdir.

See Section 8.3 [A Shared Library], page 30, for information on how to build shared libraries using libtool and the ‘LTLIBRARIES’ primary.

Each ‘_LIBRARIES’ variable is a list of the libraries to be built. For instance to create a library named libcpio.a, but not install it, you would write:

     noinst_LIBRARIES = libcpio.a

The sources that go into a library are determined exactly as they are for programs, via the ‘_SOURCES’ variables. Note that the library name is canonicalized (see Section 2.4 [Canonicalization], page 4), so the ‘_SOURCES’ variable corresponding to liblob.a is

‘liblob_a_SOURCES’, not ‘liblob.a_SOURCES’.

Extra objects can be added to a library using the ‘library_LIBADD’ variable. This should be used for objects determined by configure. Again from cpio:

     libcpio_a_LIBADD = $(LIBOBJS) $(ALLOCA)

In addition, sources for extra objects that will not exist until configure-time must be added to the BUILT_SOURCES variable (see Section 9.4 [Sources], page 51).

Building a static library is done by compiling all object files, then by invoking $(AR) $(ARFLAGS) followed by the name of the library and the list of objects, and finally by calling $(RANLIB) on that library. You should call AC_PROG_RANLIB from your configure.ac to define RANLIB (Automake will complain otherwise). AR and ARFLAGS default to ar and cru respectively; you can override these two variables my setting them in your Makefile.am, by AC_SUBSTing them from your configure.ac, or by defining a per-library maude_AR variable (see Section 8.4 [Program and Library Variables], page 35).

8.3 Building a Shared Library

Building shared libraries portably is a relatively complex matter. For this reason, GNU Libtool (see Section “Introduction” in The Libtool Manual) was created to help build shared libraries in a platform-independent way.

8.3.1 The Libtool Concept

Libtool abstracts shared and static libraries into a unified concept henceforth called libtool libraries. Libtool libraries are files using the .la suffix, and can designate a static library, a shared library, or maybe both. Their exact nature cannot be determined until ./configure is run: not all platforms support all kinds of libraries, and users can explicitly select which libraries should be built. (However the package’s maintainers can tune the default, See Section “The AC_PROG_LIBTOOL macro” in The Libtool Manual.)

Because object files for shared and static libraries must be compiled differently, libtool is also used during compilation. Object files built by libtool are called libtool objects: these are files using the .lo suffix. Libtool libraries are built from these libtool objects.

You should not assume anything about the structure of .la or .lo files and how libtool constructs them: this is libtool’s concern, and the last thing one wants is to learn about libtool’s guts. However the existence of these files matters, because they are used as targets and dependencies in Makefiles rules when building libtool libraries. There are situations where you may have to refer to these, for instance when expressing dependencies for building source files conditionally (see Section 8.3.4 [Conditional Libtool Sources], page 32).

People considering writing a plug-in system, with dynamically loaded modules, should look into libltdl: libtool’s dlopening library (see Section “Using libltdl” in The Libtool Manual). This offers a portable dlopening facility to load libtool libraries dynamically, and can also achieve static linking where unavoidable.

Before we discuss how to use libtool with Automake in details, it should be noted that the libtool manual also has a section about how to use Automake with libtool (see Section “Using Automake with Libtool” in The Libtool Manual).

8.3.2 Building Libtool Libraries

Automake uses libtool to build libraries declared with the ‘LTLIBRARIES’ primary. Each ‘_LTLIBRARIES’ variable is a list of libtool libraries to build. For instance, to create a libtool library named libgettext.la, and install it in ‘libdir’, write:

     lib_LTLIBRARIES = libgettext.la
     libgettext_la_SOURCES = gettext.c gettext.h ...

Automake predefines the variable ‘pkglibdir’, so you can use pkglib_LTLIBRARIES to

install libraries in $(libdir)/@PACKAGE@/.

8.3.3 Building Libtool Libraries Conditionally

Like conditional programs (see Section 8.1.4 [Conditional Programs], page 29), there are two main ways to build conditional libraries: using Automake conditionals or using Autoconf AC_SUBSTitutions.

The important implementation detail you have to be aware of is that the place where a library will be installed matters to libtool: it needs to be indicated at link-time using the -rpath option.

For libraries whose destination directory is known when Automake runs, Automake will automatically supply the appropriate ‘-rpath’ option to libtool. This is the case for libraries listed explicitly in some installable _LTLIBRARIES variables such as lib_LTLIBRARIES.

However, for libraries determined at configure time (and thus mentioned in EXTRA_ LTLIBRARIES), Automake does not know the final installation directory. For such libraries you must add the ‘-rpath’ option to the appropriate ‘_LDFLAGS’ variable by hand.

The examples below illustrate the differences between these two methods. Here is an example where $(WANTEDLIBS) is an AC_SUBSTed variable set at ./configure-

time to either libfoo.la, libbar.la, both, or none. Although $(WANTEDLIBS) appears in the lib_LTLIBRARIES, Automake cannot guess it relates to libfoo.la or libbar.la by the time it creates the link rule for these two libraries. Therefore the -rpath argument must be explicitly supplied.

EXTRA_LTLIBRARIES = libfoo.la libbar.la
lib_LTLIBRARIES = $(WANTEDLIBS)
libfoo_la_SOURCES = foo.c ...
libfoo_la_LDFLAGS = -rpath '$(libdir)'
libbar_la_SOURCES = bar.c ...
libbar_la_LDFLAGS = -rpath '$(libdir)'

Here is how the same Makefile.am would look using Automake conditionals named WANT_LIBFOO and WANT_LIBBAR. Now Automake is able to compute the -rpath setting itself, because it’s clear that both libraries will end up in $(libdir) if they are installed.

lib_LTLIBRARIES =
if WANT_LIBFOO
lib_LTLIBRARIES += libfoo.la
endif
if WANT_LIBBAR
lib_LTLIBRARIES += libbar.la
endif
libfoo_la_SOURCES = foo.c ...
libbar_la_SOURCES = bar.c ...

8.3.4 Libtool Libraries with Conditional Sources

Conditional compilation of sources in a library can be achieved in the same way as conditional compilation of sources in a program (see Section 8.1.3 [Conditional Sources], page 28). The only difference is that _LIBADD should be used instead of _LDADD and that it should mention libtool objects (.lo files).

So, to mimic the hello example from Section 8.1.3 [Conditional Sources], page 28, we

could build a libhello.la library using either hello-linux.c or hello-generic.c with

the following Makefile.am.

lib_LTLIBRARIES = libhello.la
libhello_la_SOURCES = hello-common.c
EXTRA_libhello_la_SOURCES = hello-linux.c hello-generic.c
libhello_la_LIBADD = $(HELLO_SYSTEM)
libhello_la_DEPENDENCIES = $(HELLO_SYSTEM)
And make sure $(HELLO_SYSTEM) is set to either hello-linux.lo or hello-generic.lo
in ./configure.

Or we could simply use an Automake conditional as follows.

lib_LTLIBRARIES = libhello.la

libhello_la_SOURCES = hello-common.c
if LINUX
libhello_la_SOURCES += hello-linux.c
else
libhello_la_SOURCES += hello-generic.c
endif

8.3.5 Libtool Convenience Libraries

Sometimes you want to build libtool libraries which should not be installed. These are called libtool convenience libraries and are typically used to encapsulate many sublibraries, later gathered into one big installed library.

Libtool convenience libraries are declared by noinst_LTLIBRARIES, check_ LTLIBRARIES, or even EXTRA_LTLIBRARIES. Unlike installed libtool libraries they do not need an -rpath flag at link time (actually this is the only difference).

Convenience libraries listed in noinst_LTLIBRARIES are always built. Those listed in check_LTLIBRARIES are built only upon make check. Finally, libraries listed in EXTRA_ LTLIBRARIES are never built explicitly: Automake outputs rules to build them, but if the library does not appear as a Makefile dependency anywhere it won’t be built (this is why EXTRA_LTLIBRARIES is used for conditional compilation).

Here is a sample setup merging libtool convenience libraries from subdirectories into one main libtop.la library.

# -- Top-level Makefile.am --
SUBDIRS = sub1 sub2 ...
lib_LTLIBRARIES = libtop.la
libtop_la_SOURCES =
libtop_la_LIBADD = \
  sub1/libsub1.la \
  sub2/libsub2.la \
  ...
# -- sub1/Makefile.am --
noinst_LTLIBRARIES = libsub1.la
libsub1_la_SOURCES = ...
# -- sub2/Makefile.am --
# showing nested convenience libraries
SUBDIRS = sub2.1 sub2.2 ...
noinst_LTLIBRARIES = libsub2.la
libsub2_la_SOURCES =
libsub2_la_LIBADD = \
  sub21/libsub21.la \
  sub22/libsub22.la \
  ...

8.3.6 Libtool Modules

These are libtool libraries meant to be dlopened. They are indicated to libtool by passing -module at link-time.

     pkglib_LTLIBRARIES = mymodule.la
     mymodule_la_SOURCES = doit.c
     mymodule_LDFLAGS = -module

Ordinarily, Automake requires that a Library’s name starts with ‘lib’. However, when building a dynamically loadable module you might wish to use a "nonstandard" name.

If ‘mymodule_la_SOURCES’ is not specified, then it defaults to the single file mymodule.c (see Section 8.5 [Default SOURCES], page 38).

8.3.7 LIBADD and LDFLAGS

As shown in previous sections, the ‘library_LIBADD’ variable should be used to list extra libtool objects (.lo files) or libtool libraries (.la) to add to library.

The ‘library_LDFLAGS’ variable is the place to list additional libtool flags, such as ‘-version-info’, ‘-static’, and a lot more. See See Section “Using libltdl” in The Libtool Manual.

8.3.8 LTLIBOBJS

Where an ordinary library might include $(LIBOBJS), a libtool library must use $(LTLIBOBJS). This is required because the object files that libtool operates on do not necessarily end in .o.

Nowadays, the computation of LTLIBOBJS from LIBOBJS is performed automatically by Autoconf (see Section “AC_LIBOBJ vs. LIBOBJS” in The Autoconf Manual).

8.3.9 Common Issues Related to Libtool’s Use

8.3.9.1 required file `./ltmain.sh' not found

Libtool comes with a tool called libtoolize that will install libtool’s supporting files into a package. Running this command will install ltmain.sh. You should execute it before

aclocal and automake.

People upgrading old packages to newer autotools are likely to face this issue because older Automake versions used to call libtoolize. Therefore old build scripts do not call

libtoolize.

Since Automake 1.6, it has been decided that running libtoolize was none of Automake’s business. Instead, that functionality has been moved into the autoreconf command (see Section “Using autoreconf” in The Autoconf Manual). If you do not want to remember what to run and when, just learn the autoreconf command. Hopefully, replacing existing bootstrap.sh or autogen.sh scripts by a call to autoreconf should also free you from any similar incompatible change in the future.

8.3.9.2 Objects created with both libtool and without

Sometimes, the same source file is used both to build a libtool library and to build another non-libtool target (be it a program or another library).

Let’s consider the following Makefile.am.

     bin_PROGRAMS = prog
     prog_SOURCES = prog.c foo.c ...
     lib_LTLIBRARIES = libfoo.la
     libfoo_la_SOURCES = foo.c ...

(In this trivial case the issue could be avoided by linking libfoo.la with prog instead of listing foo.c in prog_SOURCES. But let’s assume we really want to keep prog and libfoo.la separate.)

Technically, it means that we should build foo.$(OBJEXT) for prog, and foo.lo for libfoo.la. The problem is that in the course of creating foo.lo, libtool may erase (or replace) foo.$(OBJEXT) – and this cannot be avoided.

Therefore, when Automake detects this situation it will complain with a message such as

     object `foo.$(OBJEXT)' created both with libtool and without

A workaround for this issue is to ensure that these two objects get different basenames. As explained in Section 27.5 [renamed objects], page 86, this happens automatically when per-targets flags are used.

     bin_PROGRAMS = prog
     prog_SOURCES = prog.c foo.c ...
     prog_CFLAGS = $(AM_CFLAGS)
     lib_LTLIBRARIES = libfoo.la
     libfoo_la_SOURCES = foo.c ...

Adding prog_CFLAGS = $(AM_CFLAGS) is almost a no-op, because when the prog_CFLAGS is defined, it is used instead of AM_CFLAGS. However as a side effect it will cause prog.c and

foo.c to be compiled as prog-prog.$(OBJEXT) and prog-foo.$(OBJEXT) which solves the

issue.

8.4 Program and Library Variables

Associated with each program are a collection of variables which can be used to modify how that program is built. There is a similar list of such variables for each library. The canonical name of the program (or library) is used as a base for naming these variables.

In the list below, we use the name “maude” to refer to the program or library. In your Makefile.am you would replace this with the canonical name of your program. This list also refers to “maude” as a program, but in general the same rules apply for both static and dynamic libraries; the documentation below notes situations where programs and libraries differ.

‘maude_SOURCES’

This variable, if it exists, lists all the source files which are compiled to build the program. These files are added to the distribution by default. When building the program, Automake will cause each source file to be compiled to a single .o file (or .lo when using libtool). Normally these object files are named after the source file, but other factors can change this. If a file in the ‘_SOURCES’ variable has an unrecognized extension, Automake will do one of two things with it. If a suffix rule exists for turning files with the unrecognized extension into .o files, then automake will treat this file as it will any other source file (see Section 8.13 [Support for Other Languages], page 47). Otherwise, the file will be ignored as though it were a header file. The prefixes ‘dist_’ and ‘nodist_’ can be used to control whether files listed in a ‘_SOURCES’ variable are distributed. ‘dist_’ is redundant, as sources are distributed by default, but it can be specified for clarity if desired. It is possible to have both ‘dist_’ and ‘nodist_’ variants of a given ‘_SOURCES’ variable at once; this lets you easily distribute some files and not others, for instance:

     nodist_maude_SOURCES = nodist.c
     dist_maude_SOURCES = dist-me.c

By default the output file (on Unix systems, the .o file) will be put into the current build directory. However, if the option subdir-objects is in effect in the current directory then the .o file will be put into the subdirectory named after the source file. For instance, with subdir-objects enabled, sub/dir/file.c will be compiled to sub/dir/file.o. Some people prefer this mode of operation. You can specify subdir-objects in AUTOMAKE_OPTIONS (see Chapter 17 [Options], page 68).

‘EXTRA_maude_SOURCES’

Automake needs to know the list of files you intend to compile statically. For one thing, this is the only way Automake has of knowing what sort of language support a given Makefile.in requires.^{5} This means that, for example, you can’t put a configure substitution like ‘@my_sources@’ into a ‘_SOURCES’ variable. If you intend to conditionally compile source files and use configure to substitute the appropriate object names into, e.g., ‘_LDADD’ (see below), then you should list the corresponding source files in the ‘EXTRA_’ variable. This variable also supports ‘dist_’ and ‘nodist_’ prefixes, e.g.,

‘nodist_EXTRA_maude_SOURCES’.
‘maude_AR’

A static library is created by default by invoking $(AR) $(ARFLAGS) followed by the name of the library and then the objects being put into the library. You can override this by setting the ‘_AR’ variable. This is usually used with C++; some C++ compilers require a special invocation in order to instantiate all the templates which should go into a library. For instance, the SGI C++ compiler likes this variable set like so:

     libmaude_a_AR = $(CXX) -ar -o
‘maude_LIBADD’

Extra objects can be added to a library using the ‘_LIBADD’ variable. For instance this should be used for objects determined by configure (see Section 8.2 [A Library], page 30).

5 There are other, more obscure reasons for this limitation as well.

‘maude_LDADD’

Extra objects can be added to a program by listing them in the ‘_LDADD’ variable. For instance this should be used for objects determined by configure (see Section 8.1.2 [Linking], page 27). ‘_LDADD’ and ‘_LIBADD’ are inappropriate for passing program-specific linker flags (except for ‘-l’, ‘-L’, ‘-dlopen’ and ‘-dlpreopen’). Use the ‘_LDFLAGS’ variable for this purpose. For instance, if your configure.ac uses AC_PATH_XTRA, you could link your program against the X libraries like so:

     maude_LDADD = $(X_PRE_LIBS) $(X_LIBS) $(X_EXTRA_LIBS)
‘maude_LDFLAGS’

This variable is used to pass extra flags to the link step of a program or a shared library.

‘maude_DEPENDENCIES’

It is also occasionally useful to have a program depend on some other target which is not actually part of that program. This can be done using the ‘_DEPENDENCIES’ variable. Each program depends on the contents of such a variable, but no further interpretation is done. If ‘_DEPENDENCIES’ is not supplied, it is computed by Automake. The automatically-assigned value is the contents of ‘_LDADD’ or ‘_LIBADD’, with most configure substitutions, ‘-l’, ‘-L’, ‘-dlopen’ and ‘-dlpreopen’ options removed. The configure substitutions that are left in are only ‘$(LIBOBJS)’ and ‘$(ALLOCA)’; these are left because it is known that they will not cause an invalid value for ‘_DEPENDENCIES’ to be generated.

‘maude_LINK’

You can override the linker on a per-program basis. By default the linker is chosen according to the languages used by the program. For instance, a program that includes C++ source code would use the C++ compiler to link. The ‘_LINK’ variable must hold the name of a command which can be passed all the .o file names as arguments. Note that the name of the underlying program is not passed to ‘_LINK’; typically one uses ‘$@’:

     maude_LINK = $(CCLD) -magic -o $@
‘maude_CCASFLAGS’
‘maude_CFLAGS’
‘maude_CPPFLAGS’
‘maude_CXXFLAGS’
‘maude_FFLAGS’
‘maude_GCJFLAGS’
‘maude_LFLAGS’
‘maude_OBJCFLAGS’
‘maude_RFLAGS’
‘maude_YFLAGS’

Automake allows you to set compilation flags on a per-program (or per-library) basis. A single source file can be included in several programs, and it will potentially be compiled with different flags for each program. This works for any language directly supported by Automake. These per-target compilation flags are

‘_CCASFLAGS’, ‘_CFLAGS’, ‘_CPPFLAGS’, ‘_CXXFLAGS’, ‘_FFLAGS’, ‘_GCJFLAGS’,
‘_LFLAGS’, ‘_OBJCFLAGS’, ‘_RFLAGS’, and ‘_YFLAGS’.

When using a per-target compilation flag, Automake will choose a different name for the intermediate object files. Ordinarily a file like sample.c will be compiled to produce sample.o. However, if the program’s ‘_CFLAGS’ variable is set, then the object file will be named, for instance, maude-sample.o. (See also Section 27.5 [renamed objects], page 86.) In compilations with per-target flags, the ordinary ‘AM_’ form of the flags variable is not automatically included in the compilation (however, the user form of the variable is included). So for instance, if you want the hypothetical maude compilations to also use the value of ‘AM_CFLAGS’, you would need to write:

     maude_CFLAGS = ... your flags ... $(AM_CFLAGS)
‘maude_SHORTNAME’

On some platforms the allowable file names are very short. In order to support these systems and per-target compilation flags at the same time, Automake allows you to set a “short name” which will influence how intermediate object files are named. For instance, in the following example,

     bin_PROGRAMS = maude
     maude_CPPFLAGS = -DSOMEFLAG
     maude_SHORTNAME = m
     maude_SOURCES = sample.c ...

the object file would be named m-sample.o rather than maude-sample.o. This facility is rarely needed in practice, and we recommend avoiding it until you find it is required.

8.5 Default _SOURCES

_SOURCES variables are used to specify source files of programs (see Section 8.1 [A Program], page 26), libraries (see Section 8.2 [A Library], page 30), and Libtool libraries (see Section 8.3 [A Shared Library], page 30).

When no such variable is specified for a target, Automake will define one itself. The default is to compile a single C file whose base name is the name of the target itself, with any extension replaced by .c. (Defaulting to C is terrible but we are stuck with it for historical reasons.)

For example if you have the following somewhere in your Makefile.am with no corre-

sponding ‘libfoo_a_SOURCES’:
lib_LIBRARIES = libfoo.a sub/libc++.a

libfoo.a will be built using a default source file named libfoo.c, and sub/libc++.a will be built from sub/libc++.c. (In older versions sub/libc++.a would be built from sub_libc___a.c, i.e., the default source was the canonized name of the target, with .c appended. We believe the new behavior is more sensible, but for backward compatibility automake will use the old name if a file or a rule with that name exist.)

Default sources are mainly useful in test suites, when building many tests programs each from a single source. For instance in

     check_PROGRAMS = test1 test2 test3
test1, test2, and test3 will be built from test1.c, test2.c, and test3.c.

Another case where is this convenient is building many Libtool modules (moduleN.la), each defined in its own file (moduleN.c).

     AM_LDFLAGS = -module
     lib_LTLIBRARIES = module1.la module2.la module3.la

Finally, there is one situation where this default source computation needs to be avoided: when a target should not be built from sources. We already saw such an example in Section 3.3 [true], page 8; this happens when all the constituents of a target have already been compiled and need just to be combined using a _LDADD variable. Then it is necessary to define an empty _SOURCES variable, so that automake does not compute a default.

     bin_PROGRAMS = target
     target_SOURCES =
     target_LDADD = libmain.a libmisc.a

8.6 Special handling for LIBOBJS and ALLOCA

Automake explicitly recognizes the use of $(LIBOBJS) and $(ALLOCA), and uses this information, plus the list of LIBOBJS files derived from configure.ac to automatically include the appropriate source files in the distribution (see Chapter 14 [Dist], page 62). These source files are also automatically handled in the dependency-tracking scheme; see See Section 8.15 [Dependencies], page 48.

$(LIBOBJS) and $(ALLOCA) are specially recognized in any ‘_LDADD’ or ‘_LIBADD’ variable.

8.7 Variables used when building a program

Occasionally it is useful to know which Makefile variables Automake uses for compilations; for instance you might need to do your own compilation in some special cases.

Some variables are inherited from Autoconf; these are CC, CFLAGS, CPPFLAGS, DEFS,

LDFLAGS, and LIBS.

There are some additional variables which Automake itself defines:

AM_CPPFLAGS

The contents of this variable are passed to every compilation which invokes the C preprocessor; it is a list of arguments to the preprocessor. For instance, ‘-I’ and ‘-D’ options should be listed here. Automake already provides some ‘-I’ options automatically. In particular it generates ‘-I$(srcdir)’, ‘-I.’, and a ‘-I’ pointing to the directory holding

          config.h (if you’ve used AC_CONFIG_HEADERS or AM_CONFIG_HEADER). You

can disable the default ‘-I’ options using the ‘nostdinc’ option.

AM_CPPFLAGS is ignored in preference to a per-executable (or per-library) _ CPPFLAGS variable if it is defined.

INCLUDES This does the same job as ‘AM_CPPFLAGS’ (or any per-target ‘_CPPFLAGS’ variable if it is used). It is an older name for the same functionality. This variable is deprecated; we suggest using ‘AM_CPPFLAGS’ and per-target ‘_CPPFLAGS’ instead.

AM_CFLAGS

This is the variable which the Makefile.am author can use to pass in additional C compiler flags. It is more fully documented elsewhere. In some situations, this is not used, in preference to the per-executable (or per-library) _CFLAGS.

COMPILE This is the command used to actually compile a C source file. The filename is appended to form the complete command line.

AM_LDFLAGS

This is the variable which the Makefile.am author can use to pass in additional linker flags. In some situations, this is not used, in preference to the perexecutable (or per-library) _LDFLAGS.

LINK This is the command used to actually link a C program. It already includes ‘-o $@’ and the usual variable references (for instance, CFLAGS); it takes as “arguments” the names of the object files and libraries to link in.

8.8 Yacc and Lex support

Automake has somewhat idiosyncratic support for Yacc and Lex.

Automake assumes that the .c file generated by yacc (or lex) should be named using the basename of the input file. That is, for a yacc source file foo.y, Automake will cause the intermediate file to be named foo.c (as opposed to y.tab.c, which is more traditional).

The extension of a yacc source file is used to determine the extension of the resulting ‘C’ or ‘C++’ file. Files with the extension ‘.y’ will be turned into ‘.c’ files; likewise, ‘.yy’ will become ‘.cc’; ‘.y++’, ‘c++’; and ‘.yxx’, ‘.cxx’.

Likewise, lex source files can be used to generate ‘C’ or ‘C++’; the extensions ‘.l’, ‘.ll’, ‘.l++’, and ‘.lxx’ are recognized.

You should never explicitly mention the intermediate (‘C’ or ‘C++’) file in any ‘SOURCES’ variable; only list the source file.

The intermediate files generated by yacc (or lex) will be included in any distribution that is made. That way the user doesn’t need to have yacc or lex.

If a yacc source file is seen, then your configure.ac must define the variable ‘YACC’. This is most easily done by invoking the macro ‘AC_PROG_YACC’ (see Section “Particular Program Checks” in The Autoconf Manual).

When yacc is invoked, it is passed ‘YFLAGS’ and ‘AM_YFLAGS’. The former is a user variable and the latter is intended for the Makefile.am author.

‘AM_YFLAGS’ is usually used to pass the -d option to yacc. Automake knows what this means and will automatically adjust its rules to update and distribute the header file built by yacc -d. What Automake cannot guess, though, is where this header will be used: it is up to you to ensure the header gets built before it is first used. Typically this is necessary in order for dependency tracking to work when the header is included by another file. The common solution is listing the header file in BUILT_SOURCES (see Section 9.4 [Sources], page 51) as follows.

BUILT_SOURCES = parser.h
AM_YFLAGS = -d
bin_PROGRAMS = foo
foo_SOURCES = ... parser.y ...

If a lex source file is seen, then your configure.ac must define the variable ‘LEX’. You can use ‘AC_PROG_LEX’ to do this (see Section “Particular Program Checks” in The Autoconf Manual), but using AM_PROG_LEX macro (see Section 5.6 [Macros], page 18) is recommended.

When lex is invoked, it is passed ‘LFLAGS’ and ‘AM_LFLAGS’. The former is a user variable and the latter is intended for the Makefile.am author.

Automake makes it possible to include multiple yacc (or lex) source files in a single program. When there is more than one distinct yacc (or lex) source file in a directory, Automake uses a small program called ylwrap to run yacc (or lex) in a subdirectory. This is necessary because yacc’s output filename is fixed, and a parallel make could conceivably invoke more than one instance of yacc simultaneously. The ylwrap program is distributed with Automake. It should appear in the directory specified by ‘AC_CONFIG_AUX_DIR’ (see Section “Finding ‘configure’ Input” in The Autoconf Manual), or the current directory if that macro is not used in configure.ac.

For yacc, simply managing locking is insufficient. The output of yacc always uses the same symbol names internally, so it isn’t possible to link two yacc parsers into the same executable.

We recommend using the following renaming hack used in gdb:

#define yymaxdepth c_maxdepth
#define yyparse c_parse
#define yylex c_lex
#define yyerror c_error
#define yylval c_lval
#define yychar c_char
#define yydebug c_debug
#define yypact c_pact
#define yyr1 c_r1
#define yyr2 c_r2
#define yydef c_def
#define yychk c_chk
#define yypgo c_pgo
#define yyact c_act
#define yyexca c_exca
#define yyerrflag c_errflag
#define yynerrs c_nerrs
#define yyps c_ps
#define yypv c_pv
#define yys c_s
#define yy_yys c_yys

#define yystate c_state
#define yytmp c_tmp
#define yyv c_v
#define yy_yyv c_yyv
#define yyval c_val
#define yylloc c_lloc
#define yyreds c_reds
#define yytoks c_toks
#define yylhs c_yylhs
#define yylen c_yylen
#define yydefred c_yydefred
#define yydgoto c_yydgoto
#define yysindex c_yysindex
#define yyrindex c_yyrindex
#define yygindex c_yygindex
#define yytable  c_yytable
#define yycheck  c_yycheck
#define yyname   c_yyname
#define yyrule   c_yyrule

For each define, replace the ‘c_’ prefix with whatever you like. These defines work for bison, byacc, and traditional yaccs. If you find a parser generator that uses a symbol not covered here, please report the new name so it can be added to the list.

8.9 C++ Support

Automake includes full support for C++.

Any package including C++ code must define the output variable ‘CXX’ in configure.ac; the simplest way to do this is to use the AC_PROG_CXX macro (see Section “Particular Program Checks” in The Autoconf Manual).

A few additional variables are defined when a C++ source file is seen:

CXX The name of the C++ compiler.

CXXFLAGS Any flags to pass to the C++ compiler.

AM_CXXFLAGS

The maintainer’s variant of CXXFLAGS.

CXXCOMPILE

The command used to actually compile a C++ source file. The file name is appended to form the complete command line.

CXXLINK The command used to actually link a C++ program.

8.10 Assembly Support

Automake includes some support for assembly code.

The variable CCAS holds the name of the compiler used to build assembly code. This compiler must work a bit like a C compiler; in particular it must accept ‘-c’ and ‘-o’. The value of CCASFLAGS is passed to the compilation.

You are required to set CCAS and CCASFLAGS via configure.ac. The autoconf macro AM_PROG_AS will do this for you. Unless they are already set, it simply sets CCAS to the C compiler and CCASFLAGS to the C compiler flags.

Only the suffixes ‘.s’ and ‘.S’ are recognized by automake as being files containing assembly code.

8.11 Fortran 77 Support

Automake includes full support for Fortran 77.

Any package including Fortran 77 code must define the output variable ‘F77’ in configure.ac; the simplest way to do this is to use the AC_PROG_F77 macro (see Section “Particular Program Checks” in The Autoconf Manual).

A few additional variables are defined when a Fortran 77 source file is seen:

F77 The name of the Fortran 77 compiler.

FFLAGS Any flags to pass to the Fortran 77 compiler.

AM_FFLAGS

The maintainer’s variant of FFLAGS.

RFLAGS Any flags to pass to the Ratfor compiler.

AM_RFLAGS

The maintainer’s variant of RFLAGS.

F77COMPILE

The command used to actually compile a Fortran 77 source file. The file name is appended to form the complete command line.

FLINK The command used to actually link a pure Fortran 77 program or shared library.

Automake can handle preprocessing Fortran 77 and Ratfor source files in addition to compiling them^{6}. Automake also contains some support for creating programs and shared libraries that are a mixture of Fortran 77 and other languages (see Section 8.11.3 [Mixing Fortran 77 With C and C++], page 44).

These issues are covered in the following sections.

8.11.1 Preprocessing Fortran 77

N.f is made automatically from N.F or N.r. This rule runs just the preprocessor to convert a preprocessable Fortran 77 or Ratfor source file into a strict Fortran 77 source file. The precise command used is as follows:

.F        $(F77) -F $(DEFS) $(INCLUDES) $(AM_CPPFLAGS) $(CPPFLAGS) $(AM_
          FFLAGS) $(FFLAGS)
.r        $(F77) -F $(AM_FFLAGS) $(FFLAGS) $(AM_RFLAGS) $(RFLAGS)

6 Much, if not most, of the information in the following sections pertaining to preprocessing Fortran 77

programs was taken almost verbatim from Section “Catalogue of Rules” in The GNU Make Manual.

8.11.2 Compiling Fortran 77 Files

N.o is made automatically from N.f, N.F or N.r by running the Fortran 77 compiler. The precise command used is as follows:

.f        $(F77) -c $(AM_FFLAGS) $(FFLAGS)
.F        $(F77) -c $(DEFS) $(INCLUDES) $(AM_CPPFLAGS) $(CPPFLAGS) $(AM_
          FFLAGS) $(FFLAGS)
.r        $(F77) -c $(AM_FFLAGS) $(FFLAGS) $(AM_RFLAGS) $(RFLAGS)

8.11.3 Mixing Fortran 77 With C and C++

Automake currently provides limited support for creating programs and shared libraries that are a mixture of Fortran 77 and C and/or C++. However, there are many other issues related to mixing Fortran 77 with other languages that are not (currently) handled by Automake, but that are handled by other packages^{7}.

7 For example, the cfortran package (http://www-zeus.desy.de/~burow/cfortran/) addresses all of these

inter-language issues, and runs under nearly all Fortran 77, C and C++ compilers on nearly all platforms.

However, cfortran is not yet Free Software, but it will be in the next major release.

Automake can help in two ways:

1. Automatic selection of the linker depending on which combinations of source code.

2. Automatic selection of the appropriate linker flags (e.g. ‘-L’ and ‘-l’) to pass to the automatically selected linker in order to link in the appropriate Fortran 77 intrinsic and run-time libraries. These extra Fortran 77 linker flags are supplied in the output variable FLIBS by the AC_F77_LIBRARY_LDFLAGS Autoconf macro supplied with newer versions of Autoconf (Autoconf version 2.13 and later). See Section “Fortran 77 Compiler Characteristics” in The Autoconf .

If Automake detects that a program or shared library (as mentioned in some _PROGRAMS or _LTLIBRARIES primary) contains source code that is a mixture of Fortran 77 and C and/or C++, then it requires that the macro AC_F77_LIBRARY_LDFLAGS be called in configure.ac, and that either $(FLIBS) appear in the appropriate _LDADD (for programs) or _LIBADD (for shared libraries) variables. It is the responsibility of the person writing the Makefile.am to make sure that $(FLIBS) appears in the appropriate _LDADD or _LIBADD variable.

For example, consider the following Makefile.am:

     bin_PROGRAMS = foo
     foo_SOURCES  = main.cc foo.f
     foo_LDADD    = libfoo.la $(FLIBS)
     pkglib_LTLIBRARIES = libfoo.la
     libfoo_la_SOURCES  = bar.f baz.c zardoz.cc
     libfoo_la_LIBADD   = $(FLIBS)

In this case, Automake will insist that AC_F77_LIBRARY_LDFLAGS is mentioned in configure.ac. Also, if $(FLIBS) hadn’t been mentioned in foo_LDADD and libfoo_la_LIBADD, then Automake would have issued a warning.

8.11.3.1 How the Linker is Chosen

The following diagram demonstrates under what conditions a particular linker is chosen by Automake.

For example, if Fortran 77, C and C++ source code were to be compiled into a program, then the C++ linker will be used. In this case, if the C or Fortran 77 linkers required any special libraries that weren’t included by the C++ linker, then they must be manually added to an _LDADD or _LIBADD variable by the user writing the Makefile.am.

\              Linker
source      \
code        \     C        C++     Fortran
-----------------  +---------+---------+---------+
|         |         |         |
C                  |    x    |         |         |
|         |         |         |
+---------+---------+---------+
|         |         |         |
C++            |         |    x    |         |
|         |         |         |
+---------+---------+---------+
|         |         |         |
Fortran  |         |         |    x    |
|         |         |         |
+---------+---------+---------+
|         |         |         |
C + C++            |         |    x    |         |
|         |         |         |
+---------+---------+---------+
|         |         |         |
C +       Fortran  |         |         |    x    |
|         |         |         |
+---------+---------+---------+
|         |         |         |
C++ + Fortran  |         |    x    |         |
|         |         |         |
+---------+---------+---------+
|         |         |         |
C + C++ + Fortran  |         |    x    |         |
|         |         |         |
+---------+---------+---------+

8.12 Java Support

Automake includes support for compiled Java, using gcj, the Java front end to the GNU Compiler Collection.

Any package including Java code to be compiled must define the output variable ‘GCJ’ in configure.ac; the variable ‘GCJFLAGS’ must also be defined somehow (either in configure.ac or Makefile.am). The simplest way to do this is to use the AM_PROG_GCJ macro.

By default, programs including Java source files are linked with gcj.

As always, the contents of ‘AM_GCJFLAGS’ are passed to every compilation invoking gcj (in its role as an ahead-of-time compiler – when invoking it to create .class files, ‘AM_JAVACFLAGS’ is used instead). If it is necessary to pass options to gcj from Makefile.am, this variable, and not the user variable ‘GCJFLAGS’, should be used.

gcj can be used to compile .java, .class, .zip, or .jar files.

When linking, gcj requires that the main class be specified using the ‘--main=’ option. The easiest way to do this is to use the _LDFLAGS variable for the program.

8.13 Support for Other Languages

Automake currently only includes full support for C, C++ (see Section 8.9 [C++ Support], page 42), Fortran 77 (see Section 8.11 [Fortran 77 Support], page 43), and Java (see Section 8.12 [Java Support], page 46). There is only rudimentary support for other languages, support for which will be improved based on user demand.

Some limited support for adding your own languages is available via the suffix rule handling; see Section 18.2 [Suffixes], page 71.

8.14 Automatic de-ANSI-fication

Although the GNU standards allow the use of ANSI C, this can have the effect of limiting portability of a package to some older compilers (notably the SunOS C compiler).

Automake allows you to work around this problem on such machines by de-ANSI-fying each source file before the actual compilation takes place.

If the Makefile.am variable AUTOMAKE_OPTIONS (see Chapter 17 [Options], page 68) contains the option ansi2knr then code to handle de-ANSI-fication is inserted into the

generated Makefile.in.

This causes each C source file in the directory to be treated as ANSI C. If an ANSI C compiler is available, it is used. If no ANSI C compiler is available, the ansi2knr program is used to convert the source files into K&R C, which is then compiled.

The ansi2knr program is simple-minded. It assumes the source code will be formatted in a particular way; see the ansi2knr man page for details.

Support for de-ANSI-fication requires the source files ansi2knr.c and ansi2knr.1 to be in the same package as the ANSI C source; these files are distributed with Automake. Also, the package configure.ac must call the macro AM_C_PROTOTYPES (see Section 5.6 [Macros], page 18).

Automake also handles finding the ansi2knr support files in some other directory in the current package. This is done by prepending the relative path to the appropriate directory to the ansi2knr option. For instance, suppose the package has ANSI C code in the src and lib subdirectories. The files ansi2knr.c and ansi2knr.1 appear in lib. Then this

could appear in src/Makefile.am:
     AUTOMAKE_OPTIONS = ../lib/ansi2knr

If no directory prefix is given, the files are assumed to be in the current directory.

Note that automatic de-ANSI-fication will not work when the package is being built for a different host architecture. That is because automake currently has no way to build ansi2knr for the build machine.

Using LIBOBJS with source de-ANSI-fication used to require hand-crafted code in configure to append $U to basenames in LIBOBJS. This is no longer true today. Starting with version 2.54, Autoconf takes care of rewriting LIBOBJS and LTLIBOBJS. (see Section “AC_LIBOBJ vs. LIBOBJS” in The Autoconf Manual)

8.15 Automatic dependency tracking

As a developer it is often painful to continually update the Makefile.in whenever the include-file dependencies change in a project. Automake supplies a way to automatically track dependency changes.

Automake always uses complete dependencies for a compilation, including system headers. Automake’s model is that dependency computation should be a side effect of the build. To this end, dependencies are computed by running all compilations through a special wrapper program called depcomp. depcomp understands how to coax many different C and C++ compilers into generating dependency information in the format it requires. automake -a will install depcomp into your source tree for you. If depcomp can’t figure out how to properly invoke your compiler, dependency tracking will simply be disabled for your build.

Experience with earlier versions of Automake^{1} taught us that it is not reliable to generate dependencies only on the maintainer’s system, as configurations vary too much. So instead Automake implements dependency tracking at build time.

Automatic dependency tracking can be suppressed by putting no-dependencies in the

variable AUTOMAKE_OPTIONS, or passing no-dependencies as an argument to AM_INIT_

AUTOMAKE (this should be the preferred way). Or, you can invoke automake with the -i option. Dependency tracking is enabled by default.

The person building your package also can choose to disable dependency tracking by

configuring with --disable-dependency-tracking.

8.16 Support for executable extensions

On some platforms, such as Windows, executables are expected to have an extension such as ‘.exe’. On these platforms, some compilers (GCC among them) will automatically generate foo.exe when asked to generate foo.

Automake provides mostly-transparent support for this. Unfortunately mostly doesn’t yet mean fully. Until the English dictionary is revised, you will have to assist Automake if your package must support those platforms.

One thing you must be aware of is that, internally, Automake rewrites something like this:

     bin_PROGRAMS = liver

to this:

     bin_PROGRAMS = liver$(EXEEXT)

1 See http://sources.redhat.com/automake/dependencies.html for more information on the history

and experiences with automatic dependency tracking in Automake

The targets Automake generates are likewise given the ‘$(EXEEXT)’ extension. EXEEXT However, Automake cannot apply this rewriting to configure substitutions. This means

that if you are conditionally building a program using such a substitution, then your configure.ac must take care to add ‘$(EXEEXT)’ when constructing the output variable.

With Autoconf 2.13 and earlier, you must explicitly use AC_EXEEXT to get this support. With Autoconf 2.50, AC_EXEEXT is run automatically if you configure a compiler (say,

through AC_PROG_CC).

Sometimes maintainers like to write an explicit link rule for their program. Without executable extension support, this is easy—you simply write a rule whose target is the name of the program. However, when executable extension support is enabled, you must instead add the ‘$(EXEEXT)’ suffix.

Unfortunately, due to the change in Autoconf 2.50, this means you must always add this extension. However, this is a problem for maintainers who know their package will never run on a platform that has executable extensions. For those maintainers, the no-exeext option (see Chapter 17 [Options], page 68) will disable this feature. This works in a fairly ugly way; if no-exeext is seen, then the presence of a rule for a target named foo in Makefile.am will override an automake-generated rule for foo$(EXEEXT). Without the no-exeext option, this use will give a diagnostic.

9 Other Derived Objects

Automake can handle derived objects which are not C programs. Sometimes the support for actually building such objects must be explicitly supplied, but Automake will still automatically handle installation and distribution.

9.1 Executable Scripts

It is possible to define and install programs which are scripts. Such programs are listed using the ‘SCRIPTS’ primary name. Automake doesn’t define any dependencies for scripts; the Makefile.am should include the appropriate rules.

Automake does not assume that scripts are derived objects; such objects must be deleted by hand (see Chapter 13 [Clean], page 62).

The automake program itself is a Perl script that is generated from automake.in. Here is how this is handled:

     bin_SCRIPTS = automake
     CLEANFILES = $(bin_SCRIPTS)
     do_subst = sed -e 's,[@]datadir[@],$(datadir),g' \
                 -e 's,[@]PERL[@],$(PERL),g' \
                 -e 's,[@]PACKAGE[@],$(PACKAGE),g' \
                 -e 's,[@]VERSION[@],$(VERSION),g' \
                 ...
     automake: automake.in Makefile
             $(do_subst) < $(srcdir)/automake.in > automake

             chmod +x automake

Because—as we have just seen—scripts can be built, they are not distributed by default. Scripts that should be distributed can be specified using a dist_prefix as in other primaries. For instance the following Makefile.am declares that my_script should be distributed and installed in $(sbindir).

     dist_sbin_SCRIPTS = my_script

Script objects can be installed in bindir, sbindir, libexecdir, or pkgdatadir.

Scripts that need not being installed can be listed in noinst_SCRIPTS, and among them, those which are needed only by make check should go in check_SCRIPTS.

9.2 Header files

Header files that must be installed are specified by the ‘HEADERS’ family of variables. Headers can be installed in includedir, oldincludedir, pkgincludedir or any other directory you may have defined (see Section 2.3 [Uniform], page 3). For instance

     include_HEADERS = foo.h bar/bar.h
will install the two files as $(includedir)/foo.h and $(includedir)/bar.h.

The ‘nobase_’ prefix is also supported,

     nobase_include_HEADERS = foo.h bar/bar.h
will install the two files as $(includedir)/foo.h and $(includedir)/bar/bar.h (see

Chapter 7 [Alternative], page 26).

Usually, only header files that accompany installed libraries need to be installed. Headers used by programs or convenience libraries are not installed. The noinst_HEADERS variable can be used for such headers. However when the header actually belongs to one convenient library or program, we recommend listing it in the program’s or library’s ‘_SOURCES’ variable (see Section 8.1.1 [Program Sources], page 27) instead of in noinst_HEADERS. This is clearer for the Makefile.am reader. noinst_HEADERS would be the right variable to use in a directory containing only headers and no associated library or program.

All header files must be listed somewhere; in a ‘_SOURCES’ variable or in a ‘_HEADERS’ variable. Missing ones will not appear in the distribution.

For header files that are built and must not be distributed, use the ‘nodist_’ prefix as in

nodist_include_HEADERS or nodist_prog_SOURCES. If these generated headers are needed

during the build, you must also ensure they exist before they are used, see See Section 9.4 [Sources], page 51.

9.3 Architecture-independent data files

Automake supports the installation of miscellaneous data files using the ‘DATA’ family of variables.

Such data can be installed in the directories datadir, sysconfdir, sharedstatedir,

localstatedir, or pkgdatadir.

By default, data files are not included in a distribution. Of course, you can use the ‘dist_’ prefix to change this on a per-variable basis.

Here is how Automake declares its auxiliary data files:

     dist_pkgdata_DATA = clean-kr.am clean.am ...

9.4 Built sources

Because Automake’s automatic dependency tracking works as a side-effect of compilation (see Section 8.15 [Dependencies], page 48) there is a bootstrap issue: a target should not be compiled before its dependencies are made, but these dependencies are unknown until the target is first compiled.

Ordinarily this is not a problem, because dependencies are distributed sources: they preexist and do not need to be built. Suppose that foo.c includes foo.h. When it first compiles foo.o, make only knows that foo.o depends on foo.c. As a side-effect of this compilation depcomp records the foo.h dependency so that following invocations of make will honor it. In these conditions, it’s clear there is no problem: either foo.o doesn’t exist and has to be built (regardless of the dependencies), either accurate dependencies exist and they can be used to decide whether foo.o should be rebuilt.

It’s a different story if foo.h doesn’t exist by the first make run. For instance there might be a rule to build foo.h. This time file.o’s build will fail because the compiler can’t find foo.h. make failed to trigger the rule to build foo.h first by lack of dependency information.

The BUILT_SOURCES variable is a workaround for this problem. A source file listed in

BUILT_SOURCES is made on make all or make check (or even make install) before other

targets are processed. However, such a source file is not compiled unless explicitly requested by mentioning it in some other ‘_SOURCES’ variable.

So, to conclude our introductory example, we could use BUILT_SOURCES = foo.h to ensure foo.h gets built before any other target (including foo.o) during make all or make

check.

BUILT_SOURCES is actually a bit of a misnomer, as any file which must be created early in the build process can be listed in this variable. Moreover, all built sources do not necessarily have to be listed in BUILT_SOURCES. For instance a generated .c file doesn’t need to appear in BUILT_SOURCES (unless it is included by another source), because it’s a known dependency of the associated object.

It might be important to emphasize that BUILT_SOURCES is honored only by make all, make check and make install. This means you cannot build a specific target (e.g., make foo) in a clean tree if it depends on a built source. However it will succeed if you have run make all earlier, because accurate dependencies are already available.

The next section illustrates and discusses the handling of built sources on a toy example.

9.4.1 Built sources example

Suppose that foo.c includes bindir.h, which is installation-dependent and not distributed: it needs to be built. Here bindir.h defines the preprocessor macro bindir to the value of the make variable bindir (inherited from configure).

We suggest several implementations below. It’s not meant to be an exhaustive listing of all ways to handle built sources, but it will give you a few ideas if you encounter this issue.

First try

This first implementation will illustrate the bootstrap issue mentioned in the previous section (see Section 9.4 [Sources], page 51).

Here is a tentative Makefile.am.

# This won't work.
bin_PROGRAMS = foo
foo_SOURCES = foo.c
nodist_foo_SOURCES = bindir.h
CLEANFILES = bindir.h
bindir.h: Makefile
        echo '#define bindir "$(bindir)"' >$@

This setup doesn’t work, because Automake doesn’t know that foo.c includes bindir.h. Remember, automatic dependency tracking works as a side-effect of compilation, so the dependencies of foo.o will be known only after foo.o has been compiled (see Section 8.15 [Dependencies], page 48). The symptom is as follows.

% make
source='foo.c' object='foo.o' libtool=no \
depfile='.deps/foo.Po' tmpdepfile='.deps/foo.TPo' \
depmode=gcc /bin/sh ./depcomp \
gcc -I. -I. -g -O2 -c `test -f 'foo.c' || echo './'`foo.c
foo.c:2: bindir.h: No such file or directory
make: *** [foo.o] Error 1
Using BUILT_SOURCES

A solution is to require bindir.h to be built before anything else. This is what BUILT_ SOURCES is meant for (see Section 9.4 [Sources], page 51).

bin_PROGRAMS = foo
foo_SOURCES = foo.c
BUILT_SOURCES = bindir.h
CLEANFILES = bindir.h
bindir.h: Makefile
        echo '#define bindir "$(bindir)"' >$@

See how bindir.h get built first:

% make
echo '#define bindir "/usr/local/bin"' >bindir.h
make  all-am
make[1]: Entering directory `/home/adl/tmp'
source='foo.c' object='foo.o' libtool=no \
depfile='.deps/foo.Po' tmpdepfile='.deps/foo.TPo' \
depmode=gcc /bin/sh ./depcomp \
gcc -I. -I. -g -O2 -c `test -f 'foo.c' || echo './'`foo.c
gcc  -g -O2   -o foo  foo.o
make[1]: Leaving directory `/home/adl/tmp'

However, as said earlier, BUILT_SOURCES applies only to the all, check, and install targets. It still fails if you try to run make foo explicitly:

% make clean
test -z "bindir.h" || rm -f bindir.h
test -z "foo" || rm -f foo

     rm -f *.o
     % : > .deps/foo.Po # Suppress previously recorded dependencies
     % make foo
     source='foo.c' object='foo.o' libtool=no \
     depfile='.deps/foo.Po' tmpdepfile='.deps/foo.TPo' \
     depmode=gcc /bin/sh ./depcomp \
     gcc -I. -I. -g -O2 -c `test -f 'foo.c' || echo './'`foo.c
     foo.c:2: bindir.h: No such file or directory
     make: *** [foo.o] Error 1

Recording dependencies manually

Usually people are happy enough with BUILT_SOURCES because they never build targets such as make foo before make all, as in the previous example. However if this matters to you, you can avoid BUILT_SOURCES and record such dependencies explicitly in the Makefile.am.

     bin_PROGRAMS = foo
     foo_SOURCES = foo.c
     foo.$(OBJEXT): bindir.h
     CLEANFILES = bindir.h
     bindir.h: Makefile
             echo '#define bindir "$(bindir)"' >$@

You don’t have to list all the dependencies of foo.o explicitly, only those which might need to be built. If a dependency already exists, it will not hinder the first compilation and will be recorded by the normal dependency tracking code. (Note that after this first compilation the dependency tracking code will also have recorded the dependency between foo.o and bindir.h; so our explicit dependency is really useful to the first build only.)

Adding explicit dependencies like this can be a bit dangerous if you are not careful enough. This is due to the way Automake tries not to overwrite your rules (it assumes you know better than it). foo.$(OBJEXT): bindir.h supersedes any rule Automake may want to output to build foo.$(OBJEXT). It happens to work in this case because Automake doesn’t have to output any foo.$(OBJEXT): target: it relies on a suffix rule instead (i.e., .c.$(OBJEXT):). Always check the generated Makefile.in if you do this.

Build bindir.h from configure

It’s possible to define this preprocessor macro from configure, either in config.h (see Section “Defining Directories” in The Autoconf Manual), or by processing a bindir.h.in file using AC_CONFIG_FILES (see Section “Configuration Actions” in The Autoconf Manual).

At this point it should be clear that building bindir.h from configure work well for this example. bindir.h will exist before you build any target, hence will not cause any dependency issue.

The Makefile can be shrunk as follows. We do not even have to mention bindir.h.

     bin_PROGRAMS = foo
     foo_SOURCES = foo.c

However, it’s not always possible to build sources from configure, especially when these sources are generated by a tool that needs to be built first...

Build bindir.c, not bindir.h.

Another attractive idea is to define bindir as a variable or function exported from bindir.o, and build bindir.c instead of bindir.h.

     noinst_PROGRAMS = foo
     foo_SOURCES = foo.c bindir.h
     nodist_foo_SOURCES = bindir.c
     CLEANFILES = bindir.c
     bindir.c: Makefile
             echo 'const char bindir[] = "$(bindir)";' >$

bindir.h contains just the variable’s declaration and doesn’t need to be built, so it won’t cause any trouble. bindir.o is always dependent on bindir.c, so bindir.c will get built first.

Which is best?

There is no panacea, of course. Each solution has its merits and drawbacks.

You cannot use BUILT_SOURCES if the ability to run make foo on a clean tree is important to you.

You won’t add explicit dependencies if you are leery of overriding an Automake rule by mistake.

Building files from ./configure is not always possible, neither is converting .h files into .c files.

10 Other GNU Tools

Since Automake is primarily intended to generate Makefile.ins for use in GNU programs, it tries hard to interoperate with other GNU tools.

10.1 Emacs Lisp

Automake provides some support for Emacs Lisp. The ‘LISP’ primary is used to hold a list of .el files. Possible prefixes for this primary are ‘lisp_’ and ‘noinst_’. Note that

if lisp_LISP is defined, then configure.ac must run AM_PATH_LISPDIR (see Section 5.6

[Macros], page 18).

Automake will byte-compile all Emacs Lisp source files using the Emacs found by AM_ PATH_LISPDIR, if any was found.

Byte-compiled Emacs Lisp files are not portable among all versions of Emacs, so it makes sense to turn this off if you expect sites to have more than one version of Emacs installed. Furthermore, many packages don’t actually benefit from byte-compilation. Still, we recommend that you byte-compile your Emacs Lisp sources. It is probably better for sites with strange setups to cope for themselves than to make the installation less nice for everybody else.

There are two ways to avoid byte-compiling. Historically, we have recommended the following construct.

     lisp_LISP = file1.el file2.el

     ELCFILES =

ELCFILES is an internal Automake variable that normally lists all .elc files that must be byte-compiled. Automake defines ELCFILES automatically from lisp_LISP. Emptying this variable explicitly prevents byte-compilation to occur.

Since Automake 1.8, we now recommend using lisp_DATA instead. As in

     lisp_DATA = file1.el file2.el

Note that these two constructs are not equivalent. _LISP will not install a file if Emacs is not installed, while _DATA will always install its files.

10.2 Gettext

If AM_GNU_GETTEXT is seen in configure.ac, then Automake turns on support for GNU gettext, a message catalog system for internationalization (see Section “GNU Gettext” in GNU gettext utilities).

The gettext support in Automake requires the addition of two subdirectories to the package, intl and po. Automake insures that these directories exist and are mentioned in

SUBDIRS.

Lisp sources are not distributed by default. You can prefix the LISP primary with dist_, as in dist_lisp_LISP or dist_noinst_LISP, to indicate that these files should be distributed.

10.3 Libtool

Automake provides support for GNU Libtool (see Section “Introduction” in The Libtool Manual) with the ‘LTLIBRARIES’ primary. See Section 8.3 [A Shared Library], page 30.

10.4 Java

Automake provides some minimal support for Java compilation with the ‘JAVA’ primary.

Any .java files listed in a ‘_JAVA’ variable will be compiled with JAVAC at build time. By default, .class files are not included in the distribution.

Currently Automake enforces the restriction that only one ‘_JAVA’ primary can be used in a given Makefile.am. The reason for this restriction is that, in general, it isn’t possible to know which .class files were generated from which .java files – so it would be impossible to know which files to install where. For instance, a .java file can define multiple classes; the resulting .class file names cannot be predicted without parsing the .java file.

There are a few variables which are used when compiling Java sources:

JAVAC The name of the Java compiler. This defaults to ‘javac’.

JAVACFLAGS

The flags to pass to the compiler. This is considered to be a user variable (see Section 2.5 [User Variables], page 4).

AM_JAVACFLAGS

More flags to pass to the Java compiler. This, and not JAVACFLAGS, should be used when it is necessary to put Java compiler flags into Makefile.am.

JAVAROOT The value of this variable is passed to the ‘-d’ option to javac. It defaults to

          ‘$(top_builddir)’.
CLASSPATH_ENV

This variable is an sh expression which is used to set the CLASSPATH environment variable on the javac command line. (In the future we will probably handle class path setting differently.)

10.5 Python

Automake provides support for Python compilation with the ‘PYTHON’ primary.

Any files listed in a ‘_PYTHON’ variable will be byte-compiled with py-compile at install time. py-compile actually creates both standard (.pyc) and byte-compiled (.pyo) versions of the source files. Note that because byte-compilation occurs at install time, any files listed in ‘noinst_PYTHON’ will not be compiled. Python source files are included in the distribution by default.

Automake ships with an Autoconf macro called AM_PATH_PYTHON which will determine some Python-related directory variables (see below). If you have called AM_PATH_ PYTHON from configure.ac, then you may use the following variables to list you Python

source files in your variables: ‘python_PYTHON’, ‘pkgpython_PYTHON’, ‘pyexecdir_PYTHON’,

‘pkgpyexecdir_PYTHON’, depending where you want your files installed.

   AM_PATH_PYTHON([VERSION], [ACTION-IF-FOUND], [ACTION-IF-NOT-FOUND]) takes

three optional arguments. It will search a Python interpreter on the system. The first argument, if present, is the minimum version of Python required for this package: AM_PATH_PYTHON will skip any Python interpreter which is older than VERSION. If an interpreter is found and satisfies VERSION, then ACTION-IF-FOUND is run. Otherwise, ACTION-IF-NOT-FOUND is run.

If ACTION-IF-NOT-FOUND is not specified, the default is to abort configure. This is fine when Python is an absolute requirement for the package. Therefore if Python >= 2.2 is only optional to the package, AM_PATH_PYTHON could be called as follows.

       AM_PATH_PYTHON(2.2,, :)

AM_PATH_PYTHON creates several output variables based on the Python installation found during configuration.

PYTHON The name of the Python executable, or : if no suitable interpreter could be found. Assuming ACTION-IF-NOT-FOUND is used (otherwise ./configure will abort if Python is absent), the value of PYTHON can be used to setup a conditional in order to disable the relevant part of a build as follows.

                 AM_PATH_PYTHON(,, :)
                 AM_CONDITIONAL([HAVE_PYTHON], [test "$PYTHON" != :])

If the ACTION-IF-NOT-FOUND is specified

PYTHON_VERSION

The Python version number, in the form major.minor (e.g. ‘1.5’). This is currently the value of sys.version[:3].

PYTHON_PREFIX

The string ${prefix}. This term may be used in future work which needs the contents of Python’s sys.prefix, but general consensus is to always use the value from configure.

PYTHON_EXEC_PREFIX

The string ${exec_prefix}. This term may be used in future work which needs the contents of Python’s sys.exec_prefix, but general consensus is to always use the value from configure.

PYTHON_PLATFORM

The canonical name used by Python to describe the operating system, as given by sys.platform. This value is sometimes needed when building Python extensions.

pythondir

The directory name for the site-packages subdirectory of the standard Python install tree.

pkgpythondir

This is is the directory under pythondir which is named after the package. That is, it is ‘$(pythondir)/$(PACKAGE)’. It is provided as a convenience.

pyexecdir

This is the directory where Python extension modules (shared libraries) should be installed.

pkgpyexecdir

This is a convenience variable which is defined as ‘$(pyexecdir)/$(PACKAGE)’.

All these directory variables have values that start with either ${prefix} or ${exec_ prefix} unexpanded. This works fine in Makefiles, but it makes these variables hard to use in configure. This is mandated by the GNU coding standards, so that the user can run make prefix=/foo install. The Autoconf manual has a section with more details on this topic (see Section “Installation Directory Variables” in The Autoconf Manual).

11 Building documentation

Currently Automake provides support for Texinfo and man pages.

11.1 Texinfo

If the current directory contains Texinfo source, you must declare it with the ‘TEXINFOS’ primary. Generally Texinfo files are converted into info, and thus the info_TEXINFOS variable is most commonly used here. Any Texinfo source file must end in the .texi, .txi, or .texinfo extension. We recommend .texi for new manuals.

Automake generates rules to build .info, .dvi, .ps, .pdf and .html files from your Texinfo sources. The .info files are built by make all and installed by make install (unless you use no-installinfo, see below). The other files can be built on request by

make dvi, make ps, make pdf and make html.

If the .texi file @includes version.texi, then that file will be automatically generated. The file version.texi defines four Texinfo flag you can reference using @value{EDITION},

@value{VERSION}, @value{UPDATED}, and @value{UPDATED-MONTH}.
EDITION

VERSION Both of these flags hold the version number of your program. They are kept separate for clarity.

UPDATED This holds the date the primary .texi file was last modified.

UPDATED-MONTH

This holds the name of the month in which the primary .texi file was last modified.

The version.texi support requires the mdate-sh program; this program is supplied with Automake and automatically included when automake is invoked with the --add-

missing option.

If you have multiple Texinfo files, and you want to use the version.texi feature, then you have to have a separate version file for each Texinfo file. Automake will treat any include in a Texinfo file that matches ‘vers*.texi’ just as an automatically generated version file.

Sometimes an info file actually depends on more than one .texi file. For instance, in GNU Hello, hello.texi includes the file gpl.texi. You can tell Automake about these dependencies using the texi_TEXINFOS variable. Here is how GNU Hello does it:

     info_TEXINFOS = hello.texi
     hello_TEXINFOS = gpl.texi

By default, Automake requires the file texinfo.tex to appear in the same directory as the Texinfo source (this can be changed using the TEXINFO_TEX variable, see below). However, if you used AC_CONFIG_AUX_DIR in configure.ac (see Section “Finding ‘configure’ Input” in The Autoconf Manual), then texinfo.tex is looked for there. Automake supplies

texinfo.tex if ‘--add-missing’ is given.

The option ‘no-texinfo.tex’ can be used to eliminate the requirement for texinfo.tex. Use of the variable TEXINFO_TEX is preferable, however, because that allows the dvi, ps, and pdf targets to still work.

Automake generates an install-info rule; some people apparently use this. By default, info pages are installed by ‘make install’. This can be prevented via the no-installinfo option.

The following variables are used by the Texinfo build rules.

MAKEINFO The name of the program invoked to build .info files. This variable is defined by Automake. If the makeinfo program is found on the system then it will be used by default; otherwise missing will be used instead.

MAKEINFOHTML

The command invoked to build .html files. Automake defines this to

          $(MAKEINFO) --html.
MAKEINFOFLAGS

User flags passed to each invocation of $(MAKEINFO) and $(MAKEINFOHTML). This user variable (see Section 2.5 [User Variables], page 4) is not expected to be defined in any Makefile; it can be used by users to pass extra flags to suit their needs.

AM_MAKEINFOFLAGS
AM_MAKEINFOHTMLFLAGS

Maintainer flags passed to each makeinfo invocation. These are maintainer variables that can be overridden in Makefile.am. $(AM_MAKEINFOFLAGS) is

          passed to makeinfo when building .info files; and $(AM_MAKEINFOHTMLFLAGS)

is used when building .html files.

For instance the following setting can be used to obtain one single .html file per manual, without node separators.

               AM_MAKEINFOHTMLFLAGS = --no-headers --no-split
          By default, $(AM_MAKEINFOHTMLFLAGS) is set to $(AM_MAKEINFOFLAGS).

This means that defining $(AM_MAKEINFOFLAGS) without defining

          $(AM_MAKEINFOHTMLFLAGS) will impact builds of both .info and .html files.

TEXI2DVI The name of the command that converts a .texi file into a .dvi file. This defaults to texi2dvi, a script that ships with the Texinfo package.

TEXI2PDF The name of the command that translates a .texi file into a .pdf file. This

          defaults to $(TEXI2DVI) --pdf --batch.

DVIPS The name of the command that build a .ps file out of a .dvi file. This defaults

          to dvips.
TEXINFO_TEX

If your package has Texinfo files in many directories, you can use the variable TEXINFO_TEX to tell Automake where to find the canonical texinfo.tex for your package. The value of this variable should be the relative path from the

          current Makefile.am to texinfo.tex:
               TEXINFO_TEX = ../doc/texinfo.tex

11.2 Man pages

A package can also include man pages (but see the GNU standards on this matter, Section “Man Pages” in The GNU Coding Standards.) Man pages are declared using the ‘MANS’ primary. Generally the man_MANS variable is used. Man pages are automatically installed in the correct subdirectory of mandir, based on the file extension.

File extensions such as ‘.1c’ are handled by looking for the valid part of the extension and using that to determine the correct subdirectory of mandir. Valid section names are the digits ‘0’ through ‘9’, and the letters ‘l’ and ‘n’.

Sometimes developers prefer to name a man page something like foo.man in the source, and then rename it to have the correct suffix, e.g. foo.1, when installing the file. Automake also supports this mode. For a valid section named SECTION, there is a corresponding directory named ‘manSECTIONdir’, and a corresponding ‘_MANS’ variable. Files listed in such a variable are installed in the indicated section. If the file already has a valid suffix, then it is installed as-is; otherwise the file suffix is changed to match the section.

For instance, consider this example:

     man1_MANS = rename.man thesame.1 alsothesame.1c

In this case, rename.man will be renamed to rename.1 when installed, but the other files will keep their names.

By default, man pages are installed by ‘make install’. However, since the GNU project does not require man pages, many maintainers do not expend effort to keep the man pages up to date. In these cases, the no-installman option will prevent the man pages from being installed by default. The user can still explicitly install them via ‘make install-man’.

Here is how the man pages are handled in GNU cpio (which includes both Texinfo documentation and man pages):

     man_MANS = cpio.1 mt.1
     EXTRA_DIST = $(man_MANS)

Man pages are not currently considered to be source, because it is not uncommon for man pages to be automatically generated. Therefore they are not automatically included in the distribution. However, this can be changed by use of the ‘dist_’ prefix.

The ‘nobase_’ prefix is meaningless for man pages and is disallowed.

12 What Gets Installed

12.1 Basics of installation

Naturally, Automake handles the details of actually installing your program once it has been built. All files named by the various primaries are automatically installed in the appropriate places when the user runs make install.

A file named in a primary is installed by copying the built file into the appropriate directory. The base name of the file is used when installing.

     bin_PROGRAMS = hello subdir/goodbye

In this example, both ‘hello’ and ‘goodbye’ will be installed in $(bindir).

Sometimes it is useful to avoid the basename step at install time. For instance, you might have a number of header files in subdirectories of the source tree which are laid out precisely how you want to install them. In this situation you can use the ‘nobase_’ prefix to suppress the base name step. For example:

     nobase_include_HEADERS = stdio.h sys/types.h
   Will install stdio.h in $(includedir) and types.h in $(includedir)/sys.

12.2 The two parts of install

Automake generates separate install-data and install-exec rules, in case the installer is installing on multiple machines which share directory structure—these targets allow the machine-independent parts to be installed only once. install-exec installs platformdependent files, and install-data installs platform-independent files. The install target depends on both of these targets. While Automake tries to automatically segregate objects into the correct category, the Makefile.am author is, in the end, responsible for making sure this is done correctly.

Variables using the standard directory prefixes ‘data’, ‘info’, ‘man’, ‘include’,

‘oldinclude’, ‘pkgdata’, or ‘pkginclude’ (e.g.  ‘data_DATA’) are installed by
‘install-data’.

Variables using the standard directory prefixes ‘bin’, ‘sbin’, ‘libexec’, ‘sysconf’,

‘localstate’, ‘lib’, or ‘pkglib’ (e.g. ‘bin_PROGRAMS’) are installed by ‘install-exec’.

Any variable using a user-defined directory prefix with ‘exec’ in the name (e.g. ‘myexecbin_PROGRAMS’ is installed by ‘install-exec’. All other user-defined prefixes are installed by ‘install-data’.

12.3 Extending installation

It is possible to extend this mechanism by defining an install-exec-local or install-data-local rule. If these rules exist, they will be run at ‘make install’ time. These rules can do almost anything; care is required.

Automake also supports two install hooks, install-exec-hook and install-data-hook. These hooks are run after all other install rules of the appropriate type, exec or data, have completed. So, for instance, it is possible to perform post-installation modifications using an install hook.

12.4 Staged installs

Automake generates support for the ‘DESTDIR’ variable in all install rules. ‘DESTDIR’ is used during the ‘make install’ step to relocate install objects into a staging area. Each object and path is prefixed with the value of ‘DESTDIR’ before being copied into the install area. Here is an example of typical DESTDIR usage:

     make DESTDIR=/tmp/staging install

This places install objects in a directory tree built under /tmp/staging. If /gnu/bin/foo and /gnu/share/aclocal/foo.m4 are to be installed, the above command would install

/tmp/staging/gnu/bin/foo and /tmp/staging/gnu/share/aclocal/foo.m4.

This feature is commonly used to build install images and packages. For more information, see Section “Makefile Conventions” in The GNU Coding Standards.

Support for ‘DESTDIR’ is implemented by coding it directly into the install rules. If your Makefile.am uses a local install rule (e.g., install-exec-local) or an install hook, then you must write that code to respect ‘DESTDIR’.

12.5 Rules for the user

Automake also generates rules for targets uninstall, installdirs, and install-strip.

Automake supports uninstall-local and uninstall-hook. There is no notion of separate uninstalls for “exec” and “data”, as these features would not provide additional functionality.

Note that uninstall is not meant as a replacement for a real packaging tool.

13 What Gets Cleaned

The GNU Makefile Standards specify a number of different clean rules. See See Section “Standard Targets for Users” in The GNU Coding Standards.

Generally the files that can be cleaned are determined automatically by Automake. Of course, Automake also recognizes some variables that can be defined to specify additional

files to clean. These variables are MOSTLYCLEANFILES, CLEANFILES, DISTCLEANFILES, and
MAINTAINERCLEANFILES.

As the GNU Standards aren’t always explicit as to which files should be removed by which rule, we’ve adopted a heuristic which we believe was first formulated by François Pinard:

• If make built it, and it is commonly something that one would want to rebuild (for instance, a .o file), then mostlyclean should delete it.

• Otherwise, if make built it, then clean should delete it.

• If configure built it, then distclean should delete it.

• If the maintainer built it (for instance, a .info file), then maintainer-clean should delete it. However maintainer-clean should not delete anything that needs to exist

    in order to run ./configure && make.

We recommend that you follow this same set of heuristics in your Makefile.am.

14 What Goes in a Distribution

14.1 Basics of distribution

The dist rule in the generated Makefile.in can be used to generate a gzip’d tar file and other flavors of archive for distribution. The files is named based on the ‘PACKAGE’ and ‘VERSION’ variables defined by AM_INIT_AUTOMAKE (see Section 5.6 [Macros], page 18); more precisely the gzip’d tar file is named ‘package-version.tar.gz’. You can use the make variable ‘GZIP_ENV’ to control how gzip is run. The default setting is ‘--best’.

For the most part, the files to distribute are automatically found by Automake: all source files are automatically included in a distribution, as are all Makefile.ams and Makefile.ins. Automake also has a built-in list of commonly used files which are automatically included if they are found in the current directory (either physically, or as the target of a Makefile.am rule). This list is printed by ‘automake --help’. Also, files which are read by configure (i.e. the source files corresponding to the files specified in various Autoconf macros such as AC_ CONFIG_FILES and siblings) are automatically distributed. Files included in Makefile.ams (using include) or in configure.ac (using m4_include), and helper scripts installed with

‘automake --add-missing’ are also distributed.

Still, sometimes there are files which must be distributed, but which are not covered in the automatic rules. These files should be listed in the EXTRA_DIST variable. You can mention files from subdirectories in EXTRA_DIST.

You can also mention a directory in EXTRA_DIST; in this case the entire directory will be recursively copied into the distribution. Please note that this will also copy everything in the directory, including CVS/RCS version control files. We recommend against using this feature.

If you define SUBDIRS, Automake will recursively include the subdirectories in the distribution. If SUBDIRS is defined conditionally (see Chapter 20 [Conditionals], page 72), Automake will normally include all directories that could possibly appear in SUBDIRS in the distribution. If you need to specify the set of directories conditionally, you can set the variable DIST_SUBDIRS to the exact list of subdirectories to include in the distribution (see Chapter 6 [Top level], page 24).

14.2 Fine-grained distribution control

Sometimes you need tighter control over what does not go into the distribution; for instance you might have source files which are generated and which you do not want to distribute. In this case Automake gives fine-grained control using the ‘dist’ and ‘nodist’ prefixes. Any primary or ‘_SOURCES’ variable can be prefixed with ‘dist_’ to add the listed files to the distribution. Similarly, ‘nodist_’ can be used to omit the files from the distribution.

As an example, here is how you would cause some data to be distributed while leaving some source code out of the distribution:

     dist_data_DATA = distribute-this
     bin_PROGRAMS = foo
     nodist_foo_SOURCES = do-not-distribute.c

14.3 The dist hook

Occasionally it is useful to be able to change the distribution before it is packaged up. If the dist-hook rule exists, it is run after the distribution directory is filled, but before the actual tar (or shar) file is created. One way to use this is for distributing files in subdirectories for which a new Makefile.am is overkill:

     dist-hook:
             mkdir $(distdir)/random
             cp -p $(srcdir)/random/a1 $(srcdir)/random/a2 $(distdir)/random

Another way to to use this is for removing unnecessary files that get recursively included by specifying a directory in EXTRA DIST:

     EXTRA_DIST = doc
     dist-hook:
     rm -rf `find $(distdir)/doc -name CVS`

Two variables that come handy when writing dist-hook rules are $(distdir) and

$(top_distdir).

$(distdir) points to the directory where the dist rule will copy files from the current directory before creating the tarball. If you are at the top-level directory, then distdir = $(PACKAGE)-$(VERSION). When used from subdirectory named foo/, then distdir =

../$(PACKAGE)-$(VERSION)/foo.

$(top_distdir) always points to the root directory of the distributed tree. At the top-level it’s equal to $(distdir). In the foo/ subdirectory top_distdir =

../$(PACKAGE)-$(VERSION).

Note that when packages are nested using AC_CONFIG_SUBDIRS (see Section “Configuring Other Packages in Subdirectories” in The Autoconf Manual), then $(distdir) and $(top_ distdir) are relative to the package where make dist was run, not to any sub-packages involved.

14.4 Checking the distribution

Automake also generates a distcheck rule which can be of help to ensure that a given distribution will actually work. distcheck makes a distribution, then tries to do a VPATH build, run the test suite, and finally make another tarfile to ensure the distribution is selfcontained.

Building the package involves running ./configure. If you need to supply additional flags to configure, define them in the DISTCHECK_CONFIGURE_FLAGS variable, either in your top-level Makefile.am, or on the command line when invoking make.

If the distcheck-hook rule is defined in your top-level Makefile.am, then it will be invoked by distcheck after the new distribution has been unpacked, but before the unpacked copy is configured and built. Your distcheck-hook can do almost anything, though as always caution is advised. Generally this hook is used to check for potential distribution errors not caught by the standard mechanism. Note that distcheck-hook as well

as DISTCHECK_CONFIGURE_FLAGS are not honored in a subpackage Makefile.am, but the

DISTCHECK_CONFIGURE_FLAGS are passed down to the configure script of the subpackage.

Speaking about potential distribution errors, distcheck will also ensure that the distclean rule actually removes all built files. This is done by running make

distcleancheck at the end of the VPATH build. By default, distcleancheck will

run distclean and then make sure the build tree has been emptied by running $(distcleancheck_listfiles). Usually this check will find generated files that you forgot to add to the DISTCLEANFILES variable (see Chapter 13 [Clean], page 62).

The distcleancheck behavior should be OK for most packages, otherwise you have the possibility to override the definition of either the distcleancheck rule, or

the $(distcleancheck_listfiles) variable. For instance to disable distcleancheck

completely, add the following rule to your top-level Makefile.am:

     distcleancheck:
             @:

If you want distcleancheck to ignore built files which have not been cleaned because they are also part of the distribution, add the following definition instead:

     distcleancheck_listfiles = \
       find -type f -exec sh -c 'test -f $(srcdir)/{} || echo {}' ';'

The above definition is not the default because it’s usually an error if your Makefiles cause some distributed files to be rebuilt when the user build the package. (Think about the user missing the tool required to build the file; or if the required tool is built by your package, consider the cross-compilation case where it can’t be run.) There is a FAQ entry about this (see Section 27.4 [distcleancheck], page 84), make sure you read it before playing

with distcleancheck_listfiles.

distcheck also checks that the uninstall rule works properly, both for ordinary and ‘DESTDIR’ builds. It does this by invoking make uninstall, and then it checks the install tree to see if any files are left over. This check will make sure that you correctly coded your uninstall-related rules.

By default, the checking is done by the distuninstallcheck rule, and the list of files in the install tree is generated by $(distuninstallcheck_listfiles) (this is a variable whose value is a shell command to run that prints the list of files to stdout).

Either of these can be overridden to modify the behavior of distcheck. For instance, to disable this check completely, you would write:

     distuninstallcheck:
             @:

14.5 The types of distributions

Automake generates rules to provide archives of the project for distributions in various formats. Their targets are:

dist-bzip2

Generate a bzip2 tar archive of the distribution. bzip2 archives are frequently smaller than gzipped archives.

dist-gzip

Generate a gzip tar archive of the distribution.

dist-shar

Generate a shar archive of the distribution.

dist-zip Generate a zip archive of the distribution.

dist-tarZ

Generate a compressed tar archive of the distribution.

The rule dist (and its historical synonym dist-all) will create archives in all the enabled formats, Chapter 17 [Options], page 68. By default, only the dist-gzip target is hooked to dist.

15 Support for test suites

Automake supports two forms of test suites.

15.1 Simple Tests

If the variable TESTS is defined, its value is taken to be a list of programs to run in order to do the testing. The programs can either be derived objects or source objects; the generated rule will look both in srcdir and .. Programs needing data files should look for them in srcdir (which is both an environment variable and a make variable) so they work when building in a separate directory (see Section “Build Directories ” in The Autoconf Manual), and in particular for the distcheck rule (see Chapter 14 [Dist], page 62).

The number of failures will be printed at the end of the run. If a given test program exits with a status of 77, then its result is ignored in the final count. This feature allows non-portable tests to be ignored in environments where they don’t make sense.

The variable TESTS_ENVIRONMENT can be used to set environment variables for the test run; the environment variable srcdir is set in the rule. If all your test programs are scripts, you can also set TESTS_ENVIRONMENT to an invocation of the shell (e.g. ‘$(SHELL) -x’); this can be useful for debugging the tests.

You may define the variable XFAIL_TESTS to a list of tests (usually a subset of TESTS) that are expected to fail. This will reverse the result of those tests.

Automake ensures that each program listed in TESTS is built before any tests are run; you can list both source and derived programs in TESTS. For instance, you might want to run a C program as a test. To do this you would list its name in TESTS and also in check_PROGRAMS, and then specify it as you would any other program.

15.2 DejaGnu Tests

If ‘dejagnu’ (ftp://ftp.gnu.org/gnu/dejagnu/) appears in AUTOMAKE_OPTIONS, then a

dejagnu-based test suite is assumed. The variable DEJATOOL is a list of names which are passed, one at a time, as the --tool argument to runtest invocations; it defaults to the name of the package.

The variable RUNTESTDEFAULTFLAGS holds the --tool and --srcdir flags that are passed to dejagnu by default; this can be overridden if necessary.

The variables EXPECT and RUNTEST can also be overridden to provide project-specific values. For instance, you will need to do this if you are testing a compiler toolchain, because the default values do not take into account host and target names.

The contents of the variable RUNTESTFLAGS are passed to the runtest invocation. This is considered a “user variable” (see Section 2.5 [User Variables], page 4). If you need to set

runtest flags in Makefile.am, you can use AM_RUNTESTFLAGS instead.

Automake will generate rules to create a local site.exp file, defining various variables detected by ./configure. This file is automatically read by DejaGnu. It is OK for the user of a package to edit this file in order to tune the test suite. However this is not the place where the test suite author should define new variables: this should be done elsewhere in the real test suite code. Especially, site.exp should not be distributed.

For more information regarding DejaGnu test suites, see See The DejaGnu Manual. In either case, the testing is done via ‘make check’.

15.3 Install Tests

The installcheck target is available to the user as a way to run any tests after the package has been installed. You can add tests to this by writing an installcheck-local rule.

16 Rebuilding Makefiles

Automake generates rules to automatically rebuild Makefiles, configure, and other derived files like Makefile.in.

If you are using AM_MAINTAINER_MODE in configure.ac, then these automatic rebuilding rules are only enabled in maintainer mode.

Sometimes you need to run aclocal with an argument like -I to tell it where to find .m4 files. Since sometimes make will automatically run aclocal, you need a way to specify these arguments. You can do this by defining ACLOCAL_AMFLAGS; this holds arguments which are passed verbatim to aclocal. This variable is only useful in the top-level Makefile.am.

Sometimes it is convenient to supplement the rebuild rules for configure or config.status with additional dependencies. The variables CONFIGURE_DEPENDENCIES and CONFIG_STATUS_DEPENDENCIES can be used to list these extra dependencies. These variable should be defined in all Makefiles of the tree (because these two rebuild rules are output in all them), so it is safer and easier to AC_SUBST them from configure.ac. For instance the following statement will cause configure to be rerun each time version.sh is changed.

     AC_SUBST([CONFIG_STATUS_DEPENDENCIES], ['$(top_srcdir)/version.sh'])

Note the $(top_srcdir)/ in the filename. Since this variable is to be used in all Makefiles, its value must be sensible at any level in the build hierarchy.

   Beware not to mistake CONFIGURE_DEPENDENCIES for CONFIG_STATUS_DEPENDENCIES.

CONFIGURE_DEPENDENCIES adds dependencies to the configure rule, whose effect is to run autoconf. This variable should be seldom used, because automake already tracks m4_ included files. However it can be useful when playing tricky games with m4_esyscmd or similar non-recommendable macros with side effects.

   CONFIG_STATUS_DEPENDENCIES adds dependencies to the config.status rule, whose

effect is to run configure. This variable should therefore carry any non-standard source that may be read as a side effect of running configure, like version.sh in the example above.

Speaking of version.sh scripts, we recommend against them today. They are mainly used when the version of a package is updated automatically by a script (e.g., in daily builds). Here is what some old-style configure.acs may look like:

     AC_INIT
     . $srcdir/version.sh
     AM_INIT_AUTOMAKE([name], $VERSION_NUMBER)
     ...

Here, version.sh is a shell fragment that sets VERSION_NUMBER. The problem with this example is that automake cannot track dependencies (listing version.sh in CONFIG_STATUS_ DEPENDENCIES, and distributing this file is up to the user), and that it uses the obsolete form of AC_INIT and AM_INIT_AUTOMAKE. Upgrading to the new syntax is not straightforward, because shell variables are not allowed in AC_INIT’s arguments. We recommend that version.sh be replaced by an M4 file that is included by configure.ac:

     m4_include([version.m4])
     AC_INIT([name], VERSION_NUMBER)
     AM_INIT_AUTOMAKE
     ...
Here version.m4 could contain something like m4_define([VERSION_NUMBER], [1.2]).

The advantage of this second form is that automake will take care of the dependencies when defining the rebuild rule, and will also distribute the file automatically. An inconvenience is that autoconf will now be rerun each time the version number is bumped, when only configure had to be rerun in the previous setup.

17 Changing Automake’s Behavior

Various features of Automake can be controlled by options in the Makefile.am. Such options are applied on a per-Makefile basis when listed in a special Makefile variable named AUTOMAKE_OPTIONS. They are applied globally to all processed Makefiles when listed in the first argument of AM_INIT_AUTOMAKE in configure.ac. Currently understood options are:

gnits
gnu
foreign
cygnus

Set the strictness as appropriate. The gnits option also implies readme-alpha

          and check-news.
ansi2knr
path/ansi2knr

Turn on automatic de-ANSI-fication. See Section 8.14 [ANSI], page 47. If preceded by a path, the generated Makefile.in will look in the specified directory to find the ansi2knr program. The path should be a relative path to another directory in the same distribution (Automake currently does not check this).

check-news

Cause make dist to fail unless the current version number appears in the first few lines of the NEWS file.

dejagnu Cause dejagnu-specific rules to be generated. See Chapter 15 [Tests], page 65.

dist-bzip2
          Hook dist-bzip2 to dist.
dist-shar
          Hook dist-shar to dist.
dist-zip Hook dist-zip to dist.
dist-tarZ
          Hook dist-tarZ to dist.
no-define

This options is meaningful only when passed as an argument to AM_INIT_ AUTOMAKE. It will prevent the PACKAGE and VERSION variables to be AC_DEFINEd.

no-dependencies

This is similar to using ‘--include-deps’ on the command line, but is useful for those situations where you don’t have the necessary bits to make automatic dependency tracking work See Section 8.15 [Dependencies], page 48. In this case the effect is to effectively disable automatic dependency tracking.

no-dist Don’t emit any code related to dist target. This is useful when a package has its own method for making distributions.

no-dist-gzip
          Do not hook dist-gzip to dist.

no-exeext

If your Makefile.am defines a rule for target ‘foo’, it will override a rule for a target named ‘foo$(EXEEXT)’. This is necessary when EXEEXT is found to be empty. However, by default automake will generate an error for this use. The no-exeext option will disable this error. This is intended for use only where it is known in advance that the package will not be ported to Windows, or any other operating system using extensions on executables.

no-installinfo

The generated Makefile.in will not cause info pages to be built or installed by default. However, info and install-info targets will still be available. This option is disallowed at ‘GNU’ strictness and above.

no-installman

The generated Makefile.in will not cause man pages to be installed by default. However, an install-man target will still be available for optional installation. This option is disallowed at ‘GNU’ strictness and above.

nostdinc This option can be used to disable the standard ‘-I’ options which are ordinarily automatically provided by Automake.

no-texinfo.tex

Don’t require texinfo.tex, even if there are texinfo files in this directory.

readme-alpha

If this release is an alpha release, and the file README-alpha exists, then it will be added to the distribution. If this option is given, version numbers are expected to follow one of two forms. The first form is ‘MAJOR.MINOR.ALPHA’, where each element is a number; the final period and number should be left off for non-alpha releases. The second form is ‘MAJOR.MINORALPHA’, where ALPHA is a letter; it should be omitted for non-alpha releases.

std-options

Make the installcheck rule check that installed scripts and programs support the --help and --version options. This also provides a basic check that the program’s run-time dependencies are satisfied after installation. In a few situations, programs (or scripts) have to be exempted from this test. For instance false (from GNU sh-utils) is never successful, even for --help or --version. You can list such programs in the variable AM_INSTALLCHECK_ STD_OPTIONS_EXEMPT. Programs (not scripts) listed in this variable should be suffixed by $(EXEEXT) for the sake of Win32 or OS/2. For instance suppose we build false as a program but true.sh as a script, and that neither of them

support --help or --version:
     AUTOMAKE_OPTIONS = std-options
     bin_PROGRAMS = false ...
     bin_SCRIPTS = true.sh ...
     AM_INSTALLCHECK_STD_OPTIONS_EXEMPT = false$(EXEEXT) true.sh
subdir-objects

If this option is specified, then objects are placed into the subdirectory of the build directory corresponding to the subdirectory of the source file. For instance if the source file is subdir/file.cxx, then the output file would be

          subdir/file.o.

version A version number (e.g. ‘0.30’) can be specified. If Automake is not newer than the version specified, creation of the Makefile.in will be suppressed.

-Wcategory or --warnings=category

These options behave exactly like their command-line counterpart (see Chapter 4 [Invoking Automake], page 9). This allows you to enable or disable some warning categories on a per-file basis. You can also setup some warnings for your entire project; for instance try AM_INIT_AUTOMAKE([-Wall]) in

          your configure.ac.

Unrecognized options are diagnosed by automake.

If you want an option to apply to all the files in the tree, you can use the AM_INIT_ AUTOMAKE macro in configure.ac. See Section 5.6 [Macros], page 18.

18 Miscellaneous Rules

There are a few rules and variables that didn’t fit anywhere else.

18.1 Interfacing to etags

Automake will generate rules to generate TAGS files for use with GNU Emacs under some circumstances.

If any C, C++ or Fortran 77 source code or headers are present, then tags and TAGS rules will be generated for the directory.

At the topmost directory of a multi-directory package, a tags rule will be output which, when run, will generate a TAGS file that includes by reference all TAGS files from subdirectories.

The tags rule will also be generated if the variable ETAGS_ARGS is defined. This variable is intended for use in directories which contain taggable source that etags does not understand. The user can use the ETAGSFLAGS to pass additional flags to etags; AM_ETAGSFLAGS is also available for use in Makefile.am.

Here is how Automake generates tags for its source, and for nodes in its Texinfo file:

     ETAGS_ARGS = automake.in --lang=none \
      --regex='/^@node[ \t]+\([^,]+\)/\1/' automake.texi

If you add filenames to ‘ETAGS_ARGS’, you will probably also want to set ‘TAGS_DEPENDENCIES’. The contents of this variable are added directly to the dependencies for the tags rule.

Automake also generates a ctags rule which can be used to build vi-style tags files. The variable CTAGS is the name of the program to invoke (by default ‘ctags’); CTAGSFLAGS can be used by the user to pass additional flags, and AM_CTAGSFLAGS can be used by the

Makefile.am.

Automake will also generate an ID rule which will run mkid on the source. This is only supported on a directory-by-directory basis.

Automake also supports the GNU Global Tags program (http://www.gnu.org/ software/global/). The GTAGS rule runs Global Tags automatically and puts the result in the top build directory. The variable GTAGS_ARGS holds arguments which are passed to

gtags.

18.2 Handling new file extensions

It is sometimes useful to introduce a new implicit rule to handle a file type that Automake does not know about.

For instance, suppose you had a compiler which could compile ‘.foo’ files to ‘.o’ files. You would simply define an suffix rule for your language:

     .foo.o:
             foocc -c -o $@ $<

Then you could directly use a ‘.foo’ file in a ‘_SOURCES’ variable and expect the correct results:

     bin_PROGRAMS = doit
     doit_SOURCES = doit.foo

This was the simpler and more common case. In other cases, you will have to help Automake to figure which extensions you are defining your suffix rule for. This usually happens when your extensions does not start with a dot. Then, all you have to do is to put a list of new suffixes in the SUFFIXES variable before you define your implicit rule.

For instance the following definition prevents Automake to misinterpret ‘.idlC.cpp:’ as an attempt to transform ‘.idlC’ into ‘.cpp’.

     SUFFIXES = .idl C.cpp
     .idlC.cpp:
             # whatever

As you may have noted, the SUFFIXES variable behaves like the .SUFFIXES special target of make. You should not touch .SUFFIXES yourself, but use SUFFIXES instead and let Automake generate the suffix list for .SUFFIXES. Any given SUFFIXES go at the start of the generated suffixes list, followed by Automake generated suffixes not already in the list.

18.3 Support for Multilibs

Automake has support for an obscure feature called multilibs. A multilib is a library which is built for multiple different ABIs at a single time; each time the library is built with a different target flag combination. This is only useful when the library is intended to be cross-compiled, and it is almost exclusively used for compiler support libraries.

The multilib support is still experimental. Only use it if you are familiar with multilibs and can debug problems you might encounter.

19 Include

Automake supports an include directive which can be used to include other Makefile fragments when automake is run. Note that these fragments are read and interpreted by automake, not by make. As with conditionals, make has no idea that include is in use.

There are two forms of include:

include $(srcdir)/file

Include a fragment which is found relative to the current source directory.

include $(top_srcdir)/file

Include a fragment which is found relative to the top source directory.

Note that if a fragment is included inside a conditional, then the condition applies to the entire contents of that fragment.

Makefile fragments included this way are always distributed because there are needed to

rebuild Makefile.in.

20 Conditionals

Automake supports a simple type of conditionals.

Before using a conditional, you must define it by using AM_CONDITIONAL in the configure.ac file (see Section 5.6 [Macros], page 18).

[Macro] AM_CONDITIONAL (conditional, condition)

The conditional name, conditional, should be a simple string starting with a letter and containing only letters, digits, and underscores. It must be different from ‘TRUE’ and ‘FALSE’ which are reserved by Automake. The shell condition (suitable for use in a shell if statement) is evaluated when configure is run. Note that you must arrange for every AM_CONDITIONAL to be invoked every time configure is run – if AM_CONDITIONAL is run conditionally (e.g., in a shell if statement), then the result will confuse automake.

Conditionals typically depend upon options which the user provides to the configure script. Here is an example of how to write a conditional which is true if the user uses the

‘--enable-debug’ option.
AC_ARG_ENABLE(debug,
[  --enable-debug    Turn on debugging],
[case "${enableval}" in
  yes) debug=true ;;
  no)  debug=false ;;
  *) AC_MSG_ERROR(bad value ${enableval} for --enable-debug) ;;
esac],[debug=false])
AM_CONDITIONAL(DEBUG, test x$debug = xtrue)

Here is an example of how to use that conditional in Makefile.am:

if DEBUG
DBG = debug
else
DBG =
endif
noinst_PROGRAMS = $(DBG)

This trivial example could also be handled using EXTRA PROGRAMS (see Section 8.1.4 [Conditional Programs], page 29).

You may only test a single variable in an if statement, possibly negated using ‘!’. The else statement may be omitted. Conditionals may be nested to any depth. You may specify an argument to else in which case it must be the negation of the condition used for the current if. Similarly you may specify the condition which is closed by an end:

     if DEBUG
     DBG = debug
     else !DEBUG
     DBG =
     endif !DEBUG

Unbalanced conditions are errors.

Note that conditionals in Automake are not the same as conditionals in GNU Make. Automake conditionals are checked at configure time by the configure script, and affect the translation from Makefile.in to Makefile. They are based on options passed to configure and on results that configure has discovered about the host system. GNU Make conditionals are checked at make time, and are based on variables passed to the make program or defined in the Makefile.

Automake conditionals will work with any make program.

21 The effect of --gnu and --gnits

The ‘--gnu’ option (or ‘gnu’ in the ‘AUTOMAKE_OPTIONS’ variable) causes automake to check the following:

  • The files INSTALL, NEWS, README, AUTHORS, and ChangeLog, plus one of COPYING.LIB,

COPYING.LESSER or COPYING, are required at the topmost directory of the package.

• The options ‘no-installman’ and ‘no-installinfo’ are prohibited.

Note that this option will be extended in the future to do even more checking; it is advisable to be familiar with the precise requirements of the GNU standards. Also, ‘--gnu’ can require certain non-standard GNU programs to exist for use by various maintainer-only rules; for instance in the future pathchk might be required for ‘make dist’.

The ‘--gnits’ option does everything that ‘--gnu’ does, and checks the following as well:

• ‘make installcheck’ will check to make sure that the --help and --version really print a usage message and a version string, respectively. This is the std-options option (see Chapter 17 [Options], page 68).

• ‘make dist’ will check to make sure the NEWS file has been updated to the current version.

• ‘VERSION’ is checked to make sure its format complies with Gnits standards.

• If ‘VERSION’ indicates that this is an alpha release, and the file README-alpha appears in the topmost directory of a package, then it is included in the distribution. This is done in ‘--gnits’ mode, and no other, because this mode is the only one where version number formats are constrained, and hence the only mode where Automake can automatically determine whether README-alpha should be included.

• The file THANKS is required.

22 The effect of --cygnus

Some packages, notably GNU GCC and GNU gdb, have a build environment originally written at Cygnus Support (subsequently renamed Cygnus Solutions, and then later purchased by Red Hat). Packages with this ancestry are sometimes referred to as “Cygnus” trees.

A Cygnus tree has slightly different rules for how a Makefile.in is to be constructed. Passing ‘--cygnus’ to automake will cause any generated Makefile.in to comply with Cygnus rules.

Here are the precise effects of ‘--cygnus’:

• Info files are always created in the build directory, and not in the source directory.

• texinfo.tex is not required if a Texinfo source file is specified. The assumption is that the file will be supplied, but in a place that Automake cannot find. This assumption is an artifact of how Cygnus packages are typically bundled.

• ‘make dist’ is not supported, and the rules for it are not generated. Cygnus-style trees use their own distribution mechanism.

• Certain tools will be searched for in the build tree as well as in the user’s ‘PATH’. These

    tools are runtest, expect, makeinfo and texi2dvi.

• --foreign is implied.

  • The options ‘no-installinfo’ and ‘no-dependencies’ are implied.
  • The macros ‘AM_MAINTAINER_MODE’ and ‘AM_CYGWIN32’ are required.

• The check target doesn’t depend on all.

GNU maintainers are advised to use ‘gnu’ strictness in preference to the special Cygnus mode. Some day, perhaps, the differences between Cygnus trees and GNU trees will disappear (for instance, as GCC is made more standards compliant). At that time the special Cygnus mode will be removed.

23 When Automake Isn’t Enough

In some situations, where Automake is not up to one task, one has to resort to handwritten rules or even handwritten Makefiles.

23.1 Extending Automake Rules

With some minor exceptions (like _PROGRAMS variables being rewritten to append $(EXEEXT)), the contents of a Makefile.am is copied to Makefile.in verbatim.

These copying semantics means that many problems can be worked around by simply adding some make variables and rules to Makefile.am. Automake will ignore these additions.

Since a Makefile.in is built from data gathered from three different places (Makefile.am, configure.ac, and automake itself), it is possible to have conflicting definitions of rules or variables. When building Makefile.in the following priorities are respected by automake to ensure the user always have the last word. User defined variables in Makefile.am have priority over variables AC_SUBSTed from configure.ac, and AC_SUBSTed variables have priority over automake-defined variables. As far rules are concerned, a user-defined rule overrides any automake-defined rule for the same target.

These overriding semantics make it possible to fine tune some default settings of Automake, or replace some of its rules. Overriding Automake rules is often inadvisable, particularly in the topmost directory of a package with subdirectories. The -Woverride option (see Chapter 4 [Invoking Automake], page 9) comes handy to catch overridden definitions.

Note that Automake does not make any difference between rules with commands and rules that only specify dependencies. So it is not possible to append new dependencies to an automake-defined target without redefining the entire rule.

However, various useful targets have a ‘-local’ version you can specify in your Makefile.in. Automake will supplement the standard target with these user-supplied targets.

The targets that support a local version are all, info, dvi, ps, pdf, html, check,

install-data, install-exec, uninstall, installdirs, installcheck and the various
clean targets (mostlyclean, clean, distclean, and maintainer-clean). Note that
there are no uninstall-exec-local or uninstall-data-local targets; just use

uninstall-local. It doesn’t make sense to uninstall just data or just executables.

For instance, here is one way to install a file in /etc:

     install-data-local:
             $(INSTALL_DATA) $(srcdir)/afile $(DESTDIR)/etc/afile

Some rule also have a way to run another rule, called a hook, after their work is done. The hook is named after the principal target, with ‘-hook’ appended. The targets allowing

hooks are install-data, install-exec, uninstall, dist, and distcheck.

For instance, here is how to create a hard link to an installed program:

     install-exec-hook:
             ln $(DESTDIR)$(bindir)/program$(EXEEXT) \
                $(DESTDIR)$(bindir)/proglink$(EXEEXT)

Although cheaper and more portable than symbolic links, hard links will not work everywhere (for instance OS/2 does not have ln). Ideally you should fall back to cp -p when ln does not work. An easy way, if symbolic links are acceptable to you, is to add AC_PROG_LN_S to configure.ac (see Section “Particular Program Checks” in The Autoconf Manual) and

use $(LN_S) in Makefile.am.

For instance, here is how you could install a versioned copy of a program using $(LN_S):

     install-exec-hook:
             cd $(DESTDIR)$(bindir) && \
               mv -f prog$(EXEEXT) prog-$(VERSION)$(EXEEXT) && \
               $(LN_S) prog-$(VERSION)$(EXEEXT) prog$(EXEEXT)

Note that we rename the program so that a new version will erase the symbolic link, not the real binary. Also we cd into the destination directory in order to create relative links.

23.2 Third-Party Makefiles

In most projects all Makefiles are generated by Automake. In some cases, however, projects need to embed subdirectories with handwritten Makefiles. For instance one subdirectory could be a third-party project with its own build system, not using Automake.

It is possible to list arbitrary directories in SUBDIRS or DIST_SUBDIRS provided each of these directories has a Makefile that recognizes all the following recursive targets.

When a user runs one of these targets, that target is run recursively in all subdirectories. This is why it is important that even third-party Makefiles support them.

all Compile the entire package. This is the default target in Automake-generated Makefiles, but it does not need to be the default in third-party Makefiles.

distdir Copy files to distribute into $(distdir), before a tarball is constructed. Of course this target is not required if the no-dist option (see Chapter 17 [Options], page 68) is used. The variables $(top_distdir) and $(distdir) (see Chapter 14 [Dist], page 62) will be passed from the outer package to the subpackage when the distdir target is invoked. These two variables have been adjusted for the directory which is being recursed into, so they are ready to use.

install
install-data
install-exec
uninstall

Install or uninstall files (see Chapter 12 [Install], page 60).

install-info

Install only the Texinfo documentation (see Section 11.1 [Texinfo], page 57).

installdirs

Create install directories, but do not install any files.

check
installcheck

Check the package (see Chapter 15 [Tests], page 65).

mostlyclean
clean
distclean
maintainer-clean

Cleaning rules (see Chapter 13 [Clean], page 62).

dvi
pdf
ps
info

html Build the documentation in various formats (see Section 11.1 [Texinfo], page 57).

tags

ctags Build TAGS and CTAGS (see Section 18.1 [Tags], page 70).

If you have ever used Gettext in a project, this is a good example of how third-party Makefiles can be used with Automake. The Makefiles gettextize puts in the po/ and intl/ directories are handwritten Makefiles that implement all these targets. That way they can be added to SUBDIRS in Automake packages.

Directories which are only listed in DIST_SUBDIRS but not in SUBDIRS need only the distclean, maintainer-clean, and distdir rules (see Chapter 6 [Top level], page 24).

Usually, many of these rules are irrelevant to the third-party subproject, but they are required for the whole package to work. It’s OK to have a rule that does nothing, so if you are integrating a third-party project with no documentation or tag support, you could simply augment its Makefile as follows:

     EMPTY_AUTOMAKE_TARGETS = dvi pdf ps info html tags ctags
     .PHONY: $(EMPTY_AUTOMAKE_TARGETS)
     $(EMPTY_AUTOMAKE_TARGETS):

Another aspect of integrating third-party build systems is whether they support VPATH builds. Obviously if the subpackage does not support VPATH builds the whole package will not support VPATH builds. This in turns means that make distcheck will not work, because it relies on VPATH builds. Some people can live without this (actually, many Automake users have never heard of make distcheck). Other people may prefer to revamp the existing Makefiles to support VPATH. Doing so does not necessarily require Automake, only Autoconf is needed (see Section “Build Directories” in The Autoconf Manual). The necessary substitutions: @scrdir@, @top_srcdir@, and @top_buildir@ are defined by configure when it processes a Makefile (see Section “Preset Output Variables” in The Autoconf Manual), they are not computed by the Makefile like the aforementioned

$(distdir) and $(top_distdir) variables..

It is sometimes inconvenient to modify a third-party Makefile to introduce the above required targets. For instance one may want to keep the third-party sources untouched to ease upgrades to new versions.

Here are two other ideas. If GNU make is assumed, one possibility is to add to that subdirectory a GNUmakefile that defines the required targets and include the third-party Makefile. For this to work in VPATH builds, GNUmakefile must lie in the build directory; the easiest way to do this is to write a GNUmakefile.in instead, and have it processed with AC_CONFIG_FILES from the outer package. For example if we assume Makefile defines all targets except the documentation targets, and that the check target is actually called test,

we could write GNUmakefile (or GNUmakefile.in) like this:
     # First, include the real Makefile
     include Makefile
     # Then, define the other targets needed by Automake Makefiles.
     .PHONY: dvi pdf ps info html check
     dvi pdf ps info html:
     check: test

A similar idea that does not use include is to write a proxy Makefile that dispatches

rules to the real Makefile, either with $(MAKE) -f Makefile.real $(AM_MAKEFLAGS)

target (if it’s OK to rename the original Makefile) or with cd subdir && $(MAKE) $(AM_MAKEFLAGS) target (if it’s OK to store the subdirectory project one directory deeper). The good news is that this proxy Makefile can be generated with Automake.

All we need are -local targets (see Section 23.1 [Extending], page 74) that perform the dispatch. Of course the other Automake features are available, so you could decide to let Automake perform distribution or installation. Here is a possible Makefile.am:

     all-local:
             cd subdir && $(MAKE) $(AM_MAKEFLAGS) all
     check-local:
             cd subdir && $(MAKE) $(AM_MAKEFLAGS) test
     clean-local:
             cd subdir && $(MAKE) $(AM_MAKEFLAGS) clean
     # Assuming the package knows how to install itself
     install-data-local:
             cd subdir && $(MAKE) $(AM_MAKEFLAGS) install-data
     install-exec-local:
             cd subdir && $(MAKE) $(AM_MAKEFLAGS) install-exec
     uninstall-local:
             cd subdir && $(MAKE) $(AM_MAKEFLAGS) uninstall
     # Distribute files from here.
     EXTRA_DIST = subdir/Makefile subdir/program.c ...

Pushing this idea to the extreme, it is also possible to ignore the subproject build system and build everything from this proxy Makefile.am. This might sounds very sensible if you need VPATH builds but the subproject does not support them.

24 Distributing Makefile.ins

Automake places no restrictions on the distribution of the resulting Makefile.ins. We still encourage software authors to distribute their work under terms like those of the GPL, but doing so is not required to use Automake.

Some of the files that can be automatically installed via the --add-missing switch do fall under the GPL. However, these also have a special exception allowing you to distribute them with your package, regardless of the licensing you choose.

25 Automake API versioning

New Automake releases usually include bug fixes and new features. Unfortunately they may also introduce new bugs and incompatibilities. This makes four reasons why a package may require a particular Automake version.

Things get worse when maintaining a large tree of packages, each one requiring a different version of Automake. In the past, this meant that any developer (and sometime users) had to install several versions of Automake in different places, and switch ‘$PATH’ appropriately for each package.

Starting with version 1.6, Automake installs versioned binaries. This means you can install several versions of Automake in the same ‘$prefix’, and can select an arbitrary Automake version by running ‘automake-1.6’ or ‘automake-1.7’ without juggling with ‘$PATH’. Furthermore, Makefile’s generated by Automake 1.6 will use ‘automake-1.6’ explicitly in their rebuild rules.

The number ‘1.6’ in ‘automake-1.6’ is Automake’s API version, not Automake’s version. If a bug fix release is made, for instance Automake 1.6.1, the API version will remain 1.6. This means that a package which work with Automake 1.6 should also work with 1.6.1; after all, this is what people expect from bug fix releases.

If your package relies on a feature or a bug fix introduced in a release, you can pass this version as an option to Automake to ensure older releases will not be used. For instance, use this in your configure.ac:

       AM_INIT_AUTOMAKE(1.6.1)    dnl Require Automake 1.6.1 or better.

or, in a particular Makefile.am:

       AUTOMAKE_OPTIONS = 1.6.1   # Require Automake 1.6.1 or better.

Automake will print an error message if its version is older than the requested version.

What is in the API

Automake’s programming interface is not easy to define. Basically it should include at least all documented variables and targets that a ‘Makefile.am’ author can use, any behavior associated with them (e.g. the places where ‘-hook’’s are run), the command line interface

of ‘automake’ and ‘aclocal’, . . .

What is not in the API

Every undocumented variable, target, or command line option, is not part of the API. You should avoid using them, as they could change from one version to the other (even in bug fix releases, if this helps to fix a bug).

If it turns out you need to use such a undocumented feature, contact automake@gnu. org and try to get it documented and exercised by the test-suite.

26 Upgrading a Package to a Newer Automake Version

Automake maintains three kind of files in a package.

  • aclocal.m4
  • Makefile.ins

• auxiliary tools like install-sh or py-compile

aclocal.m4 is generated by aclocal and contains some Automake-supplied M4 macros. Auxiliary tools are installed by ‘automake --add-missing’ when needed. Makefile.ins are built from Makefile.am by automake, and rely on the definitions of the M4 macros put in aclocal.m4 as well as the behavior of the auxiliary tools installed.

Because all these files are closely related, it is important to regenerate all of them when upgrading to a newer Automake release. The usual way to do that is

     aclocal # with any option needed (such a -I m4)

     autoconf
     automake --add-missing --force-missing

or more conveniently:

     autoreconf -vfi

The use of --force-missing ensures that auxiliary tools will be overridden by new versions (see Chapter 4 [Invoking Automake], page 9).

It is important to regenerate all these files each time Automake is upgraded, even between bug fixes releases. For instance it is not unusual for a bug fix to involve changes to both the rules generated in Makefile.in and the supporting M4 macros copied to aclocal.m4.

Presently automake is able to diagnose situations where aclocal.m4 has been generated with another version of aclocal. However it never checks whether auxiliary scripts are up-to-date. In other words, automake will tell you when aclocal needs to be rerun, but it will never diagnose a missing --force-missing.

Before upgrading to a new major release, it is a good idea to read the file NEWS. This file lists all changes between releases: new features, obsolete constructs, known incompatibilities, and workarounds.

27 Frequently Asked Questions about Automake

This chapter covers some questions that often come up on the mailing lists.

27.1 CVS and generated files

27.1.1 Background: distributed generated files

Packages made with Autoconf and Automake ship with some generated files like configure or Makefile.in. These files were generated on the developer’s host and are distributed so that end-users do not have to install the maintainer tools required to rebuild them. Other generated files like Lex scanners, Yacc parsers, or Info documentation, are usually distributed on similar grounds.

Automake outputs rules in Makefiles to rebuild these files. For instance make will run autoconf to rebuild configure whenever configure.ac is changed. This makes development safer by ensuring a configure is never out-of-date with respect to configure.ac.

As generated files shipped in packages are up-to-date, and because tar preserves timestamps, these rebuild rules are not triggered when a user unpacks and builds a package.

27.1.2 Background: CVS and timestamps

Unless you use CVS keywords (in which case files must be updated at commit time), CVS preserves timestamp during cvs commit and cvs import -d operations.

When you check out a file using cvs checkout its timestamp is set to that of the revision which is being checked out.

However, during cvs update, files will have the date of the update, not the original timestamp of this revision. This is meant to make sure that make notices sources files have been updated.

This times tamp shift is troublesome when both sources and generated files are kept under CVS. Because CVS processes files in alphabetical order, configure.ac will appear older than configure after a cvs update that updates both files, even if configure was newer than configure.ac when it was checked in. Calling make will then trigger a spurious rebuild of configure.

27.1.3 Living with CVS in Autoconfiscated projects

There are basically two clans amongst maintainers: those who keep all distributed files under CVS, including generated files, and those who keep generated files out of CVS.

All files in CVS

• The CVS repository contains all distributed files so you know exactly what is distributed, and you can checkout any prior version entirely.

• Maintainers can see how generated files evolve (for instance you can see what happens to your Makefile.ins when you upgrade Automake and make sure they look OK).

• Users do not need the autotools to build a checkout of the project, it works just like a released tarball.

• If users use cvs update to update their copy, instead of cvs checkout to fetch a fresh one, timestamps will be inaccurate. Some rebuild rules will be triggered and attempt to run developer tools such as autoconf or automake.

Actually, calls to such tools are all wrapped into a call to the missing script discussed later (see Section 27.2 [maintainer-mode], page 82). missing will take care of fixing the timestamps when these tools are not installed, so that the build can continue.

• In distributed development, developers are likely to have different version of the maintainer tools installed. In this case rebuilds triggered by timestamp lossage will lead to spurious changes to generated files. There are several solutions to this:

• All developers should use the same versions, so that the rebuilt files are identical to files in CVS. (This starts to be difficult when each project you work on uses different versions.)

• Or people use a script to fix the timestamp after a checkout (the GCC folks have such a script).

• Or configure.ac uses AM_MAINTAINER_MODE, which will disable all these rebuild rules by default. This is further discussed in Section 27.2 [maintainer-mode], page 82.

• Although we focused on spurious rebuilds, the converse can also happen. CVS’s timestamp handling can also let you think an out-of-date file is up-to-date.

For instance, suppose a developer has modified Makefile.am and rebuilt Makefile.in, and then decide to do a last-minute change to Makefile.am right before checking in both files (without rebuilding Makefile.in to account for the change).

This last change to Makefile.am make the copy of Makefile.in out-of-date. Since CVS processes files alphabetically, when another developer cvs update his or her tree, Makefile.in will happen to be newer than Makefile.am. This other developer will not see Makefile.in is out-of-date.

Generated files out of CVS

One way to get CVS and make working peacefully is to never store generated files in CVS, i.e., do not CVS-control files which are Makefile targets (also called derived files).

This way developers are not annoyed by changes to generated files. It does not matter if they all have different versions (assuming they are compatible, of course). And finally, timestamps are not lost, changes to sources files can’t be missed as in the Makefile.am/Makefile.in example discussed earlier.

The drawback is that the CVS repository is not an exact copy of what is distributed and that users now need to install various development tools (maybe even specific versions) before they can build a checkout. But, after all, CVS’s job is versioning, not distribution.

Allowing developers to use different versions of their tools can also hide bugs during distributed development. Indeed, developers will be using (hence testing) their own generated files, instead of the generated files that will be released actually. The developer who prepares the tarball might be using a version of the tool that produces bogus output (for instance a non-portable C file), something other developers could have noticed if they weren’t using their own versions of this tool.

27.1.4 Third-party files

Another class of files not discussed here (because they do not cause timestamp issues) are files which are shipped with a package, but maintained elsewhere. For instance tools like gettextize and autopoint (from Gettext) or libtoolize (from Libtool), will install or update files in your package.

These files, whether they are kept under CVS or not, raise similar concerns about version mismatch between developers’ tools. The Gettext manual has a section about this, see Section “Integrating with CVS” in GNU gettext tools.

27.2 missing and AM_MAINTAINER_MODE
27.2.1 missing

The missing script is a wrapper around several maintainer tools, designed to warn users if a maintainer tool is required but missing. Typical maintainer tools are autoconf, automake, bison, etc. Because file generated by these tools are shipped with the other sources of a package, these tools shouldn’t be required during a user build and they are not checked for

in configure.

However, if for some reason a rebuild rule is triggered and involves a missing tool, missing will notice it and warn the user. Besides the warning, when a tool is missing, missing will attempt to fix timestamps in a way which allow the build to continue. For instance missing will touch configure if autoconf is not installed. When all distributed files are kept under CVS, this feature of missing allows user with no maintainer tools to build a package off CVS, bypassing any timestamp inconsistency implied by cvs update.

If the required tool is installed, missing will run it and won’t attempt to continue after failures. This is correct during development: developers love fixing failures. However, users with wrong versions of maintainer tools may get an error when the rebuild rule is spuriously triggered, halting the build. This failure to let the build continue is one of the arguments

of the AM_MAINTAINER_MODE advocates.

27.2.2 AM_MAINTAINER_MODE

AM_MAINTAINER_MODE disables the so called "rebuild rules" by default. If you have AM_

MAINTAINER_MODE in configure.ac, and run ./configure && make, then make will *never*

attempt to rebuilt configure, Makefile.ins, Lex or Yacc outputs, etc. I.e., this disables build rules for files which are usually distributed and that users should normally not have to update.

If you run ./configure --enable-maintainer-mode, then these rebuild rules will be active.

People use AM_MAINTAINER_MODE either because they do want their users (or themselves) annoyed by timestamps lossage (see Section 27.1 [CVS], page 80), or because they simply can’t stand the rebuild rules and prefer running maintainer tools explicitly.

AM_MAINTAINER_MODE also allows you to disable some custom build rules conditionally. Some developers use this feature to disable rules that need exotic tools that users may not have available.

Several years ago François Pinard pointed out several arguments against AM_ MAINTAINER_MODE. Most of them relate to insecurity. By removing dependencies you get non-dependable builds: change to sources files can have no effect on generated files and this can be very confusing when unnoticed. He adds that security shouldn’t be reserved to maintainers (what --enable-maintainer-mode suggests), on the contrary. If one user has to modify a Makefile.am, then either Makefile.in should be updated or a warning should be output (this is what Automake uses missing for) but the last thing you want is that nothing happens and the user doesn’t notice it (this is what happens when rebuild rules are disabled by AM_MAINTAINER_MODE).

Jim Meyering, the inventor of the AM_MAINTAINER_MODE macro was swayed by François’s arguments, and got rid of AM_MAINTAINER_MODE in all of his packages.

Still many people continue to use AM_MAINTAINER_MODE, because it helps them working on projects where all files are kept under CVS, and because missing isn’t enough if you have the wrong version of the tools.

27.3 Why doesn’t Automake support wildcards?

Developers are lazy. They often would like to use wildcards in Makefile.ams, so they don’t need to remember they have to update Makefile.ams every time they add, delete, or rename a file.

There are several objections to this:

• When using CVS (or similar) developers need to remember they have to run cvs add or cvs rm anyway. Updating Makefile.am accordingly quickly becomes a reflex. Conversely, if your application doesn’t compile because you forgot to add a file in Makefile.am, it will help you remember to cvs add it.

• Using wildcards makes easy to distribute files by mistake. For instance some code a developer is experimenting with (a test case, say) but which should not be part of the distribution.

• Using wildcards it’s easy to omit some files by mistake. For instance one developer creates a new file, uses it at many places, but forget to commit it. Another developer then checkout the incomplete project and is able to run ‘make dist’ successfully, even though a file is missing.

• Listing files, you control *exactly* what you distribute. If some file that should be distributed is missing from your tree, make dist will complain. Besides, you don’t distribute more than what you listed.

• Finally it’s really hard to forget adding a file to Makefile.am, because if you don’t add it, it doesn’t get compiled nor installed, so you can’t even test it.

Still, these are philosophical objections, and as such you may disagree, or find enough value in wildcards to dismiss all of them. Before you start writing a patch against Automake to teach it about wildcards, let’s see the main technical issue: portability.

Although $(wildcard ...) works with GNU make, it is not portable to other make implementations.

The only way Automake could support $(wildcard ...) is by expending $(wildcard ...) when automake is run. Resulting Makefile.ins would be portable since they would list all files and not use $(wildcard ...). However that means developers need to remember they must run automake each time they add, delete, or rename files.

Compared to editing Makefile.am, this is really little win. Sure, it’s easier and faster

to type automake; make than to type emacs Makefile.am; make. But nobody bothered

enough to write a patch add support for this syntax. Some people use scripts to generated file lists in Makefile.am or in separate Makefile fragments.

Even if you don’t care about portability, and are tempted to use $(wildcard ...) anyway because you target only GNU Make, you should know there are many places where Automake need to know exactly which files should be processed. As Automake doesn’t know how to expand $(wildcard ...), you cannot use it in these places. $(wildcard ...) is a black box comparable to AC_SUBSTed variables as far Automake is concerned.

You can get warnings about $(wildcard ...) constructs using the -Wportability flag.

27.4 Files left in build directory after distclean

This is a diagnostic you might encounter while running make distcheck.

As explained in Chapter 14 [Dist], page 62, make distcheck attempts to build and check your package for errors like this one.

make distcheck will perform a VPATH build of your package, and then call make distclean. Files left in the build directory after make distclean has run are listed after this error.

This diagnostic really covers two kinds of errors:

• files that are forgotten by distclean;

• distributed files that are erroneously rebuilt.

The former left-over files are not distributed, so the fix is to mark them for cleaning (see Chapter 13 [Clean], page 62), this is obvious and doesn’t deserve more explanations.

The latter bug is not always easy to understand and fix, so let’s proceed with an example. Suppose our package contains a program for which we want to build a man page using help2man. GNU help2man produces simple manual pages from the --help and --version output of other commands (see Section “Overview” in The Help2man Manual). Because we don’t to force want our users to install help2man, we decide to distribute the generated man page using the following setup.

     # This Makefile.am is bogus.
     bin_PROGRAMS = foo
     foo_SOURCES = foo.c
     dist_man_MANS = foo.1
     foo.1: foo$(EXEEXT)
     help2man --output=foo.1 ./foo$(EXEEXT)

This will effectively distribute the man page. However, make distcheck will fail with:

     ERROR: files left in build directory after distclean:
     ./foo.1

Why was foo.1 rebuilt? Because although distributed, foo.1 depends on a nondistributed built file: foo$(EXEEXT). foo$(EXEEXT) is built by the user, so it will always appear to be newer than the distributed foo.1.

make distcheck caught an inconsistency in our package. Our intent was to distribute foo.1 so users do not need installing help2man, however since this our rule causes this file to be always rebuilt, users do need help2man. Either we should ensure that foo.1 is not rebuilt by users, or there is no point in distributing foo.1.

More generally, the rule is that distributed files should never depend on non-distributed built files. If you distribute something generated, distribute its sources.

One way to fix the above example, while still distributing foo.1 is to not depend on foo$(EXEEXT). For instance, assuming foo --version and foo --help do not change unless foo.c or configure.ac change, we could write the following Makefile.am:

     bin_PROGRAMS = foo
     foo_SOURCES = foo.c
     dist_man_MANS = foo.1
     foo.1: foo.c $(top_srcdir)/configure.ac
             $(MAKE) $(AM_MAKEFLAGS) foo$(EXEEXT)
     help2man --output=foo.1 ./foo$(EXEEXT)

This way, foo.1 will not get rebuilt every time foo$(EXEEXT) changes. The make call makes sure foo$(EXEEXT) is up-to-date before help2man. Another way to ensure this would be to use separate directories for binaries and man pages, and set SUBDIRS so that binaries are built before man pages.

We could also decide not to distribute foo.1. In this case it’s fine to have foo.1 dependent upon foo$(EXEEXT), since both will have to be rebuilt. However it would be impossible to build the package in a cross-compilation, because building foo.1 involves an

execution of foo$(EXEEXT).

Another context where such errors are common is when distributed files are built by tools which are built by the package. The pattern is similar:

     distributed-file: built-tools distributed-sources
             build-command

should be changed to

     distributed-file: distributed-sources
             $(MAKE) $(AM_MAKEFLAGS) built-tools
             build-command

or you could choose not to distribute distributed-file, if cross-compilation does not matter.

The points made through these examples are worth a summary:

• Distributed files should never depend upon non-distributed built files.

• Distributed files should be distributed will all their dependencies.

• If a file is intended be rebuilt by users, there is no point in distributing it.

For desperate cases, it’s always possible to disable this check by setting distcleancheck_ listfiles as documented in Chapter 14 [Dist], page 62. Make sure you do understand the

reason why make distcheck complains before you do this. distcleancheck_listfiles is

a way to hide errors, not to fix them. You can always do better.

27.5 Why are object files sometimes renamed?

This happens when per-target compilation flags are used. Object files need to be renamed just in case they would clash with object files compiled from the same sources, but with different flags. Consider the following example.

     bin_PROGRAMS = true false
     true_SOURCES = generic.c
     true_CPPFLAGS = -DEXIT_CODE=0
     false_SOURCES = generic.c
     false_CPPFLAGS = -DEXIT_CODE=1

Obviously the two programs are built from the same source, but it would be bad if they shared the same object, because generic.o cannot be built with both -DEXIT_CODE=0 *and* -DEXIT_CODE=1. Therefore automake outputs rules to build two different objects:

true-generic.o and false-generic.o.

automake doesn’t actually look whether sources files are shared to decide if it must rename objects. It will just rename all objects of a target as soon as it sees per-target compilation flags are used.

It’s OK to share object files when per-target compilation flags are not used. For instance true and false will both use version.o in the following example.

     AM_CPPFLAGS = -DVERSION=1.0
     bin_PROGRAMS = true false
     true_SOURCES = true.c version.c
     false_SOURCES = false.c version.c

Note that the renaming of objects is also affected by the _SHORTNAME variable (see Section 8.4 [Program and Library Variables], page 35).

27.6 Handling Tools that Produce Many Outputs

This section describes a make idiom that can be used when a tool produces multiple output files. It is not specific to Automake and can be used in ordinary Makefiles.

Suppose we have a program called foo that will read one file called data.foo and produce two files named data.c and data.h. We want to write a Makefile rule that captures this one-to-two dependency.

The naive rule is incorrect:

     # This is incorrect.
     data.c data.h: data.foo
             foo data.foo

What the above rule really says is that data.c and data.h each depend on data.foo, and can each be built by running foo data.foo. In other words it is equivalent to:

     # We do not want this.
     data.c: data.foo
             foo data.foo
     data.h: data.foo
             foo data.foo

which means that foo can be run twice. Usually it will not be run twice, because make implementations are smart enough to check for the existence of the second file after the first one has been built; they will therefore detect that it already exists. However there are a few situations where it can run twice anyway:

• The most worrying case is when running a parallel make. If data.c and data.h are built in parallel, two foo data.foo commands will run concurrently. This is harmful.

• Another case is when the dependency (here data.foo) is (or depends upon) a phony target.

A solution that works with parallel make but not with phony dependencies is the following:

     data.c data.h: data.foo
             foo data.foo
     data.h: data.c

The above rules are equivalent to

     data.c: data.foo
             foo data.foo
     data.h: data.foo data.c
             foo data.foo

therefore a parallel make will have to serialize the builds of data.c and data.h, and will detect that the second is no longer needed once the first is over.

Using this pattern is probably enough for most cases. However it does not scale easily to more output files (in this scheme all output files must be totally ordered by the dependency relation), so we will explore a more complicated solution.

Another idea is to write the following:

     # There is still a problem with this one.
     data.c: data.foo

             foo data.foo
     data.h: data.c

The idea is that foo data.foo is run only when data.c needs to be updated, but we further state that data.h depends upon data.c. That way, if data.h is required and data.foo is out of date, the dependency on data.c will trigger the build.

This is almost perfect, but suppose we have built data.h and data.c, and then we erase data.h. Then, running make data.h will not rebuild data.h. The above rules just state that data.c must be up-to-date with respect to data.foo, and this is already the case.

What we need is a rule that forces a rebuild when data.h is missing. Here it is:

     data.c: data.foo
             foo data.foo
     data.h: data.c
             @if test -f $@; then :; else \
               rm -f data.c; \
               $(MAKE) $(AM_MAKEFLAGS) data.c; \
             fi

The above scales easily to more outputs and more inputs. One of the output is picked up to serve as a witness of the run of the command, it depends upon all inputs, and all other outputs depend upon it. For instance if foo should additionally read data.bar and also produce data.w and data.x, we would write:

     data.c: data.foo data.bar
             foo data.foo data.bar
     data.h data.w data.x: data.c
             @if test -f $@; then :; else \
               rm -f data.c; \
               $(MAKE) $(AM_MAKEFLAGS) data.c; \
             fi

There is still a minor problem with this setup. foo outputs four files, but we do not know in which order these files are created. Suppose that data.h is created before data.c. Then we have a weird situation. The next time make is run, data.h will appear older than data.c, the second rule will be triggered, a shell will be started to execute the if...fi command, but actually it will just execute the then branch, that is: nothing. In other words, because the witness we selected is not the first file created by foo, make will start a shell to do nothing each time it is run.

A simple riposte is to fix the timestamps when this happens.

     data.c: data.foo data.bar
             foo data.foo data.bar
     data.h data.w data.x: data.c
             @if test -f $@; then \
               touch $@; \
             else \
               rm -f data.c; \
               $(MAKE) $(AM_MAKEFLAGS) data.c; \
             fi

Another solution, not incompatible with the previous one, is to use a different and dedicated file as witness, rather than using any of foo’s outputs.

data.stamp: data.foo data.bar
@rm -f data.tmp
@touch data.tmp
foo data.foo data.bar
@mv -f data.tmp $@
data.c data.h data.w data.x: data.stamp
@if test -f $@; then \
  touch $@; \
else \
  rm -f data.stamp; \
  $(MAKE) $(AM_MAKEFLAGS) data.stamp; \
fi

data.tmp is created before foo is run, so it has a timestamp older than output files output by foo. It is then renamed to data.stamp after foo has run, because we do not want to update data.stamp if foo fails.

Using a dedicated witness like this is very handy when the list of output files is not known beforehand. As an illustration, consider the following rules to compile many *.el files into *.elc files in a single command. It does not matter how ELFILES is defined (as long as it is not empty: empty targets are not accepted by POSIX).

ELFILES = one.el two.el three.el ...
ELCFILES = $(ELFILES:=c)
elc-stamp: $(ELFILES)
@rm -f elc-temp
@touch elc-temp
$(elisp_comp) $(ELFILES)
@mv -f elc-temp $@
$(ELCFILES): elc-stamp
@if test -f $@; then \
  touch $@; \
else \
  rm -f elc-stamp; \
  $(MAKE) $(AM_MAKEFLAGS) elc-stamp; \
fi

For completeness it should be noted that GNU make is able to express rules with multiple output files using pattern rules (see Section “Pattern Rule Examples” in The GNU Make Manual). We do not discuss pattern rules here because they are not portable, but they can be convenient in packages that assume GNU make.

Appendix A Copying This Manual

A.1 GNU Free Documentation License

Version 1.2, November 2002

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000,2001,2002 Free Software Foundation, Inc. 59 Temple Place, Suite 330, Boston, MA 02111-1307, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them.

The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition.

The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warrany Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided for under this License. Any other attempt to copy, modify, sublicense or distribute the Document is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation.

A.1.1 ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.2
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover Texts.
        A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with...Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Appendix B: Indices 97

Appendix B Indices

B.1 Macro and Variable Index

_LDADD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 _LDFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 _LIBADD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 _SOURCES. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27, 38 _TEXINFOS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

A

AC_CANONICAL_HOST . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 AC_CANONICAL_SYSTEM. . . . . . . . . . . . . . . . . . . . . . . . . . 13 AC_CONFIG_AUX_DIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 AC_CONFIG_FILES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 AC_CONFIG_HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 AC_CONFIG_LINKS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 AC_DEFUN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 AC_F77_LIBRARY_LDFLAGS . . . . . . . . . . . . . . . . . . . . . . 14 AC_LIBOBJ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13, 34 AC_LIBSOURCE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 AC_LIBSOURCES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 AC_OUTPUT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 AC_PREREQ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 AC_PROG_CXX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 AC_PROG_F77 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 AC_PROG_LEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 AC_PROG_LIBTOOL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 AC_PROG_RANLIB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 AC_PROG_YACC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 AC_SUBST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 ACLOCAL_AMFLAGS. . . . . . . . . . . . . . . . . . . . . . . . . . . 22, 66

am_cv_sys_posix_termios . . . . . . . . . . . . . . . . . . . . . 20

AM_C_PROTOTYPES . . . . . . . . . . . . . . . . . . . . . . . 14, 18, 47 AM_CFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 AM_CONDITIONAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 AM_CONFIG_HEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 AM_CPPFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 AM_CXXFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 AM_ETAGSFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 AM_FFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 AM_GCJFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 AM_GNU_GETTEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

AM_HEADER_TIOCGWINSZ_NEEDS_SYS_IOCTL . . . . . . 18

AM_INIT_AUTOMAKE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

AM_INSTALLCHECK_STD_OPTIONS_EXEMPT . . . . . . . . 69

AM_JAVACFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 AM_LDFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27, 40 AM_MAINTAINER_MODE . . . . . . . . . . . . . . . . . . . 15, 66, 83 AM_MAKEINFOFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 AM_MAKEINFOHTMLFLAGS . . . . . . . . . . . . . . . . . . . . . . . . 59 AM_PATH_LISPDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 AM_PROG_GCJ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 AM_RFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 AM_RUNTESTFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 AUTOCONF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 AUTOM4TE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 AUTOMAKE_OPTIONS . . . . . . . . . . . . . . . . . . . . . . 47, 48, 68

B

bin_PROGRAMS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 bin_SCRIPTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 build_alias . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 BUILT_SOURCES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

C

CC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 CCAS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 CCASFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 CFLAGS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 check_LTLIBRARIES . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 check_PROGRAMS. . . . . . . . . . . . . . . . . . . . . . . . . . . . 27, 38 check_SCRIPTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 CLASSPATH_ENV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 CLEANFILES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 COMPILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

                                                CONFIG_STATUS_DEPENDENCIES. . . . . . . . . . . . . . . . . . 67

CONFIGURE_DEPENDENCIES . . . . . . . . . . . . . . . . . . . . . . 67 CPPFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 CXX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 CXXCOMPILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 CXXFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 CXXLINK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

D

data_DATA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 DATA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 50 DEFS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 DEJATOOL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 DESTDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 dist_. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 dist_lisp_LISP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 dist_noinst_LISP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 DIST_SUBDIRS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25, 63

                                                DISTCHECK_CONFIGURE_FLAGS. . . . . . . . . . . . . . . . . . . 64
                                                distcleancheck_listfiles . . . . . . . . . . . . . . . . 64, 86

DISTCLEANFILES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 distdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63, 76

                                                distuninstallcheck_listfiles . . . . . . . . . . . . . . . 64

DVIPS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

E

ETAGS_ARGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 ETAGSFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 EXPECT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 EXTRA_DIST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 EXTRA_PROGRAMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

F

F77 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 F77COMPILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 FFLAGS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 FLINK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

G

GCJFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 GTAGS_ARGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

H

HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 host_alias . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 host_triplet. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

I

include_HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 INCLUDES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 info_TEXINFOS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

J

JAVA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 JAVAC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 JAVACFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 JAVAROOT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

L

LDADD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 LDFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 lib_LIBRARIES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 lib_LTLIBRARIES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 LIBADD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 libexec_PROGRAMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 libexec_SCRIPTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 LIBOBJS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14, 34 LIBRARIES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LIBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 LINK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 lisp_LISP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 LISP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 54 localstate_DATA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 LTLIBOBJS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

M

m4_include . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

m4_include, distribution . . . . . . . . . . . . . . . . . . . . 62

MAINTAINERCLEANFILES . . . . . . . . . . . . . . . . . . . . . . . . 62 MAKE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 MAKEFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 MAKEINFO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 MAKEINFOFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 MAKEINFOHTML. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 man_MANS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 MANS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 59 MOSTLYCLEANFILES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

N

nodist_. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 noinst_HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 noinst_LIBRARIES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 noinst_LISP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 noinst_LTLIBRARIES . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 noinst_PROGRAMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 noinst_SCRIPTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

O

oldinclude_HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

P

PACKAGE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 PACKAGE, directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

PACKAGE, prevent definition . . . . . . . . . . . . . . . . . 19

pkgdata_DATA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 pkgdata_SCRIPTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 pkgdatadir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 pkginclude_HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 pkgincludedir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 pkglib_LIBRARIES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 pkglib_LTLIBRARIES . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 pkglib_PROGRAMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 pkglibdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 pkgpyexecdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 pkgpythondir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 PROGRAMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 4 pyexecdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 pythondir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 PYTHON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 56 PYTHON_EXEC_PREFIX . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 PYTHON_PLATFORM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 PYTHON_PREFIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 PYTHON_VERSION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

R

RFLAGS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 RUNTEST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 RUNTESTDEFAULTFLAGS. . . . . . . . . . . . . . . . . . . . . . . . . . 66 RUNTESTFLAGS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

S

sbin_PROGRAMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 sbin_SCRIPTS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 SCRIPTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 49 sharedstate_DATA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 SOURCES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27, 38 SUBDIRS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 SUFFIXES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 sysconf_DATA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

T

TAGS_DEPENDENCIES . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 target_alias. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 TESTS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 TESTS_ENVIRONMENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 TEXI2DVI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 TEXI2PDF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 TEXINFO_TEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 TEXINFOS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 57, 58 top_distdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63, 76

V

VERSION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

                                                VERSION, prevent definition . . . . . . . . . . . . . . . . . 19

W

WARNINGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 WITH_DMALLOC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 WITH_REGEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

X

XFAIL_TESTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

Y

YACC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

B.2 General Index

#

## (special Automake comment) . . . . . . . . . . . . . . . 2

–

--acdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 --add-missing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 --copy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 --cygnus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 –enable-debug, example . . . . . . . . . . . . . . . . . . . . . . . . 72

--enable-maintainer-mode . . . . . . . . . . . . . . . . . . . . 15

--force . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 --force-missing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 --foreign . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 --gnits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 –gnits, complete description . . . . . . . . . . . . . . . . . . . . 73 --gnu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 –gnu, complete description . . . . . . . . . . . . . . . . . . . . . 73 –gnu, required files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 --help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 16 --include-deps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 --libdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 --no-force . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 --output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 --output-dir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 --print-ac-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 --verbose . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 16 --version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 16 --warnings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 --with-dmalloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 --with-regex. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 -a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 -c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 -f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 -hook targets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 -i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 -I . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 -local targets. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 -module, libtool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 -o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 -v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 -W . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

.

.la suffix, defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

DATA primary, defined . . . . . . . . . . . . . . . . . . . . . . . 50 DEPENDENCIES, defined . . . . . . . . . . . . . . . . . . . . 28 HEADERS primary, defined. . . . . . . . . . . . . . . . . . . 50 JAVA primary, defined . . . . . . . . . . . . . . . . . . . . . . . . 55 LDFLAGS, defined . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 LDFLAGS, libtool . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 LIBADD primary, defined . . . . . . . . . . . . . . . . . . . . . 30 LIBADD, libtool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 LIBRARIES primary, defined . . . . . . . . . . . . . . . . . 30 LISP primary, defined. . . . . . . . . . . . . . . . . . . . . . . . . 54 LTLIBRARIES primary, defined . . . . . . . . . . . . . . 31 MANS primary, defined . . . . . . . . . . . . . . . . . . . . . . . 59 PROGRAMS primary variable . . . . . . . . . . . . . . . . . 3 PYTHON primary, defined. . . . . . . . . . . . . . . . . . . . 56 SCRIPTS primary, defined . . . . . . . . . . . . . . . . . . . . 49 SOURCES and header files. . . . . . . . . . . . . . . . . . . . 27 SOURCES primary, defined . . . . . . . . . . . . . . . . . . . 27 _SOURCES, default . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 _SOURCES, empty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 TEXINFOS primary, defined . . . . . . . . . . . . . . . . . . 57

A

AC_SUBST and SUBDIRS . . . . . . . . . . . . . . . . . . . . . . . . . 25 acinclude.m4, defined . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 aclocal program, introduction . . . . . . . . . . . . . . . . . . . 6 aclocal search path. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 aclocal’s scheduled death . . . . . . . . . . . . . . . . . . . . . . . 23 aclocal, extending . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 aclocal, Invoking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 aclocal, Options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 aclocal.m4, preexisting . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Adding new SUFFIXES. . . . . . . . . . . . . . . . . . . . . . . . 71 all . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 all-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 ALLOCA, special handling . . . . . . . . . . . . . . . . . . . . . . . 39 AM_CONDITIONAL and SUBDIRS . . . . . . . . . . . . . . . . . . 25 AM INIT AUTOMAKE, example use . . . . . . . . . . . 6 AM MAINTAINER MODE, purpose . . . . . . . . . . 83 ansi2knr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 ansi2knr and LIBOBJS . . . . . . . . . . . . . . . . . . . . . . . . . 48 ansi2knr and LTLIBOBJS . . . . . . . . . . . . . . . . . . . . . . . 48 Append operator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 autogen.sh and autoreconf . . . . . . . . . . . . . . . . . . . . . 34 Automake constraints . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Automake options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Automake requirements . . . . . . . . . . . . . . . . . . . . . 1, 12 Automake, invoking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Automake, recursive operation . . . . . . . . . . . . . . . . . . 2 Automatic dependency tracking . . . . . . . . . . . . . . . . 48 Automatic linker selection. . . . . . . . . . . . . . . . . . . . . . 46 autoreconf and libtoolize . . . . . . . . . . . . . . . . . . . . . . . 34 Auxiliary programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Avoiding path stripping . . . . . . . . . . . . . . . . . . . . . . . . 26

B

bootstrap.sh and autoreconf. . . . . . . . . . . . . . . . . . . . 34 BUGS, reporting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 BUILT SOURCES, defined . . . . . . . . . . . . . . . . . . . . 51

C

C++ support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 canonicalizing Automake variables. . . . . . . . . . . . . . . 4 cfortran. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 check. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 check primary prefix, definition. . . . . . . . . . . . . . . . . . 4 check-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 check_PROGRAMS example . . . . . . . . . . . . . . . . . . . . . . 38 clean. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 clean-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Comment, special to Automake . . . . . . . . . . . . . . . . . 2 Complete example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Conditional example, –enable-debug . . . . . . . . . . . . 72 conditional libtool libraries . . . . . . . . . . . . . . . . . . . . . 31 Conditional programs . . . . . . . . . . . . . . . . . . . . . . . . . . 29 Conditional subdirectories. . . . . . . . . . . . . . . . . . . . . . 24 Conditional SUBDIRS . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 Conditionals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 config.guess . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 configure.ac, from GNU Hello . . . . . . . . . . . . . . . . . . . 7 configure.ac, scanning . . . . . . . . . . . . . . . . . . . . . . . . . . 12 conflicting definitions . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Constraints of Automake . . . . . . . . . . . . . . . . . . . . . . . . 1 convenience libraries, libtool . . . . . . . . . . . . . . . . . . . 33 copying semantics. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 cpio example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 cvs-dist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 cvs-dist, non-standard example . . . . . . . . . . . . . . . . . . 1 CVS and generated files. . . . . . . . . . . . . . . . . . . . . . . . 81 CVS and third-party files . . . . . . . . . . . . . . . . . . . . . . 82 CVS and timestamps . . . . . . . . . . . . . . . . . . . . . . . . . . 80 Cygnus strictness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

D

DATA primary, defined . . . . . . . . . . . . . . . . . . . . . . . . 50 de-ANSI-fication, defined. . . . . . . . . . . . . . . . . . . . . . . 47 default _SOURCES. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 default source, Libtool modules example. . . . . . . . 39 definitions, conflicts . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 dejagnu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 depcomp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 dependencies and distributed files . . . . . . . . . . . . . . 84 Dependency tracking. . . . . . . . . . . . . . . . . . . . . . . . . . . 48 Dependency tracking, disabling. . . . . . . . . . . . . . . . . 48 dirlist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Disabling dependency tracking . . . . . . . . . . . . . . . . . 48 dist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 dist-bzip2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65, 68 dist-gzip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 dist-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63, 75 dist-shar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65, 68 dist-tarZ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65, 68 dist-zip. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65, 68 dist and nobase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 DIST_SUBDIRS, explained . . . . . . . . . . . . . . . . . . . . . . . 25 distcheck . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 distclean . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75, 84 distclean, diagnostic . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 distclean-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 distcleancheck. . . . . . . . . . . . . . . . . . . . . . . . . . . . 64, 84 distdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 dmalloc, support for . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 dvi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 dvi-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 DVI output using Texinfo . . . . . . . . . . . . . . . . . . . . . . 57

E

E-mail, bug reports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 EDITION Texinfo flag . . . . . . . . . . . . . . . . . . . . . . . . . 57 else . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 empty _SOURCES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 endif. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 Example conditional –enable-debug . . . . . . . . . . . . 72 Example of recursive operation . . . . . . . . . . . . . . . . . . 2 Example of shared libraries . . . . . . . . . . . . . . . . . . . . 31 Example, EXTRA PROGRAMS . . . . . . . . . . . . . . . . 3 Example, false and true . . . . . . . . . . . . . . . . . . . . . . . . . 8 Example, GNU Hello. . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Example, handling Texinfo files . . . . . . . . . . . . . . . . . 7 Example, mixed language . . . . . . . . . . . . . . . . . . . . . . 45 Example, regression test. . . . . . . . . . . . . . . . . . . . . . . . . 8 Executable extension. . . . . . . . . . . . . . . . . . . . . . . . . . . 48 Exit status 77, special interpretation . . . . . . . . . . . 65 Expected test failure . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 Extending aclocal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Extending list of installation directories . . . . . . . . . 3 Extension, executable . . . . . . . . . . . . . . . . . . . . . . . . . . 48 Extra files distributed with Automake . . . . . . . . . . 10 EXTRA , prepending . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 EXTRA prog SOURCES, defined . . . . . . . . . . . . . . 28 EXTRA PROGRAMS, defined . . . . . . . . . . . . . . 3, 29

F

false Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 FDL, GNU Free Documentation License. . . . . . . . 90 Files distributed with Automake . . . . . . . . . . . . . . . 10 First line of Makefile.am . . . . . . . . . . . . . . . . . . . . . . . . 2 FLIBS, defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 foreign strictness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Fortran 77 support. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 Fortran 77, mixing with C and C++ . . . . . . . . . . . . 44 Fortran 77, Preprocessing . . . . . . . . . . . . . . . . . . . . . . 43

G

generated files and CVS. . . . . . . . . . . . . . . . . . . . . . . . 81 generated files, distributed . . . . . . . . . . . . . . . . . . . . . 80 Gettext support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 gnits strictness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 gnu strictness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 GNU Gettext support. . . . . . . . . . . . . . . . . . . . . . . . . . 55 GNU Hello, configure.ac. . . . . . . . . . . . . . . . . . . . . . . . . 7 GNU Hello, example . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 GNU make extensions. . . . . . . . . . . . . . . . . . . . . . . . . . . 1 GNU Makefile standards . . . . . . . . . . . . . . . . . . . . . . . . 1 GNUmakefile including Makefile . . . . . . . . . . . . . . . 77

H

Header files in SOURCES . . . . . . . . . . . . . . . . . . . . . 27 HEADERS primary, defined. . . . . . . . . . . . . . . . . . . . 50 HEADERS, installation directories . . . . . . . . . . . . . 50 Hello example. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Hello, configure.ac . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 hook targets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 HP-UX 10, lex problems . . . . . . . . . . . . . . . . . . . . . . . 20 html . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 html-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 HTML installation, example. . . . . . . . . . . . . . . . . . . . . 3 HTML output using Texinfo . . . . . . . . . . . . . . . . . . . 57

I

id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 if . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 include . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 INCLUDES, example usage . . . . . . . . . . . . . . . . . . . . . 8 Including Makefile fragment . . . . . . . . . . . . . . . . . . . . 71 info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69, 75 info-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 install . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60, 75 Install hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 Install, two parts of . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 install-data. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 install-data-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 install-data-local . . . . . . . . . . . . . . . . . . . . . . . 61, 75 install-exec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60, 75 install-exec-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 install-exec-local . . . . . . . . . . . . . . . . . . . . . . . 61, 75 install-info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58, 69 install-info target . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 install-man . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60, 69 install-man target . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 install-strip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 Installation directories, extending list . . . . . . . . . . . . 3 Installation support . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 installcheck. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 installcheck-local . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 installdirs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61, 75 installdirs-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Installing headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 Installing scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 installing versioned binaries . . . . . . . . . . . . . . . . . . . . 75 Interfacing with third-party packages . . . . . . . . . . . 76 Invoking aclocal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Invoking Automake . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

J

Java support. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 JAVA primary, defined . . . . . . . . . . . . . . . . . . . . . . . . . 55 JAVA restrictions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

L

lex problems with HP-UX 10 . . . . . . . . . . . . . . . . . . . 20 lex, multiple lexers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 LIBADD primary, defined . . . . . . . . . . . . . . . . . . . . . . 30 libltdl, introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 LIBOBJS and ansi2knr . . . . . . . . . . . . . . . . . . . . . . . . . 48 LIBOBJS, special handling . . . . . . . . . . . . . . . . . . . . . . 39 LIBRARIES primary, defined . . . . . . . . . . . . . . . . . . 30 libtool convenience libraries . . . . . . . . . . . . . . . . . . . . 33 libtool libraries, conditional . . . . . . . . . . . . . . . . . . . . 31 libtool library, definition . . . . . . . . . . . . . . . . . . . . . . . 31 libtool modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 Libtool modules, default source example. . . . . . . . 39 libtool, introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 libtoolize and autoreconf . . . . . . . . . . . . . . . . . . . . . . . 34 libtoolize, no longer run by Automake . . . . . . . . . . 34 Linking Fortran 77 with C and C++ . . . . . . . . . . . . 44 LISP primary, defined. . . . . . . . . . . . . . . . . . . . . . . . . . 54 LN S example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 local targets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 LTLIBOBJS and ansi2knr . . . . . . . . . . . . . . . . . . . . . . . 48 LTLIBOBJS, special handling . . . . . . . . . . . . . . . . . . . . 34 LTLIBRARIES primary, defined . . . . . . . . . . . . . . . 31 ltmain.sh not found . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

M

Macro search path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Macros Automake recognizes . . . . . . . . . . . . . . . . . . . 13 make check . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 make clean support . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 make dist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 make distcheck. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 make distcleancheck. . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 make distuninstallcheck . . . . . . . . . . . . . . . . . . . . . . . . 64 make install support . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 make installcheck . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 Make rules, overriding. . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Make targets, overriding . . . . . . . . . . . . . . . . . . . . . . . . 2 Makefile fragment, including . . . . . . . . . . . . . . . . . . . 71 Makefile.am, first line . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 MANS primary, defined . . . . . . . . . . . . . . . . . . . . . . . . 59 many outputs, rules with. . . . . . . . . . . . . . . . . . . . . . . 87 mdate-sh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 missing, purpose . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 Mixed language example . . . . . . . . . . . . . . . . . . . . . . . 45 Mixing Fortran 77 with C and C++ . . . . . . . . . . . . . 44 Mixing Fortran 77 with C and/or C++. . . . . . . . . . 44 modules, libtool. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 mostlyclean . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 mostlyclean-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Multiple configure.ac files . . . . . . . . . . . . . . . . . . . . . . . 9 Multiple lex lexers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 multiple outputs, rules with . . . . . . . . . . . . . . . . . . . . 87 Multiple yacc parsers . . . . . . . . . . . . . . . . . . . . . . . . . . 41

N

no-dependencies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 no-dist-gzip. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 no-installinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 no-installman . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 no-texinfo.tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 nobase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 nobase and dist or nodist . . . . . . . . . . . . . . . . . . . 26 nodist and nobase . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 noinst primary prefix, definition . . . . . . . . . . . . . . . . . 4 noinstall-info target . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 noinstall-man target . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 Non-GNU packages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Non-standard targets. . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

O

Objects in subdirectory . . . . . . . . . . . . . . . . . . . . . . . . 36 Option, ansi2knr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Option, check-news . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Option, cygnus. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Option, dejagnu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Option, dist-bzip2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Option, dist-shar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Option, dist-tarZ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Option, dist-zip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Option, foreign . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Option, gnits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Option, gnu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Option, no-define . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Option, no-dependencies . . . . . . . . . . . . . . . . . . . . . . . 68 Option, no-dist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Option, no-dist-gzip. . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Option, no-exeext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 Option, no-installinfo . . . . . . . . . . . . . . . . . . . . . . . . . . 69 Option, no-installman . . . . . . . . . . . . . . . . . . . . . . . . . . 69 Option, no-texinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 Option, nostdinc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 Option, readme-alpha . . . . . . . . . . . . . . . . . . . . . . . . . . 69 Option, version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 Option, warnings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 Options, aclocal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Options, Automake . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Options, std-options . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 Overriding make rules . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Overriding make targets. . . . . . . . . . . . . . . . . . . . . . . . . 2 Overriding make variables . . . . . . . . . . . . . . . . . . . . . . . 2 overriding rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 overriding semantics . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

P

Path stripping, avoiding. . . . . . . . . . . . . . . . . . . . . . . . 26 pdf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 pdf-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 PDF output using Texinfo . . . . . . . . . . . . . . . . . . . . . 57 per-target compilation flags, defined . . . . . . . . . . . . 37 pkgdatadir, defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 pkgincludedir, defined . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 pkglibdir, defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 POSIX termios headers . . . . . . . . . . . . . . . . . . . . . . . . 20 Preprocessing Fortran 77 . . . . . . . . . . . . . . . . . . . . . . . 43 Primary variable, DATA . . . . . . . . . . . . . . . . . . . . . . . 50 Primary variable, defined. . . . . . . . . . . . . . . . . . . . . . . . 3 Primary variable, HEADERS . . . . . . . . . . . . . . . . . . 50 Primary variable, JAVA . . . . . . . . . . . . . . . . . . . . . . . . 55 Primary variable, LIBADD. . . . . . . . . . . . . . . . . . . . . 30 Primary variable, LIBRARIES . . . . . . . . . . . . . . . . . 30 Primary variable, LISP . . . . . . . . . . . . . . . . . . . . . . . . 54 Primary variable, LTLIBRARIES . . . . . . . . . . . . . . 31 Primary variable, MANS . . . . . . . . . . . . . . . . . . . . . . . 59 Primary variable, PROGRAMS . . . . . . . . . . . . . . . . . 3 Primary variable, PYTHON. . . . . . . . . . . . . . . . . . . . 56 Primary variable, SCRIPTS . . . . . . . . . . . . . . . . . . . . 49 Primary variable, SOURCES . . . . . . . . . . . . . . . . . . . 27 Primary variable, TEXINFOS . . . . . . . . . . . . . . . . . . 57 prog LDADD, defined . . . . . . . . . . . . . . . . . . . . . . . . . 27 Programs, auxiliary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Programs, conditional . . . . . . . . . . . . . . . . . . . . . . . . . . 29 PROGRAMS primary variable . . . . . . . . . . . . . . . . . . 3 PROGRAMS, bindir . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 Proxy Makefile for third-party packages . . . . . . . 77 ps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 ps-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 PS output using Texinfo . . . . . . . . . . . . . . . . . . . . . . . 57 PYTHON primary, defined. . . . . . . . . . . . . . . . . . . . . 56

R

Ratfor programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 README-alpha . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 rebuild rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66, 80 Recognized macros by Automake . . . . . . . . . . . . . . . 13 Recursive operation of Automake. . . . . . . . . . . . . . . . 2 recursive targets and third-party Makefiles …. 76 regex package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Regression test example . . . . . . . . . . . . . . . . . . . . . . . . . 8 Reporting BUGS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Requirements of Automake. . . . . . . . . . . . . . . . . . . . . 12 Requirements, Automake. . . . . . . . . . . . . . . . . . . . . . . . 1 Restrictions for JAVA . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Rule, install-info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 Rule, install-man . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 Rule, noinstall-info. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 Rule, noinstall-man . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 rules with multiple outputs. . . . . . . . . . . . . . . . . . . . . 87 rules, conflicting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 rules, overriding. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 rx package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

S

Scanning configure.ac . . . . . . . . . . . . . . . . . . . . . . . . . . 12 SCRIPTS primary, defined . . . . . . . . . . . . . . . . . . . . . 49 SCRIPTS, installation directories . . . . . . . . . . . . . . 50 Selecting the linker automatically . . . . . . . . . . . . . . 46 Shared libraries, support for. . . . . . . . . . . . . . . . . . . . 30 site.exp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 SOURCES primary, defined . . . . . . . . . . . . . . . . . . . . 27 Special Automake comment . . . . . . . . . . . . . . . . . . . . . 2 Strictness, command line . . . . . . . . . . . . . . . . . . . . . . . 10 Strictness, defined. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Strictness, foreign . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Strictness, gnits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Strictness, gnu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Subdirectories, building conditionally. . . . . . . . . . . 24 Subdirectory, objects in . . . . . . . . . . . . . . . . . . . . . . . . 36 SUBDIRS and AC_SUBST . . . . . . . . . . . . . . . . . . . . . . . . . 25 SUBDIRS and AM_CONDITIONAL . . . . . . . . . . . . . . . . . . 25 SUBDIRS, conditional . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 SUBDIRS, explained . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 suffix .la, defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 suffix .lo, defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 SUFFIXES, adding . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 Support for C++ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 Support for Fortran 77 . . . . . . . . . . . . . . . . . . . . . . . . . 43 Support for GNU Gettext . . . . . . . . . . . . . . . . . . . . . . 55 Support for Java . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

T

tags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 TAGS support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 Target, install-info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 Target, install-man. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 Target, noinstall-info . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 Target, noinstall-man . . . . . . . . . . . . . . . . . . . . . . . . . . 60 termios POSIX headers . . . . . . . . . . . . . . . . . . . . . . . . 20 Test suites . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 Tests, expected failure . . . . . . . . . . . . . . . . . . . . . . . . . 66 Texinfo file handling example. . . . . . . . . . . . . . . . . . . . 7 Texinfo flag, EDITION. . . . . . . . . . . . . . . . . . . . . . . . . 57 Texinfo flag, UPDATED . . . . . . . . . . . . . . . . . . . . . . . 57 Texinfo flag, UPDATED-MONTH. . . . . . . . . . . . . . 57 Texinfo flag, VERSION . . . . . . . . . . . . . . . . . . . . . . . . 57 texinfo.tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 TEXINFOS primary, defined . . . . . . . . . . . . . . . . . . . 57 third-party files and CVS . . . . . . . . . . . . . . . . . . . . . . 82 Third-party packages, interfacing with. . . . . . . . . . 76 timestamps and CVS . . . . . . . . . . . . . . . . . . . . . . . . . . 80 true Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

U

underquoted AC DEFUN . . . . . . . . . . . . . . . . . . . . . . 21 Uniform naming scheme. . . . . . . . . . . . . . . . . . . . . . . . . 3 uninstall . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61, 75 uninstall-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 uninstall-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 UPDATED Texinfo flag . . . . . . . . . . . . . . . . . . . . . . . . 57 UPDATED-MONTH Texinfo flag . . . . . . . . . . . . . . 57 user variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

V

variables, conflicting . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Variables, overriding . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 variables, reserved for the user . . . . . . . . . . . . . . . . . . 4 version.m4, example . . . . . . . . . . . . . . . . . . . . . . . . . . 67 version.sh, example . . . . . . . . . . . . . . . . . . . . . . . . . . 67 VERSION Texinfo flag . . . . . . . . . . . . . . . . . . . . . . . . . 57 versioned binaries, installing . . . . . . . . . . . . . . . . . . . 75

W

wildcards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

Y

yacc, multiple parsers . . . . . . . . . . . . . . . . . . . . . . . . . . 41 ylwrap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

Z

zardoz example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

Table of Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 General ideas . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2.1 General Operation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2.2 Strictness. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.3 The Uniform Naming Scheme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.4 How derived variables are named . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.5 Variables reserved for the user . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.6 Programs automake might require . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3 Some example packages . . . . . . . . . . . . . . . . . . . . . . . . . 6

3.1 A simple example, start to finish . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.2 A classic program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.3 Building true and false. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

4 Creating a Makefile.in . . . . . . . . . . . . . . . . . . . . . . . . . . 9

5 Scanning configure.ac . . . . . . . . . . . . . . . . . . . . . . . . . 12

5.1 Configuration requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 5.2 Other things Automake recognizes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 5.3 Auto-generating aclocal.m4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 5.4 aclocal options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 5.5 Macro search path. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

5.5.1 Modifying the macro search path: --acdir. . . . . . . . . . . . . . . . 17 5.5.2 Modifying the macro search path: -I dir . . . . . . . . . . . . . . . . . 17 5.5.3 Modifying the macro search path: dirlist. . . . . . . . . . . . . . . . 17

5.6 Autoconf macros supplied with Automake . . . . . . . . . . . . . . . . . . . . . 18 5.6.1 Public macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 5.6.2 Private macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

5.7 Writing your own aclocal macros. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 5.8 Handling Local Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 5.9 The Future of aclocal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

6 The top-level Makefile.am. . . . . . . . . . . . . . . . . . . . . . 24

6.1 Recursing subdirectories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 6.2 Conditional subdirectories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

6.2.1 Conditional subdirectories with AM_CONDITIONAL . . . . . . . . . . 25 6.2.2 Conditional subdirectories with AC_SUBST . . . . . . . . . . . . . . . . . 25 6.2.3 How DIST_SUBDIRS is used . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

7 An Alternative Approach to Subdirectories . . 26 8 Building Programs and Libraries . . . . . . . . . . . . . 26

8.1 Building a program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 8.1.1 Defining program sources . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 8.1.2 Linking the program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 8.1.3 Conditional compilation of sources . . . . . . . . . . . . . . . . . . . . . . . . 28

8.1.3.1 Conditional compilation using _LDADD substitutions . . 28 8.1.3.2 Conditional compilation using Automake conditionals . . 29

8.1.4 Conditional compilation of programs . . . . . . . . . . . . . . . . . . . . . . 29 8.1.4.1 Conditional programs using configure substitutions . . 29 8.1.4.2 Conditional programs using Automake conditionals… 30

8.2 Building a library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 8.3 Building a Shared Library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

8.3.1 The Libtool Concept. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 8.3.2 Building Libtool Libraries. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 8.3.3 Building Libtool Libraries Conditionally . . . . . . . . . . . . . . . . . . 31 8.3.4 Libtool Libraries with Conditional Sources . . . . . . . . . . . . . . . . 32 8.3.5 Libtool Convenience Libraries. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 8.3.6 Libtool Modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 8.3.7 LIBADD and LDFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 8.3.8 LTLIBOBJS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 8.3.9 Common Issues Related to Libtool’s Use . . . . . . . . . . . . . . . . . . 34

     8.3.9.1 required file `./ltmain.sh' not found . . . . . . . . . . . 34
     8.3.9.2 Objects created with both libtool and without . . . 34

8.4 Program and Library Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 8.5 Default _SOURCES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 8.6 Special handling for LIBOBJS and ALLOCA . . . . . . . . . . . . . . . . . . 39 8.7 Variables used when building a program . . . . . . . . . . . . . . . . . . . . . . . 39 8.8 Yacc and Lex support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 8.9 C++ Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 8.10 Assembly Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 8.11 Fortran 77 Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

8.11.1 Preprocessing Fortran 77 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 8.11.2 Compiling Fortran 77 Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 8.11.3 Mixing Fortran 77 With C and C++. . . . . . . . . . . . . . . . . . . . . 44

8.11.3.1 How the Linker is Chosen . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 8.12 Java Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 8.13 Support for Other Languages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 8.14 Automatic de-ANSI-fication . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 8.15 Automatic dependency tracking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 8.16 Support for executable extensions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

9 Other Derived Objects . . . . . . . . . . . . . . . . . . . . . . . . 49

9.1 Executable Scripts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 9.2 Header files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 9.3 Architecture-independent data files . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 9.4 Built sources . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

9.4.1 Built sources example. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 First try. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 Using BUILT_SOURCES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Recording dependencies manually . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Build bindir.h from configure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Build bindir.c, not bindir.h. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 Which is best? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

10 Other GNU Tools. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

10.1 Emacs Lisp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 10.2 Gettext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 10.3 Libtool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 10.4 Java . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 10.5 Python. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

11 Building documentation . . . . . . . . . . . . . . . . . . . . . . 57

11.1 Texinfo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 11.2 Man pages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

12 What Gets Installed . . . . . . . . . . . . . . . . . . . . . . . . . . 60

12.1 Basics of installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 12.2 The two parts of install. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 12.3 Extending installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 12.4 Staged installs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 12.5 Rules for the user . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

13 What Gets Cleaned. . . . . . . . . . . . . . . . . . . . . . . . . . . 62

14 What Goes in a Distribution . . . . . . . . . . . . . . . . 62

14.1 Basics of distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 14.2 Fine-grained distribution control. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 14.3 The dist hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 14.4 Checking the distribution. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 14.5 The types of distributions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

15 Support for test suites. . . . . . . . . . . . . . . . . . . . . . . . 65

15.1 Simple Tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 15.2 DejaGnu Tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 15.3 Install Tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

16 Rebuilding Makefiles . . . . . . . . . . . . . . . . . . . . . . . . . 66

17 Changing Automake’s Behavior . . . . . . . . . . . . . 68 18 Miscellaneous Rules . . . . . . . . . . . . . . . . . . . . . . . . . . 70

18.1 Interfacing to etags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 18.2 Handling new file extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 18.3 Support for Multilibs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

19 Include. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

20 Conditionals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

21 The effect of --gnu and --gnits . . . . . . . . . . . . . . 73

22 The effect of --cygnus . . . . . . . . . . . . . . . . . . . . . . . . 74

23 When Automake Isn’t Enough. . . . . . . . . . . . . . . 74

23.1 Extending Automake Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 23.2 Third-Party Makefiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76

24 Distributing Makefile.ins . . . . . . . . . . . . . . . . . . . . 78

25 Automake API versioning. . . . . . . . . . . . . . . . . . . . 78

26 Upgrading a Package to a Newer Automake Version. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

27 Frequently Asked Questions about Automake . . 80

27.1 CVS and generated files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 27.1.1 Background: distributed generated files . . . . . . . . . . . . . . . . . . 80 27.1.2 Background: CVS and timestamps. . . . . . . . . . . . . . . . . . . . . . . 80 27.1.3 Living with CVS in Autoconfiscated projects. . . . . . . . . . . . . 81 27.1.4 Third-party files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82

27.2 missing and AM_MAINTAINER_MODE . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 27.2.1 missing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 27.2.2 AM_MAINTAINER_MODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

27.3 Why doesn’t Automake support wildcards? . . . . . . . . . . . . . . . . . . . 83 27.4 Files left in build directory after distclean. . . . . . . . . . . . . . . . . . . . . 84 27.5 Why are object files sometimes renamed? . . . . . . . . . . . . . . . . . . . . . 86 27.6 Handling Tools that Produce Many Outputs . . . . . . . . . . . . . . . . . . 87

Appendix A Copying This Manual. . . . . . . . . . . . . . 90

A.1 GNU Free Documentation License. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 A.1.1 ADDENDUM: How to use this License for your documents . . 96

Appendix B Indices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

B.1 Macro and Variable Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 B.2 General Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99